### v2

# Perverse microsheaves

Abstract.
On a complex contact manifold, or complex symplectic manifold with weight-1 circle action, we construct a sheaf of stable categories carrying a $t$-structure which is locally equivalent to a microlocalization of the perverse $t$-structure.
1. Introduction
For a complex manifold $M$, let us write $sh_{\mathbb{C}-c}(M)$ for a derived category of sheaves on $M$, whose objects are each locally constant on the strata of a locally finite stratification by complex subvarieties. Perverse sheaves are those $F$ with the following property:
(1) $\dim_{\mathbb{C}}\{\,p\in M\,|\,H^{i}(p^{*}F)\neq 0\}\leq i\qquad\qquad\dim_{\mathbb{C}}\{\,p\in M\,|\,H^{i}(p^{!}F)\neq 0\}\leq i$
Here $p^{*}$ and $p^{!}$ are the restriction functors to the point $p$.
Perverse sheaves have played a pivotal role in many results in algebraic geometry and geometric representation theory. They turn to be natural both in terms of considerations of Frobenius eigenvalues in positive characteristic [3], and in terms of analytic considerations in characteristic zero, where they are the sheaves of solutions to regular holonomic differential equations, or more generally, D-modules [10, 21]. This latter equivalence highlights a key feature, not immediately apparent from the definition: despite being a seemingly arbitrarily demarcated subcategory of a category of complexes, perverse sheaves form an abelian category.
There are microlocal versions (living on $T^{*}M$ or $\mathbb{P}T^{*}M$) of the category of $D$-modules [25, 12], perverse sheaves [2, 27], and the equivalence between them [1, 23, 28]. More generally still, Kashiwara has constructed a sheaf of categories on any complex contact manifold, locally equivalent to the microlocalization of $D$-modules [11], see also [24]. The variant appropriate to conic complex symplectic geometry has allowed the methods of geometric representation theory to be extended beyond cotangent bundles to more general symplectic resolutions and similar spaces [14, 4].
The purpose of the present article is to construct categories of complex-constructible microsheaves on complex contact manifolds which naturally carry a perverse $t$-structure, globalizing the construction of Waschkies [27]. In the sequel we establish a Riemann-Hilbert equivalence with the canonical stack of $\mathcal{E}$-modules defined by Kashiwara [11].
Our starting point is the globalization [26, 22] of the microlocal sheaf theory of Kashiwara and Schapira [15]. We recall the relevant notions in Section 4. In brief, the theory takes as input a real contact or exact symplectic manifold $V$, a choice of symmetric monoidal stable presentable coefficient category $\mathcal{C}$, and certain topological ‘Maslov data’. Maslov data is by definition a choice of trivialization for a certain canonical obstruction $V\to B^{2}Pic(\mathcal{C})$, and in fact agrees with the choices made in defining Floer-theoretic invariants in the same target spaces. The output of the theory is a sheaf of stable categories $\mu sh_{V}$ on $V$. For a locally closed subset $X\subset V$, we write $\mu sh_{X}\subset\mu sh_{V}|_{X}$ for the subsheaf of full subcategories on objects locally supported in $X$. A Legendrian or conic Lagrangian $L\subset V$ determines an obstruction $L\to BPic(\mathcal{C})$, a choice of trivialization for which (‘secondary Maslov data’) yields an equivalence $\mu sh_{L}\cong Loc_{L}$ with the sheaf of categories of local systems along $L$. Here we show that in the complex case:
Theorem 1.1 (see Section 3).
Let $W$ be an exact complex symplectic manifold, and $\mathcal{C}$ the category of modules over a (discrete) commutative ring $R$. Then there is a canonical Maslov datum for $W$, with respect to which secondary Maslov data for a Lagrangian $L\subset W$ are identified with spin structures on $L$.
From this and the existing theory [22], we deduce:
Theorem 1.2 (see Section 6).
Let $W$ be an exact complex symplectic manifold. For a discrete commutative ring $R$, there is a canonical sheaf of stable $R$-linear categories $\mu sh_{W}$ on $W$. For a complex Lagrangian $L\subset X$, a spin structure on $L$ determines an equivalence $\mu sh_{L}\cong Loc_{L}$.
Taking $D$ to be a disk around a smooth Lagrangian point $x\in X$, we obtain a canonical (up to noncanonical automorphism) ‘microstalk’ functor $\mu_{x}:\mu sh_{X}(W)\to\mu sh_{X}(D)\xrightarrow{\sim}R-mod$.
We write $\mu sh_{W,\mathbb{C}-c}\subset\mu sh_{W}$ for the sheaf of full subcategories on objects whose (micro)support is a complex analytic Lagrangian subset of $W$. We define $(\mu sh_{W,\mathbb{C}-c})^{\leq 0}\subset\mu sh_{W,\mathbb{C}-c}$ as the sheaves of full subcategories on those objects all of whose microstalks, as elements of $R-mod$, have cohomology concentrated in degrees $\leq 0$. We define similarly $(\mu sh_{W,\mathbb{C}-c})^{\geq 0}$.
We do not know in general whether $((\mu sh_{W,\mathbb{C}-c})^{\leq 0},(\mu sh_{W,\mathbb{C}-c})^{\geq 0})$ is a $t$-structure on $\mu sh_{W,\mathbb{C}-c}$. However, we will use these subcategories to construct $t$-structures on certain related categories.
To introduce these other categories, we return to the basic definitions of microsheaves. For $M$ a manifold, we have the presheaf of categories $\mu sh^{pre}$ on $T^{*}M$ defined by
$\mu sh^{pre}(U):=sh(M)/\{F\,|\,ss(F)\cap U=\emptyset\}.$
We write $\mu sh$ for its sheafification.
Let us write $T^{*}M_{\mathbb{R}}$ for the topological space whose open sets are the $\mathbb{R}^{>0}$-invariant open subsets of $T^{*}M$, and $\gamma_{\mathbb{R}}:T^{*}M\to T^{*}M_{\mathbb{R}}$ for the continous map given by the set theoretic identity. Let us write $\mu_{\mathbb{R}}sh$ for the sheafification of $(\gamma_{\mathbb{R}})_{*}\mu sh^{pre}$. Because microsupports are conic, there is an isomorphism of the sheaf pullback $\gamma_{\mathbb{R}}^{*}\mu_{\mathbb{R}}sh\xrightarrow{\sim}\mu sh$. In fact, because $\mathbb{R}^{>0}$ is contractible, the presheaf preimage $\gamma_{\mathbb{R}}^{*,pre}\mu_{\mathbb{R}}sh$ is already a sheaf, so already isomorphic to $\mu sh$. That is, nothing is gained or lost by considering $\mu_{\mathbb{R}}sh$ instead of $\mu sh$.
By contrast, when $M$ is a complex manifold, consider similarly the topological space $T^{*}M_{[\mathbb{C}^{*}]}$ on $\mathbb{C}^{*}$-invariant open sets in $T^{*}M$, and define $\mu_{\mathbb{C}}sh$ etc. similarly. Now there are maps
$\gamma_{\mathbb{C}}^{*,pre}\mu_{\mathbb{C}}sh\to\gamma_{\mathbb{C}}^{*}\mu_{\mathbb{C}}sh\to\mu sh$
These maps are all fully faithful on sections [27, Sec. 2.4], but we do not know when or whether (or after imposing what microsupport conditions) they are essentially surjective.
Consider the projectivized cotangent bundle, $\pi:T^{*}M\setminus 0_{M}\to\mathbb{P}^{*}M$. Then we have an inclusion of sheaves of full subcategories:
(2) $\mathbb{P}\mu sh:=\pi_{*}\gamma_{\mathbb{C}^{*}}^{*,pre}\mu_{\mathbb{C}}sh\hookrightarrow\pi_{*}\mu sh$
More generally, for $V$ a complex contact manifold and $\pi:\tilde{V}\to V$ its symplectization, we will identify a sheaf of full subcategories $\mathbb{P}\mu sh_{V}\subset\pi_{*}\mu sh_{\tilde{V}}$ which agrees with (2) in local charts (see Definition 5.14). We write $\mathbb{P}\mu sh_{V,\mathbb{C}-c}\subset\mathbb{P}\mu sh_{V}$ for the subsheaf of full subcategories on objects with complex Legendrian microsupport; note $\mathbb{P}\mu sh_{V,\mathbb{C}-c}=\mathbb{P}\mu sh_{V}\cap\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}$.
For complex exact symplectic manifolds $(W,\lambda)$, we may use the contactization $(W\times\mathbb{C},\lambda+dz)$ to define a sheaf of categories $\mathbb{P}\mu sh_{W\times 0}$. We do not understand the relationship of this with $\mu sh_{W}$ in general. However, if the Liouville flow on $(W,\lambda)$ integrates to a weight-1 $\mathbb{C}^{*}$-action, then $\mathbb{C}^{*}$ naturally acts on $W\times\mathbb{C}$ by contactomorphism. Let $\gamma_{\mathbb{C}}$ be the set-theoretic identity on $W\times\mathbb{C}^{*}$ (resp. on $W$) where the source carries the Euclidean topology but the target is endowed with the $\mathbb{C}^{*}$ invariant topology. In Theorem 6.3, we show that there is a natural $\mathbb{Z}[t,t^{-1}]$-linear structure on $\lambda_{*}\mathbb{P}\mu sh_{W\times\{0\}}$, and there are fully faithful embeddings
(3) $(\gamma_{\mathbb{C}})_{*}\mathbb{P}\mu sh_{W\times\{0\}}\hookrightarrow(\gamma_{\mathbb{C}})_{*}\mu sh_{W}\otimes\mathbb{Z}[t,t^{-1}]$
(4) $\mu_{\mathbb{C}}sh_{W}:=(\gamma_{\mathbb{C}})_{*}\mathbb{P}\mu sh_{W\times\{0\}}\otimes_{\mathbb{Z}[t,t^{-1}]}\mathbb{Z}\hookrightarrow(\gamma_{\mathbb{C}})_{*}\mu sh_{W}$
We can now state our results on $t$-structures.
Theorem 1.3 (7.16).
Let $V$ be a complex contact manifold. Then the pair
$\mathbb{P}\mu sh_{V,\mathbb{C}-c}^{\geq 0}:=\mathbb{P}\mu sh_{V}\cap\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}^{\geq 0}\qquad\qquad\qquad\mathbb{P}\mu sh_{V,\mathbb{C}-c}^{\leq 0}:=\mathbb{P}\mu sh_{V}\cap\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}^{\leq 0}$
determines a $t$-structure on $\mathbb{P}\mu sh_{V,\mathbb{C}-c}$.
Theorem 1.4.
Let $W$ be a complex exact symplectic manifold whose Liouville vector field integrates to a weight-1 $\mathbb{C}^{*}$ action. Then the pair
$\mu_{\mathbb{C}}sh_{W,\mathbb{C}-c}^{\geq 0}:=\mu_{\mathbb{C}}sh_{W}\cap\mu sh_{W,\mathbb{C}-c}^{\geq 0}\qquad\qquad\qquad\mu_{\mathbb{C}}sh_{W,\mathbb{C}-c}^{\leq 0}:=\mu_{\mathbb{C}}sh_{W}\cap\mu sh_{W,\mathbb{C}-c}^{\leq 0}$
determines a $t$-structure on $\mu_{\mathbb{C}}sh_{W,\mathbb{C}-c}$. Moreover, the Hom sheaf of two objects in the heart is a ($\frac{1}{2}\dim W$-shifted) perverse sheaf.
Example 1.5.
Take $(W,\lambda)$ as in Theorem 1.4, and suppose in addition $(W,\operatorname{re}\lambda)$ is a Weinstein manifold admiting a complete collection of generalized cocores, so that the main theorem of [7] can be applied. Many examples coming from geometric representation theory – quiver varieties, hypertoric varieties, moduli of Higgs bundles, etc. satisfy this assumption; cf. [29]. Let $L$ be the core of $W$. Then $\mu sh_{L}(W)$ is equivalent to (the module category of) $Fuk(W)$ [7], and so Theorem 1.4 asserts we have constructed a $t$-structure on the full subcategory
$\mu_{\mathbb{C}}sh_{L}(W)\hookrightarrow\mu sh_{L}(W)\cong Fuk(W)$
and correspondingly the perverse microsheaves determine a full abelian subcategory of $H^{0}Fuk(W)$.
Let us conclude by explicitly advertising the question:
Definition 1.6.
Let $W$ be a complex exact symplectic manifold whose Liouville vector field integrates to a weight-1 $\mathbb{C}^{*}$ action. We say a complex conic Lagrangian $L\subset W$ has only complex global microsheaves if $\mu_{\mathbb{C}}sh_{L}(W)\hookrightarrow\mu sh_{L}(W)$ is essentially surjective.
Question 1.7.
Which conic complex Lagrangians have only complex global microsheaves?
We do not know that the answer is not “all of them”. (In fact, in a previous draft of this article, we blithely assumed this was the case.) Example 1.5 provides one motivation for the question, since in the notation there, whenever $L$ has only complex global microsheaves, we have constructed a $t$-structure on the Fukaya category of $W$.
Here is what little we know about 1.7. It is not difficult to see that $\mathbb{C}^{*}$-conic subsets of cotangent bundles are completely complex, and slightly more generally, that the same holds for Lagrangians which admit a cover by $\mathbb{C}^{*}$-conic subsets of cotangent bundles. (This implies the existence of $t$-structures on Fukaya categories of hypertoric varieties, ALE spaces, and a few other quiver varieties.) We do not presently know much more in general.
Acknowledgements. We wish to thank Roman Bezrukavnikov, Sanath Devalapurkar, Peter Haine, Benjamin Gammage, Justin Hilburn, Michael McBreen, Semon Rezchikov, Marco Robalo, Pierre Schapira and Filip Živanović for enlightening conversations.
LC was partially supported by Simons Foundation grant 385573 (Simons Collaboration on Homological Mirror Symmetry). DN was partially supported by NSF grant DMS-2101466. CK and VS were partially supported by: VILLUM FONDEN grant VILLUM Investigator 37814, and VS was supported in addition by Novo Nordisk Foundation grant NNF20OC0066298, Danish National Research Foundation grant DNRF157, and the USA NSF grant CAREER DMS-1654545.
2. Complex contact and symplectic manifolds
We review some standard properties of complex contact and symplectic manifolds. A classical reference in the contact setting is [16].
If $X$ is a complex manifold, its holomorphic tangent bundle is denoted by $\mathcal{T}_{X}$. The holomorphic cotangent bundle is denoted by $\Omega_{X}$, and its exterior powers are denoted by $\Omega^{k}_{X}:=\bigwedge^{k}\Omega_{X}$.
Definition 2.1.
A complex (or holomorphic) symplectic manifold is a complex manifold $X$ along with a closed holomorphic $2$-form $\omega\in H^{0}(X,\Omega^{2}_{X})$.
A complex symplectic manifold $(X,\omega)$ determines a family, parameterized by $\hbar\in\mathbb{C}^{*}$, of real symplectic manifolds $(X,\operatorname{re}(\hbar\omega))$. By an exact complex symplectic manifold, we mean a pair $(X,\lambda)$ where $X$ is a complex manifold and $\lambda$ is a holomorphic 1-form such that $d\lambda=\partial\lambda$ is symplectic.
Example 2.2.
Let $X$ be a complex manifold. Then the holomorphic cotangent bundle $\Omega_{X}$ carries a canonical holomorphic $1$-form $\lambda_{can,\mathbb{C}}=ydx$, where $(x_{1},\dots,x_{n},y_{1},\dots,y_{n})$ are canonical holomorphic coordinates. Then $(\Omega_{X},d\lambda_{can,\mathbb{C}})$ is a complex symplectic manifold.
Meanwhile, the real cotangent bundle $T^{*}M$ carries the canonical $1$-form $\lambda_{can}$. There is a natural identification $\Omega_{X}\to T^{*}M$ defined in local holomorphic coordinates by
$(x,y)\mapsto(\operatorname{re}(x),\operatorname{im}(x),\operatorname{re}(y),-\operatorname{im}(y)).$
One computes that this identification pulls back $\lambda_{can}$ to $\operatorname{re}(\lambda_{can,\mathbb{C}})$.
If $(X,\omega)$ is a complex symplectic manifold, a half-dimensional complex submanifold $L\subset X$ is said to be complex Lagrangian if $\omega|_{L}=0$. Clearly complex Lagrangian submanifolds are automatically (real) Lagrangian with respect to $\operatorname{re}(\hbar\omega)$, for all $\hbar\in\mathbb{C}^{*}$.
Definition 2.3.
A complex (or holomorphic) contact manifold is a complex manifold $V$ of complex dimension $12n+$ 1 along with a holomorphic hyperplane field $\mathcal{H}\hookrightarrow\mathcal{T}_{V}$ which is maximally non-integrable. Concretely, this means that if $\alpha\in\Omega_{V}(U)$ is a holomorphic $1$-form for which $\mathcal{H}=\ker\alpha$ in some local chart $U\subset V$, then $\alpha\wedge(d\alpha)^{n}\neq 0$.
Given a complex contact manifold $(V,\xi)$, there is a holomorphic line bundle $\mathcal{T}_{V}/\mathcal{H}\to V$. Contact forms are naturally sections of $(\mathcal{T}_{V}/\mathcal{H})^{\vee}$. Note the existence of holomorphic global sections is a rather more stringent condition than the corresponding condition in real smooth geometry; correspondingly complex contact manifolds typically do not have global contact forms.
The bundle $(\mathcal{T}_{V}/\mathcal{H})^{\vee}$ is naturally a holomorphic sub-bundle of $\Omega_{V}$ (indeed, for $v\in V$, we have $(\mathcal{T}_{V}/\mathcal{H})^{\vee}_{v}=\{\alpha\in\Omega_{V,v}\mid\alpha(\xi)=0\}\subset\Omega_{V,v}$). We consider the $\mathbb{C}^{*}$-bundle
$\pi:\tilde{V}:=(\mathcal{T}_{V}/\mathcal{H})^{\vee}\setminus 0_{V}\to V$
which we call the complex symplectization of $V$ and let $\lambda_{\tilde{V}}$ denote the pullback of $\lambda_{can,\mathbb{C}}$ under the inclusion $\tilde{V}\hookrightarrow\Omega_{V}$.
We also consider the projectivized $S^{1}$-bundle $p:\tilde{V}/\mathbb{R}_{+}\to V$ and set $\xi_{\hbar}:=\ker(\operatorname{re}(\hbar\lambda_{\tilde{V}}))$ for $\hbar\in\mathbb{C}^{*}$. The relation between these spaces is summarized by the following diagram:
[…]
Lemma 2.4.
With the above notation:
  1. (i)
    $(\tilde{V},\hbar\lambda_{\tilde{V}})$ is an exact complex symplectic manifold
  2. (ii)
    $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$ is a real contact manifold
  3. (iii)
    $(\tilde{V},\operatorname{re}(\hbar\lambda_{\tilde{V}}))$ is canonically isomorphic to the real symplectization of $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$. (This isomorphism intertwines the $1$-forms and the $\mathbb{R}_{+}$-bundle structure over $\tilde{V}/\mathbb{R}_{+}$).
Proof.
We compute $\hbar\lambda_{\tilde{V}}$ locally by choosing a holomorphic contact $1$-form $\alpha$ defined on some open set $U\subset V$. Such a choice induces a holomorphic embedding
$\iota_{\alpha}:\mathbb{C}^{*}\times U$ $\hookrightarrow\tilde{V}\subset\Omega_{V}$
$(z,x)$ $\mapsto z\alpha_{x}$
To compute the pullback of $\hbar\lambda_{\tilde{V}}$ under $\iota_{\alpha}$, choose $(z,x)\in\mathbb{C}^{*}\times U$ and $d\iota_{\alpha}(v)\in T\tilde{V}_{z\alpha_{x}}\subset T\Omega_{V}$. Then we have $\hbar\lambda_{\tilde{V}}(d\iota_{\alpha}(v))=\hbar\lambda_{can,\mathbb{C}}(d\iota_{\alpha}(v))=z\hbar\alpha_{x}(d\pi\circ d\iota_{\alpha}(v))$. Hence
(6) $\iota_{\alpha}^{*}\hbar\lambda_{\tilde{V}}=z\hbar\alpha.$
Both (i) and (ii) can be checked immediately using (6). The proof of (iii) is an exercise in chasing definitions so is omitted.
Example 2.5 (Example 2.2 continued).
The complex projectivization $V:=(\Omega_{X}-0_{X})/\mathbb{C}^{*}$ is a complex contact manifold with respect to $\lambda_{can,\mathbb{C}}$ and we have $\tilde{V}=(\Omega_{X}-0_{X})$. The associated real projectivization $\tilde{V}/\mathbb{R}_{+}$ carries a circle of real contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.
Example 2.6.
Given an exact complex symplectic manifold $(X,\lambda)$, then $(X\times\mathbb{C},\lambda+dz)$ is a complex contact manifold. The contact form $\lambda+dz$ defines a section of the $\mathbb{C}^{*}$-bundle $\widetilde{X\times\mathbb{C}}=X\times\mathbb{C}\times\mathbb{C}^{*}\to X\times\mathbb{C}$. Similarly, $\operatorname{re}(\hbar(\lambda+dz))$ defines a section of $\widetilde{X\times\mathbb{C}}/\mathbb{R}^{+}=X\times\mathbb{C}\times S^{1}$.
Observe that there is a fiber-preserving $\mathbb{C}^{*}$-action on $\tilde{V}$: over some fiber $\tilde{V}_{v}\subset\Omega_{V,v}$, it sends $\alpha\mapsto z\alpha$ for $z\in\mathbb{C}^{*}$ and $v\in V$. For concreteness, we write $z=e^{t+i\theta}$ for $(t,\theta)\in\mathbb{R}_{+}\times S^{1}$ and let $\partial_{t},\partial_{\theta}$ denote the vector fields on $\tilde{V}$ generated by the $\mathbb{R}_{+}$ and $S^{1}$ actions.
We have the following morphisms of bundles over $\tilde{V}$:
(7) $\pi^{*}\xi\equiv\operatorname{ker}(\lambda_{\tilde{V}})/\langle\partial_{t},\partial_{\theta}\rangle\twoheadleftarrow\operatorname{ker}(\lambda_{\tilde{V}})\hookrightarrow\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\tilde{V}}))\hookrightarrow T\tilde{V}$
Lemma 2.7.
There is a splitting of vector bundles over $\tilde{V}$ (well-defined up to contractible choice)
$\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\tilde{V}}))=\partial_{t}\oplus\langle\partial_{\theta},X\rangle\oplus\pi^{*}\xi,$
where $X$ is a non-vanishing section of $\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\tilde{V}}))/\operatorname{ker}(\hbar\lambda_{\tilde{V}})$ and
  • (i)
    $\langle\partial_{\theta},X\rangle$ is a trivial $2$-dimensional real symplectic vector bundle with respect to $d(\operatorname{re}(\hbar\lambda_{\tilde{V}}))$
  • (ii)
    $\pi^{*}\xi$ is a complex symplectic vector bundle with respect to $d(\lambda_{\tilde{V}})$
Proof.
The existence of this splitting is essentially a restatement of (7). The other statements can be checked locally using (6). ∎
Given a complex contact manifold $(V,\mathcal{H})$ of complex dimension $12n+$ 1, a complex submanifold $L\subset V$ of complex dimension $n$ which is everywhere tangent to $\mathcal{H}$ is said to be a complex Legendrian.
Lemma 2.8.
If $L\subset(V,\xi)$ is complex Legendrian, then its preimage under $\tilde{V}\to V$ is denoted by $\tilde{L}$ and is complex exact Lagrangian. The quotient $\tilde{L}/\mathbb{R}_{+}\subset\tilde{V}/\mathbb{R}_{+}$ is a real Legendrian with respect to any of the $\mathbb{C}^{*}$ of contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.
3. Grading and orientation data
The material in this section is mostly algebraic topology. We will need it in Section 5 when we discuss how the condition of being a complex contact manifold/complex Legendrian interacts with orientations/gradings in microlocal sheaf theory.
Consider the following inclusions of groups ($n/2$ ones defined only when $n$ even):
${U(n/2,\mathbb{H})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${U(n/2)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\sqrt{SU}(n)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${U(n)}$${O(n,\mathbb{R})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Here $\sqrt{SU}(n)$ is defined as the kernel of $U(n)\xrightarrow{det^{2}}U(1)$. Note that $U(n)$ is the maximal compact subgroup of both $Sp(2n,\mathbb{R})$ and $GL(n,\mathbb{C})$. Meanwhile $U(n/2,\mathbb{H})$ is also known as the ‘compact symplectic group’ $Sp(n/2)=Sp(n,\mathbb{C})\cap U(n)$, the maximal compact subgroup of $Sp(n,\mathbb{C})$.
Definition 3.1.
Consider a topological space $X$ carrying a hermitian bundle classified by a map $X\to BU(n)$. We give names to the following sorts of structures:
  • A grading (or grading datum) is a lift to $B(\sqrt{SU}(n))$
  • A polarization is a lift to $BO(n,\mathbb{R})$
  • A quaternionic structure is a lift to $BU(n/2,\mathbb{H})$
  • A complex polarization is a lift to $BU(n/2)$.
When $X$ is a symplectic manifold or contact manifold, by a polarization (etc.) on $X$, we always mean a polarization (etc.) on the symplectic tangent bundle of $X$ or the contact distribution.
Lemma 3.2.
A polarization induces a grading. A quaternionic structure induces a grading. A complex polarization induces both a quaternionic structure and a polarization, each of which induces the same grading. ∎
Remark 3.3.
We emphasize that the identifications in Definition 3.1 should be understood as holding up to coherent homotopy. So, for example, a polarization for $X\to BU(n)$ is a map $X\to BO(n)$ and a homotopy between the composition $X\to BO(n)\to BU(n)$ and the original structure map $X\to BU(n)$.
Remark 3.4.
Strictly speaking, a symplectic vector bundle $E$ on $X$ is classified by a map $X\to BSp(2n,\mathbb{R})$. However, $BU(n)\to BSp(2n,\mathbb{R})$ is a homotopy equivalence and we may always lift to some $X\to BU(n)$ uniquely up to contractible choice. (I.e. the space of Hermitian structures compatible with a given symplectic bundle is contractible.) Any Lagrangian distribution in $E$ will be totally real for any compatible Hermitian structure, hence define a reduction of structure to $O(n)$, i.e. what we have here called a polarization.
Likewise, a complex symplectic bundle carries a quaternionic structure, uniquely up to contractible choices. We typically omit discussion of such issues in the sequel. In particular, for $(X,\omega)$ is complex symplectic, we simply say $TX$ carries a quaternionic structure.
Remark 3.5.
Let us recall the relationship between null-homotopies and lifts. Recall that a fiber sequence $P\to Q\to R$ is a Cartesian diagram $P\xrightarrow{\sim}Q\times_{R}\bullet$, where $\bullet$ is a point. Mapping spaces preserve limits, so $Map(X,P)\xrightarrow{\sim}Map(X,Q)\times_{Map(X,R)}Map(X,\bullet)$. That is, given such a fiber sequence of pointed spaces and a map $X\to Q$, a null-homotopy of the composite map $X\to Q\to R$ is equivalent to a lift $X\to P$.
Example 3.6.
The sequence $O(n)\to U(n)\xrightarrow{det^{2}}S^{1}$ is a fiber sequence, which we may deloop to a fiber sequence $BO(n)\to BU(n)\xrightarrow{Bdet^{2}}BS^{1}$. So, a symplectic vector bundle $E$ admits a grading iff $02c_{1}(E)=$ 0. In particular, a symplectic manifold $X$ admits a grading iff $02c_{1}(X)=$ 0.
We consider the classical stablized compact groups
$Sp=\lim_{n\to\infty}U(n,\mathbb{H})\qquad\qquad U=\lim_{n\to\infty}U(n,\mathbb{C})\qquad\qquad O=\lim_{n\to\infty}O(n,\mathbb{R})$
and the Lagrangian Grassmannians
$LGr_{\mathbb{C}}=Sp/U\qquad\qquad LGr=U/O$
The natural inclusion $LGr_{\mathbb{C}}\to LGr$ is the limit of $U(n,\mathbb{H})/U(n,\mathbb{C})\hookrightarrow U(2n,\mathbb{C})/O(2n,\mathbb{R})$.
There are evident stable analogues of the notions of Def. 3.1, and the stable analogue of Lemma 3.2 also holds.
Lemma 3.7.
Let $(V,\xi)$ be a complex contact manifold and consider the real contact manifold $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$. Then the contact distribution $\xi_{\hbar}\to\tilde{V}/\mathbb{R}_{+}$ carries a stable quaternionic structure.
Proof.
It is equivalent to prove that $q^{*}\xi_{\hbar}$ carries a stable quaternionic structure, where $q:\tilde{V}\to\tilde{V}/\mathbb{R}_{+}$ is the quotient map. This follows from Lemma 2.7. ∎
We write $w_{2}:BO\to B^{2}(\mathbb{Z}/2\mathbb{Z})$ for the map classified by the second Stiefel-Whitney class.
Definition 3.8.
Let $X$ be a topological space and consider a map $\nu:X\to BU$, i.e. a stable complex or symplectic vector bundle on $X$. Orientation data is a null-homotopy of the composite map:
[…]
The canonical orientation data is the null-homotopy arising from the canonical null-homotopy of $BU\to B(U/O)\to B^{2}O$.
Evidently, the space of orientation data is identified (using the canonical orientation datum) with $Maps(X,B^{2}(\mathbb{Z}/2\mathbb{Z}))$; in particular, the choices up to homotopy are $H^{2}(X,\mathbb{Z}/2\mathbb{Z})$.
[…]
(8) $X\to BO\to BU\to B(U/O)\to B^{2}O\xrightarrow{Bw_{2}}B^{3}(\mathbb{Z}/2\mathbb{Z})$
there are two evident null-homotopies, coming from the compositions $BO\to BU\to B(U/O)$ and $BU\to B(U/O)\to B^{2}O$. We term the first the polarization orientation data; the second is what we called above the canonical orientation data.
Lemma 3.9.
Fix a stable polarization $X\to BO$ of a given $X\to BU$. The space of homotopies between the canonical and polarization orientation data is equivalent to the space of lifts to $X\to BPin$.
Proof.
First we recall some general facts about null-homotopies and exact triangles. Given two null-homotopies $n_{1},n_{2}$ of a given map $f:Q\to R$, we produce a pointed map $[n_{1},n_{2}]\in Q\times S^{1}\to R$ by taking $S^{1}=[-\pi,\pi]$, taking the map $f$ on $Q\times 0$ and applying the null-homotopy $n_{1}$ along $[-\pi,0]$ and the null-homotopy $n_{2}$ along $[0,\pi]$. Note that $\operatorname{Hom}(Q\times S^{1},R)=\operatorname{Hom}(Q,\Omega R)$. A homotopy between $n_{1}$ and $n_{2}$ is a null-homotopy of $[n_{1},n_{2}]\in\operatorname{Hom}(Q,\Omega R)$.
Suppose now given any exact triangle $P\to Q\to R\to BP$ in a stable category (for the application here, the stable category of spectra). The compositions $P\to Q\to R$ and $Q\to R\to BP$ give two null-homotopies of the composite map $P\to BP$. The definition of exact triangles [19, Def. 1.1.2.11] promises that the comparison of these two null-homotopies is identified with the identity of $\operatorname{Hom}(P,\Omega BP)=\operatorname{Hom}(P,P)$.
[…]
given by $[s\circ n_{1}\circ p,s\circ n_{2}\circ p]\in\operatorname{Hom}(X,\Omega BS)$ is identified with $s\circ p\in\operatorname{Hom}(X,S)$. Here, $X$ need only be a space, not a spectrum.
Now specializing to (8), we find that the space of homotopies between polarization and canonical orientation data is identified with the space of null-homotopies of $X\to BO\xrightarrow{w_{2}}B^{2}(\mathbb{Z}/2\mathbb{Z})$, i.e., the space of lifts to $X\to BPin$. ∎
Definition 3.10.
Let $V$ be a symplectic or contact manifold. Fix a polarization / grading / orientation data $m$ on $V$.
Consider $L\subset V$ a Lagrangian or Legendrian. There is a canonical isomorphism of $T^{*}L$ with the restriction of the symplectic tangent bundle or contact distribution. We write $\phi_{L}$ for the fiber polarization.
A relative polarization for $L$ is a homotopy between $m|_{L}$ and $\phi_{L}$. Relative grading / orientation data are similarly homotopies between the grading / orientation data induced by $m|_{L}$ and $\phi_{L}$.
Note that the obstruction to existence of a relative grading is a class in $[L,B\mathbb{Z}]=H^{1}(L,\mathbb{Z})$, and the space of relative gradings is a torsor for $Map(L,\mathbb{Z})$, hence in particular, is discrete. Thus we simply ask whether relative gradings are equal, rather than discuss homotopies between them.
Lemma 3.11.
Let $V$ is a complex contact manifold, $\tilde{V}\to V$ the corresponding real contact manifold equipped with the canonical grading constructed in Lemma 3.2. Let $L\subset V$ be a smooth complex Legendrian. Then there is a canonical choice for the relative grading of $\tilde{L}\subset\tilde{V}$.
Proof.
On $L$, the fiber polarization provides a complex polarization of the restriction of the contact distribution, which, by Lemma 3.2, agrees with the canonical grading. ∎
Lemma 3.12.
Fix the canonical orientation data on $V$. Then relative orientation data for $L\subset V$ is equivalent to a pin structure on $L$. (If $L$ is oriented, then a pin structure on $L$ is equivalent to a spin structure on $L$ compatible with the given orientation.)
Proof.
This is a special case of Proposition 3.9. ∎
4. Microsheaves on real contact manifolds
Here we review ideas from the microlocal theory of sheaves as formulated for cotangent bundles of manifolds in [15] and globalized to arbitrary contact manifolds in [26, 22].
4.1. Sheaves on manifolds
Let $M$ be a real manifold. Fix a symmetric monoidal stable presentable category , $C$. The reader will not lose much of the point of the paper taking throughout $C$ to be the derived category of dg modules over some commutative ring $R$. We write $sh(M)$ for the (stable) category of sheaves on $M$ with values in $C$.
In this subsection we review ideas from Kashiwara and Schapira [15]. Often these were originally formulated for bounded derived categories, viewed as triangulated categories. Modern foundations [18, 19] allow one to work directly in the stable setting, and in addition for the boundedness hypothesis to be removed for many purposes; we do so when appropriate without further comment.
4.1.1. Microsupport
Given $F\in sh(M)$, we say that a smooth function $f:M\to\mathbb{R}$ has a cohomological $F$-critical point at $x\in M$ if $(j^{!}F)_{x}\neq 0$ for $j:\{f\geq 0\}\hookrightarrow M$ the inclusion. The microsupport of $F$ (also called the singular support) is the closure of the locus of differentials of functions at their cohomological $F$-critical points. We denote it by $ss(F)$.
The microsupport is easily seen to be conical and satisfy $ss(Cone(F\to G))\subset ss(F)\cup ss(G)$. A deep result of [15, Thm. 6.5.4] is that the microsupport is coisotropic (also called involutive; see [15, Def. 6.5.1]). For a conic subset $K\subset T^{*}M$, we write $sh_{K}(M)$ for the full subcategory on objects microsupported in $K$. For any subset $\Lambda\subset S^{*}M$, we write $sh_{\Lambda}(M):=sh_{\mathbb{R}_{+}\Lambda\cup 0_{M}}(M)$, with $0_{M}$ the zero section of the cotangent bundle.
The assignment $U\mapsto sh(U)$ defines a sheaf of categories on $M$; we denote it $sh$. Similarly, $U\mapsto sh_{K\cap T^{*}U}(U)$ defines a subsheaf of full subcategories, we denote it $sh_{K}$. Similarly, $sh_{\Lambda}$.
4.1.2. Constructibility
Definition 4.1.
A stratification of a topological space $X$ is a locally finite decomposition $X=\sqcup_{\alpha}X_{\alpha}$, where the $X_{\alpha}$ are pairwise disjoint locally closed subsets called strata. This decomposition must satisfy the frontier condition: the boundary $\overline{X}_{\alpha}\setminus X_{\alpha}$ is a union of other strata.
Theorem 4.2 (Thm. 8.4.2 of [15]).
Let $M$ be a real analytic manifold, and $F$ a sheaf on $M$. Then the following are equivalent:
  • There is a subanalytic stratification $M=\coprod M_{i}$ such that $F|_{M_{i}}$ is locally constant
  • $ss(F)$ is subanalytic (singular) Lagrangian
Sheaves satisfying these equivalent conditions are said to be $\mathbb{R}$-constructible. We write $Sh_{\mathbb{R},c}(M)$ for the category of $\mathbb{R}$-constructible sheaves.
Theorem 4.3 (Thm. 8.5.5 of [15]).
Let $M$ be a complex analytic manifold, and $F$ a sheaf on $M$. Then the following are equivalent:
  • There is a complex analytic stratification $M=\coprod M_{i}$ such that $F|_{M_{i}}$ is locally constant
  • $ss(F)$ is contained in a closed $\mathbb{C}^{*}$-conic subanalytic isotropic subset
  • $ss(F)$ is a complex analytic (singular) Lagrangian
Sheaves satisfying these equivalent conditions are said to be $\mathbb{C}$-constructible. We write $Sh_{\mathbb{C},c}(M)$ for the category of $\mathbb{C}$-constructible sheaves.
4.2. Microsheaves on cotangent bundles
We consider the presheaf of stable categories on $T^{*}M$:
(9) $\mu sh_{T^{*}M}^{pre}(U):=sh(M)/sh_{T^{*}M\setminus U}(M)$
Definition 4.4.
Let $\mu sh_{T^{*}M}$ be the sheaf of categories on $T^{*}M$ defined by sheafifying the presheaf $sh_{T^{*}M}^{pre}(U)$ in (9). Similarly, let $\mu sh_{S^{*}M}$ be the presheaf of categories on $S^{*}M$ obtained by sheafifying $\mu sh^{pre}_{S^{*}M}$.
For $F\in sh(M)$, one finds that the support of the image of $F$ in $\mu sh(\Omega)$ is $ss(F)\cap\Omega$. For this reason, for any object $G\in\mu sh(\Omega)$, we sometimes write $ss(G)$ for the support of $G$.
The sheaf $\mu sh_{T^{*}M}$ is conic, i.e. equivariant for the $\mathbb{R}^{+}$ scaling action. In particular, $\mu sh_{T^{*}M}|_{T^{\circ}M}$ is locally constant in the radial direction. This being a contractible $\mathbb{R}^{+}$, we may define a sheaf of categories $\mu sh_{S^{*}M}$ on $S^{*}M$ equivalently by pushforward or pullback along an arbitrary section of $T^{\circ}M\to S^{*}M$.
Definition 4.5.
For conic $K\subset T^{*}M$, let $\mu sh_{K}\subset\mu sh_{T^{*}M}$ be the subsheaf of full subcategories on objects supported in $K$. Similarly, for $\Lambda\subset S^{*}M$, let $\mu sh_{\Lambda}\subset\mu sh_{S^{*}M}$ be the subsheaf of full subcategories on objects supported in $\Lambda$.
The following is shown in [15, Prop. 6.6.1]:
Proposition 4.6.
Let $M$ be a manifold and let $N\subset M$ be a submanifold. Let $L=T^{*}_{N}M$ and fix a point $p\in L$. Then there is an equivalence of categories:
(10) $\omega_{p}:C$ $\xrightarrow{\sim}(\mu sh_{L,\phi})_{p}$
$A$ $\mapsto A_{N},$
where $A_{N}$ is the image in $\mu sh$ of the constant sheaf on $N$ with value $A$. The corresponding result of course holds, and we use the same notations, in $S^{*}M$.
Remark 4.7.
Obviously $\omega_{p}$ depends only on the germ of $M$ near the image of $p$; for this remark we regard $M,N$ as germs. Consider another triple $M^{\prime},N^{\prime},L^{\prime}=T^{*}_{N^{\prime}}M^{\prime},p^{\prime}\in L^{\prime}$. If $\dim M=\dim M^{\prime}$ and $\dim N=\dim N^{\prime}$, then there is a diffeomorphism $(M,N)\to(M^{\prime},N^{\prime})$ whose canonical lift to the cotangent bundle sends $p\to p^{\prime}$. (Indeed, note that $p$ may be moved arbitrarily in $L\setminus 0$ by diffeomorphisms of $M$ fixing $N$.) Obviously the construction of $\omega_{p}$ is functorial in such diffeomorphisms.
In any sheaf of categories $\mathcal{X}$ on a topological space $T$, given $F,G\in\mathcal{X}(T)$, the assignment $U\mapsto\operatorname{Hom}_{\mathcal{X}(U)}(F|_{U},G|_{U})$ is a sheaf on $T$; let us denote it as $\mathcal{H}om_{\mathcal{X}}(F,G)$. Per [15, Thm. 6.1.2]: [^1] More precisely, [15] show there is a morphism of this kind for $\mu sh^{pre}$ which is an isomorphism at stalks; the stated result follows upon sheafification. See [22] for some detailed discussions about the sheafification of $\mu sh^{pre}$.
(11) $\mathcal{H}om_{\mu sh}(F,G)=\mu hom(F,G)$
where $\mu hom$ is a functor explicitly defined in terms of standard sheaf theoretic operations, namely the Fourier-Sato transformation along the diagonal of the external Hom:
$\mu hom(F,G):=\mu_{\Delta}\operatorname{Hom}_{M\times M}(\pi_{1}^{!}F,\pi_{2}^{*}G)$
Definition 4.8.
For $M$ real analytic, we define $\mu sh_{\mathbb{R}-c}\subset\mu sh$ as the subsheaf of full subcategories on objects whose support is subanalytic and the closure of its smooth Lagrangian locus.
For $M$ complex analytic, we define $\mu sh_{\mathbb{C}-c}\subset\mu sh_{\mathbb{R}-c}$ as the subsheaf of full subcategories on objects whose support is complex analytic and the closure of its smooth Lagrangian locus.
4.3. Microsheaves on polarized real contact manifolds
A key tool of [15] is the functoriality of $\mu sh$ under (quantized) contact transformation [15, Sec. 7]. One readily obtains:
Lemma 4.9.
Suppose given a contact manifold $V$ and a contact embedding $\iota:V\times T^{*}D^{n}$, for some disk $D$. Then $\mu sh_{\iota(V)\times D^{n}}$ is locally constant along $D^{n}$, hence the pullback of a sheaf of categories along $V\times D^{n}\to V$.
The basic idea of [26] is to use Lemma 4.9 as a definition of $\mu sh_{V}$. Let $(V,\xi)$ be a contact manifold. Suppose given a (possibly positive codimensional) contact embedding $\iota:V\to S^{*}M$, and any Lagrangian distribution $\eta$ of the symplectic normal bundle to $\iota$. Let us write $V(\eta)$ for a thickening of $\iota(V)$ along $\eta$ in a tubular neighborhood. Then $\iota^{*}\mu sh_{V(\eta)}$ provides a sheaf of categories on $V$. By contact transformations, it is easy to see that this sheaf of categories is locally equivalent to microsheaves on a contact cosphere bundle of the same dimension.
To eliminate the dependence on $\iota$ and $M$, one notes that Gromov’s $h$-principle for contact embeddings implies the existence of high codimension embeddings of $V$ into the standard contact $\mathbb{R}^{2n+1}$ for large enough $n$, the space of which moreover becomes arbitrarily connected as $n\to\infty$. There remains the (continuous) dependence on a choice of polarization of the stable symplectic normal bundle (“stable normal polarization”).
Theorem 4.10.
[26] Given a contact manifold $(V,\xi)$ and a polarization $\eta$ of the stable symplectic normal bundle to $V$, there is a canonical sheaf of categories $\mu sh_{V,-\eta}$ on $V$, locally isomorphic to $\mu sh_{S^{*}M}$ for any $M$ of the appropriate dimension.
Above, $-\eta$ can be regarded as just a notational choice. To give it an actual meaning, note that $BO\to BU$ is a morphism of spectra, so given a stable symplectic vector bundle $E$ on some topological space $X$, say classified by some map $E:X\to BU$, and a polarization of $E$, i.e. lift to some $F:X\to BO$, then $-F$ gives a polarization of $-E$, where -- is the pullback by the canonical ‘inverse’ involution on $BO$ or $BU$.
In fact [22, Sec. 10.2], it is always possible to build a contact manifold, $V^{T^{*}LGr(\xi)}$, which is a bundle over $V$ with fibers the cotangent bundles to the Lagrangian Grassmannians of the contact distribution $\xi$. The fiberwise zero section is the Lagrangian Grassmannian bundle $V^{LGr(\xi)}$. The virtue of $V^{T^{*}LGr(\xi)}$ is that it has a canonical polarization of the contact distribution, hence a canonical stable normal polarization. Thus we define $\mu sh_{V^{T^{*}LGr(\xi)}}$, and, restricting supports, $\mu sh_{V^{LGr(\xi)}}$.
Let us explain how this recovers the previous notion. A polarization $\rho$ of the contact distribution is (by definition) a section $\rho:V\to V^{LGr(\xi)}$. Now, the normal bundle to $\rho(V)\subset V^{T^{*}LGr(\xi)}$ is $T^{*}LGr(\xi)$, this has the canonical polarization by the cotangent fiber. We may combine this with canonical polarization of the symplectic stable normal bundle of $V^{T^{*}LGr(\xi)}$ to obtain a polarization of the symplectic stable normal bundle of $V$. It’s an exercise to see that this polarization is canonically identified with $-\rho$. We conclude:
Lemma 4.11.
For $\rho$ a polarization of the contact distribution, there is a canonical isomorphism $\rho^{*}\mu sh_{V^{LGr(\xi)}}=\mu sh_{V,\rho}$.
Proof.
We may choose the embedding of $V$ by first embedding $V^{LGr(\xi)}$. To obtain $\mu sh_{V,\rho}$, we must then thicken $V$ along a polarization of its normal bundle which stabilizes to $-\rho$. We obtain such by thickening $V$ along $V^{LGr(\xi)}\subset V^{T^{*}LGr(\xi)}$ and then along the canonical polarization of the normal bundle to $V^{T^{*}LGr(\xi)}$. But the same procedure defines $\mu sh_{V^{LGr(\xi)}}$. ∎
In case $V=S^{*}M$, there is a polarization $\phi$ given by the cotangent fiber. Now we have two notions of $\mu sh_{S^{*}M}$: the original given by (9), and then the construction of Theorem 4.10.
Lemma 4.12.
Let $M$ be a manifold, and $\nu$ the stable normal bundle of $M$. Then the stable symplectic normal bundle to $S^{*}M$ is $\nu\oplus\nu^{*}$, and the stable normal polarization by $\nu$ is canonically identified with $-\phi$.
Corollary 4.13.
There is a canonical equivalence of categories $\mu sh_{S^{*}M}\cong\mu sh_{S^{*}M,\phi}$, where the LHS is defined by Equation 9 and the right hand side by Theorem 4.10.
Proof.
Said differently, we officially define $\mu sh$ by embedding into $J^{1}\mathbb{R}^{n}\subset S^{*}\mathbb{R}^{n+1}$, but the construction makes sense for any embedding into any cosphere bundle, e.g. the embedding of $S^{*}M$ into itself. We should check these give the same result.
So embed $i:M\hookrightarrow\mathbb{R}^{n}$; we denote the normal bundle $\nu$; note this is a representative of the stable normal bundle. Let $\lambda_{T^{*}\mathbb{R}^{n}}$ and $\lambda_{T^{*}M}$ be the canonical one forms on cotangent bundles. Any splitting $\sigma$ of the vector bundle map $T^{*}\mathbb{R}^{n}|_{M}\twoheadrightarrow T^{*}M$ will satisfy $\sigma^{*}\lambda_{T^{*}\mathbb{R}^{n}}=\lambda_{T^{*}M}$, hence define a contact embedding $\sigma:S^{*}M\hookrightarrow S^{*}\mathbb{R}^{n}$.
The symplectic normal bundle to $\sigma$ is the restriction of $T^{*}_{T^{*}M}T^{*}\mathbb{R}^{n}=\nu\oplus\nu^{*}$. Let $S^{*}M(\nu)$ be a thickening of $\sigma(S^{*}M)$ in the direction $\nu$. By definition, $\mu sh_{S^{*}M,-\nu}=\sigma^{*}\mu sh_{S^{*}M(\nu)}$. Lemma 4.12 gives $\phi=-\nu$.
Meanwhile, $i^{*}:sh(\mathbb{R}^{n})\to sh(M)$ microlocalizes to define a map $i^{*}:\sigma^{*}\mu sh_{S^{*}M(\nu)}\to\mu sh_{S^{*}M}$; using e.g. Lemma 4.9, one readily sees this is an isomorphism. ∎
4.4. Maslov data
It is evident from the construction that $\mu sh_{V^{LGr(\xi)}}$ is locally constant in the Lagrangian Grassmannian direction. We stabilize $\xi\to\xi\oplus T^{*}\mathbb{R}^{n}$; taking $n\to\infty$, we have a sheaf of categories $\mu sh_{V^{LGr}}$ on the (stable) Lagrangian Grassmannian bundle $V^{LGr}$, locally constant along the Lagrangian Grassmannian direction. In fact:
Theorem 4.14.
[22, Sec. 11] There is a map of infinite loop spaces $\mu:LGr\to BPic(C)$ such that the sheaf of categories $\mu sh_{V^{LGr}}$ descends to the $BPic(C)$ bundle over $V$ classified by the map
[…]
Definition 4.15.
By $C$-Maslov data for $V$, we mean a null-homotopy of the map $V\to B^{2}Pic(C)$.
A choice $\eta$ of $C$-Maslov data gives a trivialization of the $BPic(C)$ bundle to which Theorem 4.14 asserts that $\mu sh_{V^{LGr}}$ descends. We write $\mu sh_{V,\eta}$ for the sheaf of categories on $V$ obtained by the pullback along the zero section of this trivial bundle (of groups).
[…]
For a subset $X\subset V$ we write $\mu sh_{X,\eta}\subset\mu sh_{V,\eta}$ for the sheaf of full subcategories on objects supported in $X$. For any smooth Legendrian $L\subset V$, Theorem 4.14 plus the Weinstein neighborhood theorem imply $\mu sh_{L,\eta}\subset\mu sh_{V,\eta}$ depends only on $\eta|_{L}$ (and not on $V$).
Remark 4.16.
There is a canonical choice of Maslov data on (or in a neighborhood of) a real Legendrian $L$. Indeed, canonically $\xi|_{L}\cong T^{*}L$, and taking the fiber polarization $\phi_{L}$ gives a canonical null-homotopy of $L\to V\to BLGr$. By composition we obtain a canonical null-homotopy $\phi_{L}$ of $L\to B^{2}Pic(C)$.
Given Maslov data $\eta$ for $V$, we obtain another such null-homotopy $\eta|_{L}$, we say that a homotopy $\phi_{L}\sim\eta|_{L}$ is secondary Maslov data for $L$. Evidently, secondary Maslov data induces an isomorphism $\mu sh_{L,\eta|_{L}}\cong\mu sh_{L,\phi_{L}}$.
It is possible to further fix an isomorphism $\mu sh_{L,\phi_{L}}\cong Loc_{L}$, the right hand side being the sheaf of categories of local systems on $L$. To do so, consider the embedding $L=L\times 0\subset L\times\mathbb{R}$. Let $L^{+}\subset S^{*}(L\times\mathbb{R})$ be the positive conormal to $L$.
Observe that the fiber polarization of $S^{*}(L\times\mathbb{R})$ restricts to $\phi_{L^{+}}$. Taking the pushforward along the inclusion and then microlocalizing along the positive conormal gives a map $Loc_{L}\to\mu sh_{L}\cong\mu sh_{L,\phi_{L}}$, which we can check locally is an equivalence (e.g. using [15, Proposition 6.6.1]).
Remark 4.17.
The category of co-orientable contact manifolds is equivalent to the category of symplectic manifolds with a free transitive $\mathbb{R}_{>0}$ action scaling the symplectic form: given a contact manifold $V$, we may form the corresponding symplectic manifold $\mathbb{R}_{>0}V$ by taking the locus in $T^{*}V$ of all graphs of contact 1-forms respecting the co-orientation. An embedding $V\hookrightarrow S^{*}M$ always lifts to a $\mathbb{R}_{>0}$-equivariant embedding $\mathbb{R}_{>0}V\to T^{\circ}M=\mathbb{R}_{>0}S^{*}M$, uniquely up to shift by a function $f:V\to\mathbb{R}_{>0}$, and a thickening of $V$ canonically lifts to a thickening of $\mathbb{R}_{>0}V$, etc. Stable polarizations and Maslov data being topological, these notions are equivalent on $V$ and $\mathbb{R}_{>0}V$.
Thus by embedding, thickening, descending, etc., we find that from Maslov data $\eta$ on $V$, we obtain an $\mathbb{R}_{>0}$-equivariant sheaf of categories $\mu sh_{\mathbb{R}_{>0}V,\eta}$ on $\mathbb{R}_{>0}V$. It is of course canonically isomorphic to the pullback of $\mu sh_{V,\eta}$ along the projection $\mathbb{R}_{>0}V\to V$.
In the discussion thus far, we have been agnostic as far as the choice of the category $C$, and we have also not needed to compute the map $\mu_{C}:U/O\to BPic(C)$. We now turn to this question. Note first that given a map of symmetric monoidal stable categories $C\to D$, it is clear that $\mu_{D}$ is the composition of $\mu_{C}$ with the natural map $Pic(C)\to Pic(D)$.
In particular, when $R$ is an ordinary commutative ring ($R=\pi_{0}(R)$), the map $\mu_{R-mod}$ factors through $\mu_{\mathbb{Z}-mod}$. We have $Pic(\mathbb{Z}-mod)=\mathbb{Z}\oplus B(\mathbb{Z}/2\mathbb{Z})$, where $\mathbb{Z}/2\mathbb{Z}$ is really the group of invertible integers, {±$\{\pm 1\}$. The map $\mu_{\mathbb{Z}-mod}$ has been calculated:
Theorem 4.18.
[8] The map $\mu_{\mathbb{Z}-mod}$ of Theorem 4.14 is $LGr\xrightarrow{\det^{2}\oplus w_{2}}S^{1}\oplus B^{2}(\mathbb{Z}/2\mathbb{Z})$. That is, $\mathbb{Z}-mod$ Maslov data is precisely grading and orientation data.
It follows that grading and orientation data provide $R-mod$ Maslov data for any commutative ring $R$, although of course there may be $R-mod$ Maslov data which does not arise in this way.
Remark 4.19.
Let $C=\mathbb{Z}-mod$. Let $M$ be a manifold. Let $m_{can}$ be the Maslov datum whose grading comes from the fiber polarization, and whose orientation data is the canonical orientation data of Def. 3.8. By Lemma 3.9, Corollary 4.13, and Theorem 4.18, a spin structure on $M$ gives an equivalence of sheaves of categories on $T^{\circ}M$: $\mu sh_{S^{*}M}\cong\mu sh_{S^{*}M,m_{can}}$ where the right hand side is defined by Equation (9) and the LHS by Theorem 4.14.
If $M=N\times\mathbb{R}$ is not spin, then in fact $\mu sh_{S^{*}M}\not\cong\mu sh_{S^{*}M,m_{can}}$. Indeed, in this case the conormal Legendrian to $N\times 0$ can be seen to support a microsheaf in $\mu sh_{S^{*}M}$ but not in $\mu sh_{S^{*}M,m_{can}}$.
This corresponds to the observation of Kragh that $Fuk(T^{*}M)\neq Loc(M)$ when $M$ is not spin [17, Remark 1.3].
Remark 4.20.
We return to the case of general $C$. Consider first the universal case $C=\mathbb{S}-mod$ of spectra. The map $\mu_{\mathbb{S}-mod}$ was calculated in [9]; the result is that it is a delooping of the $J$-homomorphism [9]. The map $\mathbb{S}\to\pi_{0}(\mathbb{S})=\mathbb{Z}$ induces $Pic(\mathbb{S}-mod)\to\pi_{\leq 1}Pic(\mathbb{S}-mod)=Pic(\mathbb{Z}-mod)$. In particular, the inclusion of the shift $\mathbb{Z}\to Pic(\mathbb{S}-mod)$ splits and we can write $Pic(\mathbb{S}-mod)=\mathbb{Z}\oplus Pic(\mathbb{S})_{0}$, where $Pic(\mathbb{S})_{0}$ is the identity connected component. (It is known that $Pic(\mathbb{S})_{0}=B\mathbb{S}^{\times}$ for some $\mathbb{S}^{\times}$ with a natural interpretation as “the units of $\mathbb{S}$”.) We write $w_{\mathbb{S}-mod}$ for the composition $LGr\to BPic(\mathbb{S}-mod)\to B^{2}\mathbb{S}^{\times}$, and define $\mathbb{S}-mod$ orientation data to be a null-homotopy of the appropriate composition with $Bw_{\mathbb{S}-mod}$. Tautologically, then, $\mathbb{S}$-mod Maslov data is the same as a grading and $\mathbb{S}-mod$ orientation data.
For more general $C$, let us write $Pic(C)_{0}$ for the identity connected component. Then the canonical map $\mathbb{S}-mod\to C$ induces a map $Pic(\mathbb{S})_{0}\to Pic(C)_{0}$, and we write $w_{C}$ for the composition of $w_{\mathbb{S}-mod}$ with the delooping of this map. We define $C$-orientation data to be a null-homotopy of the appropriate composition with $Bw_{C}$. Evidently a grading and $C$-orientation datum yield a $C$-Maslov datum, though not all $C$-Maslov data need arise in this way.
Still more generally, suppose given any partition $\Pi$ of the set of isomorphism classes of objects of $C$. We write $Pic(C)_{\Pi}\subset Pic(C)$ for the full subcategory on invertible elements which preserve the partition $\Pi$. By $(C,\Pi)$-orientation data, we mean a null-homotopy of the composite map to $BPic(C)_{\Pi}$. ($C$-orientation data being the case of the total partition). In particular, for $D\subset C$ a subcategory, we consider the partition into objects in versus not in $D$, and term the corresponding orientation data as $(C,D)$-orientation data.
5. Microsheaves on complex contact manifolds
Let $V$ be a complex contact manifold. Recall from Section 2 that we have maps
[…]
where $\tilde{V}$ is an exact complex symplectic manifold with $1$-form $\lambda_{\tilde{V}}$, and $\tilde{V}/\mathbb{R}_{+}$ carries a $\mathbb{C}^{*}$ of real contact forms $\xi_{\hbar}:=\operatorname{re}(\hbar\lambda_{\tilde{V}})$. Evidently $(\tilde{V},\operatorname{re}(\hbar\lambda_{\tilde{V}}))$ is the real symplectization of $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$. For specificity we fix $\hbar=0$ (but see Remark 5.7).
For a complex Legendrian $L\subset V$, we similarly write $\tilde{L}:=\pi^{-1}(L)$ and $\tilde{L}/\mathbb{R}_{+}:=p^{-1}(L)$.
Definition 5.1.
Let $V$ be a complex contact manifold and $\tilde{V}$ its symplectization. Per Lemma 3.2 and Remark 3.4, there is a canonical grading for $\tilde{V}$. Choose $C$-orientation data $o$ on $V$ and pull it back to $\tilde{V}$; combining this with with the canonical grading, we obtain Maslov data. We also write $o$ for this Maslov data, and correspondingly $\mu sh_{\tilde{V},o}$ for the corresponding sheaf of categories on $\tilde{V}$. In case $C=R-mod$ for a commutative ring $R$, then we have available the canonical orientation data; when we make this choice, we omit $o$ from the notation entirely, writing simply $\mu sh_{\tilde{V}}$.
For $\tilde{\Lambda}\subset\tilde{V}$ a $\mathbb{R}_{>0}$-invariant complex-analytic subset which is the closure of its smooth Lagrangian locus, we write $\mu sh_{\tilde{\Lambda}}\subset\mu sh_{\tilde{V}}$ for the subsheaf of full subcategories on objects supported in $\tilde{\Lambda}$. We write $\mu sh_{\tilde{V},\mathbb{C}-c,o}$ for the union of all such $\mu sh_{\tilde{\Lambda},o}$.
Remark 5.2.
For any fixed subanalytic Lagrangian $\Lambda$, the category $sh_{\Lambda}(M)$ is presentable, but $sh_{\mathbb{C}-c}(M)$ is not (arbitrary sums of constructible sheaves certainly need not be constructible). The situation for the microsheaf categories is entirely analogous. The distinction is rarely relevant: for example, in the present article, while we state theorems for $\mu sh_{\tilde{V},\mathbb{C}-c,o}$, their proofs quickly reduce to statements about $\mu sh_{\tilde{\Lambda},o}$.
Remark 5.3.
A closed and complex analytic subset $\tilde{\Lambda}\subset\tilde{V}$ is $\mathbb{R}_{>0}$-invariant iff it is $\mathbb{C}^{*}$-invariant. We use the formulation above because we may want to consider open sets which are $\mathbb{R}_{>0}$-invariant but not $\mathbb{C}^{*}$-invariant, and so need a corresponding local notion.
Remark 5.4.
Per Remark 4.17, we could equivalently have defined a category of microsheaves on the real contact manifold $\tilde{V}/\mathbb{R}_{>0}$ and pulled it back along $q:\tilde{V}\to\tilde{V}/\mathbb{R}_{+}$; our choice of formulation above is solely to avoid the minor cognitive dissonance of passing constantly to a non-complex manifold.
Remark 5.5.
One could also consider the subsheaf of full subcategories of $\mu sh_{\tilde{V}}$ on objects whose microsupport is complex-analytic, but this would e.g. not be closed under taking summands.
Lemma 5.6.
Let $f:U\to V$ be a complex contact embedding, and $\tilde{f}:\tilde{U}\to\tilde{V}$ its lift to the complex symplectization. Then there is a canonical isomorphism $\mu sh_{U,f^{*}o}=f^{*}\mu sh_{V,o}$.
In particular, when $o$ is the canonical orientation datum, there is a canonical isomorphism $\mu sh_{U}=f^{*}\mu sh_{V}$.
Proof.
It suffices to argue that the Maslov data of $V$ pulls back to the Maslov data of $U$, up to canonical homotopy. The orientation data is respected by definition, while the grading is canonically induced by the complex structure, hence behaves naturally under the above pullback. ∎
Remark 5.7.
$\tilde{V}$ is a $\mathbb{C}^{*}$-principal bundle over $V$, and the natural $\mathbb{C}^{*}$ action evidently satisfies $\hbar^{*}\lambda_{\tilde{V}}=\hbar^{-1}\lambda_{\tilde{V}}$ for $\hbar\in\mathbb{C}^{*}$. In particular, multiplication by $\hbar$ descends to a real contactomorphism $(\tilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\lambda)\cong(\tilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\lambda)$. The action of $\hbar$ is also canonically trivial on Maslov data, as this is pulled back from $V$. Thus by similar reasoning as in Lemma 5.6, there is a canonical isomorphism $\hbar^{*}\mu sh_{(\tilde{V},\hbar\lambda)}\cong\mu sh_{(\tilde{V},\lambda)}$.
Lemma 5.8.
Let $M$ be a complex manifold. Consider the complex contact manifold $V=\mathbb{P}T^{*}M$, and the corresponding $\tilde{V}=T^{\circ}M$, with canonical orientation data. A spin structure on $M$ determines an equivalence of sheaves of categories on $T^{\circ}M$: $\mu sh\cong\mu sh_{T^{\circ}M}$ where the left hand side is as in (9) and the right hand side is defined as in Definition 5.1.
Proof.
Follows from Corollary 4.13 by Lemma 3.11 and Lemma 3.12. ∎
Even if we had used some arbitrary choice of Maslov data $\mu$ on $V$, not necessarily coming from orientation data and the canonical complex grading, the (complex) Darboux theorem and general functoriality of microsheaves would nevertheless imply that $\mathbb{P}\mu sh_{V,\mu}$ is locally isomorphic to any $\mathbb{P}\mu sh_{\mathbb{P}T^{*}M}$ of the correct dimension, since the space of Maslov data, being a torsor for $Maps(V,BPic(C))$, is certainly connected when $V$ is contractible. However, the resulting isomorphism is not canonical, even locally. What we gain from using the canonical grading is:
Lemma 5.9.
Let $U$ be a contractible complex contact manifold, and $u:U\to V$ be any complex contact embedding. Then for any choices of $C$-orientation data $o,o^{\prime}$, there is an isomorphism $\mu sh_{U,o}\cong u^{*}\mu sh_{V,o^{\prime}}$ which is canonical up to a non-canonical natural transformation.
Proof.
Using Lemma 5.6, a homotopy of orientation data $u^{*}o^{\prime}\sim o$ will determine such an isomorphism. This is defined up to the action of loops in the space of orientation data; as $U$ is contractible, this is just (by definition of $C$-orientation data, see Remark 4.20) the connected component of the identity $Pic(C)_{0}\subset Pic(C)$, which acts by natural transformations. ∎
If $\Lambda\subset V$ is a Legendrian in a real contact manifold equipped with appropriate orientation and grading data, it follows, ultimately from Proposition 4.6, that there is a non-canonical isomorphism $(\mu sh_{\Lambda})_{p}\xrightarrow{\sim}C$ at any smooth point $p\in\Lambda$. Any choice of such an isomorphism is termed a microstalk functor; there is a $Pic(C)$ worth of choices.
In (10), we saw how to define a distinguished microstalk functor for Legendrians which occur as conormals to embedded submanifolds, where the orientation data is induced by the fiber polarization. However, leaving this class of submanifolds, the torsor of microstalk functors has some monodromy (classified by the map $\mu_{C}$ from Theorem 4.14), typically involving nontrivial shifts.
A central observation of this paper is that the situation is better in the complex world:
Definition-Proposition 5.10.
Let $V$ be a complex contact manifold equipped with some orientation data $o$. Let $L\subset V$ be a complex Legendrian, and $p\in L$ a smooth point with lift $\tilde{p}\in\tilde{L}$.
Choose a codimension zero embedding $\iota:Nbd(p)\hookrightarrow\mathbb{P}T^{*}M$ for a complex manifold $M$, such that the projection $L\cap Nbd(p)\to M$ is an embedding onto its image. Consider the composition of isomorphisms:
$C\xrightarrow{\omega_{\iota(\tilde{p})}}(\mu sh_{\iota(\tilde{L})})_{\iota(\tilde{p})}=(\mu sh_{\iota(\tilde{L}),\phi})_{\tilde{p}}\cong(\mu sh_{\tilde{L},o})_{\tilde{p}}$
here, the first arrow is (10), the second is (the $T^{\circ}M$ version of) Corollary 4.13, and the last is Lemma 5.9. We have seen the first two maps are independent of choices, and the final map depends on choices only through (noncanonical) natural transformation. We denote the composition by $\omega_{\tilde{p}}$, and, henceforth, when we refer in this context to microstalk functors, we mean only those of the form $\omega_{\tilde{p}}^{-1}$.
In particular, given $F\in(\mu sh_{\tilde{L},o})_{\tilde{p}}$, the object $\omega_{\tilde{p}}^{-1}(F)\in C$ is well defined (i.e. independent of the embedding $\iota$). We term this object the microstalk at $\tilde{p}$.
Definition 5.11.
Fix a subcategory $D$ of our coefficient category $C$, and fix $C$-orientation data $o$ (or more generally, $(C,D)$-orientation data). Let $V$ be a complex contact manifold. We write $(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{D}\subset\mu sh_{\tilde{V},\mathbb{C}-c,o}$ for the full subcategory on objects $F$ such that at every smooth point $p\in supp(F)$, one has $\omega_{\mathbb{C}}^{-1}(F_{\tilde{p}})\in D$.
Remark 5.12.
By the same argument as in Lemma 5.9, the isomorphism class of $\omega_{\mathbb{C}}^{-1}(F_{\tilde{p}})$ is constant in the smooth locus of the microsupport. So we need only check $\omega_{\mathbb{C}}^{-1}(F_{\tilde{p}})\in D$ at one point in each connected component of the smooth locus.
Legendrians give objects:
Proposition 5.13.
Let $V$ be a complex contact manifold and $L\subset V$ a complex Legendrian. Then relative orientation data on $L$ (e.g. a spin structure if we use the canonical orientation datum on $V$) determines an equivalence of categories $\mu sh_{\tilde{V},\mathbb{C}}|_{\tilde{L}}\cong Loc_{\tilde{L}}$ where $Loc_{\tilde{L}}$ is the sheaf of categories of local systems.
Proof.
As described in Remark 4.16, it suffices to provide secondary Maslov data. The grading part comes from Lemma 3.11. Lemma 3.12 explains why, for the canonical orientation datum, the corresponding notion of relative orientation data reduces is identified with spin structure. ∎
We now introduce sheaves of categories which will carry the microlocal perverse $t$-structure.
Definition 5.14.
Let $V$ be a complex-contact manifold. Given an object $K\in\pi_{*}\mu sh_{\tilde{V}}(U),U\subset V,$ and a point $p\in ss(K)\cap U$, we say that $K$ is representable at $p$ by a sheaf if there exists a contact embedding $\iota:Nbd(p)\hookrightarrow\mathbb{P}^{*}M$ such that (the image of) $K$ in $(\pi_{*}\mu sh_{T^{*}M-0_{M}})_{p}$ is represented by a sheaf $\mathcal{F}_{K}\in sh(M)$.222i.e. the image of $K$ in $(\pi_{*}\mu sh_{T^{*}M-0_{M}})_{p}$ coincides with the image of $\mathcal{F}_{K}$ under the quotient $sh(M)\to(\pi_{*}\mu sh_{T^{*}M-0_{M}})_{p}$. Here $M$ is any complex manifold and we recall that $\pi$ is the projection $T^{*}M-0_{M}=\widetilde{\mathbb{P}^{*}M}\to\mathbb{P}^{*}M$ – not projection to the base. We say that $K$ is pointwise representable by a sheaf if it is so at all points in its microsupport.
Remark 5.15.
By the discussion around [15, (11.4.8)] this definition is independent of the choice of the Darboux chart.
Definition 5.16.
Let $V$ be a contact manifold, $o$ orientation data, and $\Lambda$ a complex (singular) Legendrian. We denote by:
$\mathbb{P}\mu sh_{V,o}\subset\pi_{*}\mu sh_{\tilde{V},o}\qquad\qquad\mathbb{P}\mu sh_{\Lambda,o}\subset\pi_{*}\mu sh_{\pi^{-1}(\Lambda),o}\qquad\qquad\mathbb{P}\mu sh_{V,\mathbb{C}-c,o}\subset\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c,o}$
the sheaves of full subcategories of $\pi_{*}\mu sh_{\tilde{V},o}$ (resp. $\pi_{*}\mu sh_{\pi^{-1}(\Lambda),o}$, $\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c,o}$) on objects which are pointwise representable by sheaves.
Remark 5.17.
It is a tautology that $\mathbb{P}\mu sh_{V,\mathbb{C}-c}$ is a union over all complex (singular) legendrians $\Lambda$ of $\mathbb{P}\mu sh_{\Lambda}$.
Remark 5.18.
We could equivalently have pushed forward the corresponding sheaves of categories defined on the circle bundle $\tilde{V}/\mathbb{R}_{>0}$.
6. Microsheaves on complex exact symplectic manifolds
Let $X=(X,\lambda)$ be an exact complex symplectic manifold. Consider the contact thickening $(X\times\mathbb{C},\lambda+dz)$ and the associated symplectization
(13) $(X\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz)).$
Let $V_{\mathbb{R}}$ denote the Liouville vector field for $(X,\lambda)$ and let $I$ denote the almost-complex structure induced by complex multiplication. Integrating the flow of $Z$ in the first component and the flow of $V_{S^{1}}:=IV_{\mathbb{R}}$ in second, we obtain an action $A:\mathbb{R}\times\mathbb{R}\times X\to X$ which satisfies $A_{(t,\theta)}^{*}\lambda=e^{t+2\pi i\theta}\lambda$.
If $U\subset\mathbb{C}^{*}$ is contractible, then we can define a action $U\times X\to X$ by lifting $U$ to $\mathbb{R}\times\mathbb{R}$ by the covering map $(t,\theta)\mapsto e^{t+2\pi i\theta}$. In many situations, the flow in the $IZ$ direction factors through $k\mathbb{Z}\subset\mathbb{R},k\geq 1$; in this case, we can lift the action via the covering map $(t,\theta)\mapsto e^{k(t+2\pi i\theta)}$ to define a weight $k$ action of $\mathbb{C}^{*}$.
Let us now suppose that $U\subset\mathbb{C}^{*}$ is a ball containing $1$, and consider the induced weight-$1$ action $U\times X\to X,(w,x)\mapsto A_{w}(x)$, $A_{w}^{*}(\lambda)=w\lambda$. We record the following convenient change of variables:
(14) $(X\times\mathbb{C}\times U,\lambda+wdz)$ $\xrightarrow{\simeq}(X\times\mathbb{C}\times U,w(\lambda+dz))$
$(x,z,w)$ $\mapsto(A_{w^{-1}}(x),z,w)$
We let $\pi:X\times\mathbb{C}\times\mathbb{C}^{*}\to X\times\mathbb{C}$ be the projection which forgets the last component.
Recall that in 5.10, we show that, for a complex contact manifold $V$ and a complex Legendrian $L\subseteq V$, we can fix a canonical identification $(\mu sh_{\tilde{L}})_{\tilde{p}}\simeq C$ where $\tilde{L}$ is its Lagrangian lift and $\tilde{p}$ a smooth point. The above change of coordinates implies that a similar choice can be made for the symplectic case.
Lemma 6.1.
Suppose now that $L\subset X$ is a (possibly singular) conic complex-Lagrangian. Given $p\in L$, we have
$(\mu sh_{L})_{p}=\left(\mu sh_{\pi^{-1}(L\times\{0\})}\right)_{(p,0,1)}$
where the left hand side is defined as in [22, Section 9.1], for any exact symplectic manifolds, and the right hand side is defined in Definition 5.1, only for symplectizations.
Proof.
We have $\pi^{-1}(L\times\{0\})=L\times T^{\circ}_{0}\mathbb{C}=L\times\{0\}\times\mathbb{C}^{*}$. As a consequence of the local splitting (14), we have
(15) $\left(\mu sh_{\pi^{-1}(L\times\{0\})}\right)_{(p,0,1)}\simeq(\mu sh_{L})_{p}\otimes(\mu sh_{T^{\circ}_{0}\mathbb{C}})_{(0,1)}\simeq(\mu sh_{L})_{p}$
We now define the microstalk functor for exact complex-symplectic manifolds.
Definition 6.2.
Let $(X,\lambda)$ be an exact complex-symplectic manifold equipped with some orientation data $o$. Let $L\subset X$ a (germ of a) conical complex Lagrangian submanifold. For $p\in L$ a smooth point, we have a microstalk funtor defined as the composition
$C\to\left(\mu sh_{\pi^{-1}(L\times\{0\}),o}\right)_{(p,0,1)}=(\mu sh_{L,o})_{p}$
where the first arrow is the microstalk functor introduced in 5.10 and the equality is given by (15) above.
Proof of Theorem 1.2.
Per Lemma 3.2 and Remark 3.4, there is a canonical grading for $\tilde{V}$. Choose the canonical orientation data. Combining these, we obtain Maslov canonical data and hence a canonical sheaf of categories $\mu sh_{X}$. Note that here we are viewing $X=(X,\operatorname{re}\lambda)$ as a real exact symplectic manifold, so that one can define $\mu sh_{X}$ as in [22, Section 9.1]. More precisely, we identify $X$ with the subset $X\times\{0\}$ of its contactization $X\times\mathbb{R}$, and we then define $\mu sh_{X}$ as the restriction $\mu sh_{X\times\{0\}}|_{X\times\{0\}}$.
Finally, if $L$ is a smooth smooth conic complex spin Lagrangian, the canonical equivalence $\mu sh_{W,L}(-)\simeq loc(L)$ is established exactly like Proposition 5.13. ∎
Given an exact complex-symplectic manifold $(X,\lambda)$, consider the contactization $(X\times\mathbb{C},\lambda+dz)$. This is a complex-contact manifold, and so we can consider $\mathbb{P}\mu sh_{X\times\{0\}}\subset\mathbb{P}\mu sh_{X\times\mathbb{C}}$, which is the full subcategory on objects microsupported on $X\times\{0\}$.
We now have two sheaves of categories on $X$, namely $\mu sh_{X}$ and $\mathbb{P}\mu sh_{X\times\{0\}}$. They are not the same: for example, if $X$ is a point, then $\mu sh_{X}=C$ while $\mathbb{P}\mu sh_{X\times\{0\}}$ is the category of local systems on $\mathbb{C}^{*}$. In general, we do not know how to relate $\mu sh_{X}$ and $\mathbb{P}\mu sh_{X\times\{0\}}$. However:
Theorem 6.3.
When the Liouville vector field of $(X,\lambda)$ integrates to a weight-1 $\mathbb{C}^{*}$-action, there is a natural $\mathbb{Z}=\Omega S^{1}$-linear structure on $\mathbb{P}\mu sh_{X\times\{0\}}$, and there are fully faithful embeddings
(16) $\mathbb{P}\mu sh_{X\times\{0\}}\hookrightarrow\mu sh_{X}\otimes\Omega S^{1}$
(17) $\mu_{\mathbb{C}}sh_{X}:=\mathbb{P}\mu sh_{X\times\{0\}}\otimes_{\Omega S^{1}}\bullet\hookrightarrow\mu sh_{X}$
Of course, when working over a discrete ring, we may replace $\Omega S^{1}$ and $\bullet$ by their linearizations $\mathbb{Z}[\mathbb{Z}]=\mathbb{Z}[t,t^{-1}]$ and $\mathbb{Z}$.
Proof.
The change of variables (14) is global:
(18) $(\widetilde{X\times\mathbb{C}}=X\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz))$ $\xrightarrow{\simeq}(X\times\mathbb{C}\times\mathbb{C}^{*},\lambda+wdz)$
The Liouville structure on the right hand side is a product; we have the Künneth isomorphism:
$\mu sh_{X\times 0\times\mathbb{C}^{*}}=\mu sh_{X}\boxtimes\mu sh_{\mathbb{C}^{*}\subset T^{*}\mathbb{C}^{*}}.$
Pushing forward and composing with the embedding $\mathbb{P}\mu sh_{X\times 0}\hookrightarrow\pi_{*}\mu sh_{X\times 0\times\mathbb{C}^{*}}$ gives the first assertion, from which the second follows formally. ∎
Remark 6.4.
(17) amounts to taking invariants of a topological $\mathbb{C}^{*}$-action; see [20, Sec. 6]. With respect to the split symplectic form (i.e. the right hand side of (18)), this action is $\theta\cdot(x,z,w)\mapsto(x,\theta z,\theta^{-1}w)$. With respect to the coupled symplectic form (i.e. the left hand side of (18)), this corresponds to the diagonal $\mathbb{C}^{*}$ action $\theta(x,z,w)=(\theta\cdot x,\theta z,\theta^{-1}w)$. [^3] Note the slight abuse of notation: $\theta\cdot x$ refers to the image of $x\in X$ under the time $\theta$ flow of the Liouville vector field, which integrates by assumption to a $\mathbb{C}^{*}$ action. Observe that this is just the lift to the symplectization of the $\mathbb{C}^{*}$-action $\theta\cdot(x,z)=(\theta\cdot x,\theta z)$ on $X\times\mathbb{C}$ by contactomorphism.
Remark 6.5.
If $(X,\omega)$ admits a weight $\mathbb{C}^{*}$-action of weight $k>1$, we expect a related statement, except that $\mathbb{P}\mu sh_{X}$ should be a full subcategory of $(\mu sh_{X})^{\mathbb{Z}/k}$.
7. The perverse t-structure
7.1. t-structures
The notion of a $t$-structure on a triangulated category was introduced in [3]. We recall the definition and some basic properties.
Definition 7.1.
Let $\mathcal{T}$ be a triangulated category. A pair of subcategories $\mathcal{T}^{\leq 0},\mathcal{T}^{\geq 0}$ determine a t-structure if the following conditions are satisfied:
  • (i)
    For any $K^{\prime}\in T^{\leq 0}$ and $K^{\prime\prime}\in T^{\geq 0}$, we have $\operatorname{Hom}(K^{\prime},K^{\prime\prime}[-1])=0$.
  • (ii)
    If $K^{\prime}\in\mathcal{T}^{\leq 0}$ then $K^{\prime}[1]\in\mathcal{T}^{\leq 0}$; similarly if $K^{\prime\prime}\in\mathcal{T}^{\geq 0}$ then $K^{\prime\prime}[-1]\in\mathcal{T}^{\geq 0}$.
  • (iii)
    Given $K\in\mathcal{T}$, there exist $K^{\prime}\in\mathcal{T}^{\leq 0}$ and $K^{\prime\prime}\in\mathcal{T}^{\geq 0}$, and a distinguished triangle
    $K^{\prime}\to K\to K^{\prime\prime}[-1]\xrightarrow{[1]}$
We write $\mathcal{T}^{\geq n}:=\mathcal{T}^{\geq 0}[-n]$ and $\mathcal{T}^{\leq n}:=\mathcal{T}^{\leq 0}[-n]$. The heart of the t-structure is $\mathcal{T}^{\heartsuit}:=\mathcal{T}^{\leq 0}\cap\mathcal{T}^{\geq 0}\subset T$.
It is shown that there are truncation functors $\tau^{\leq n}:\mathcal{T}\to\mathcal{T}^{\leq n}$ and $\tau^{\geq n}:\mathcal{T}\to\mathcal{T}^{\geq n}$ which are right and left adjoint to the inclusions of the corresponding subcategories. The truncation functors commute in an appropriate sense (e.g. when composed with the inclusions so as to define endomorphisms of $\mathcal{T}$). Then $H^{0}:=\tau^{\leq 0}\tau^{\geq 0}=\tau^{\geq 0}\tau^{\leq 0}$ defines a map $\mathcal{T}\to\mathcal{T}^{\heartsuit}$, and one writes $H^{n}:\mathcal{T}\to\mathcal{T}^{\heartsuit}$ for the appropriate composition with the shift functor. Finally, $\mathcal{T}^{\heartsuit}$ is an abelian category, closed under extensions [3, Thm. 1.3.6].
The prototypical example is when $\mathcal{T}$ is a derived category of chain complexes, and $\mathcal{T}^{\leq 0}$ consists of the complexes whose cohomology is concentrated in degrees $\leq 0$, etc.
Let us recall a result about when $t$-structures pass to quotient categories.
Lemma 7.2 (Lem. 3.3 in [5]).
Let $\mathcal{T}^{\leq 0},\mathcal{T}^{\geq 0}$ determine a $t$-structure on $\mathcal{T}$. Let $\mathcal{I}\subset\mathcal{T}$ be a triangulated subcategory, closed under taking direct summands (“thick subcategory”) and let $Q:\mathcal{T}\to\mathcal{T}/\mathcal{I}$ be the Verdier quotient. Then:
  1. (1)
    $\mathcal{T}^{\leq 0}\cap\mathcal{I},\mathcal{T}^{\geq 0}\cap\mathcal{I}$ determine a $t$-structure if and only if $\tau_{\leq 0}\mathcal{I}\subset\mathcal{I}$
  2. (2)
    if the equivalent assertions of (1) hold, $Q(\mathcal{T}^{\leq 0}),Q(\mathcal{T}^{\geq 0})$ determine a $t$-structure if and only if $\mathcal{I}\cap\mathcal{T}^{\heartsuit}\subset\mathcal{T}^{\heartsuit}$ is a “Serre subcategory” (meaning it is closed under extensions, quotients and sub-objects).
The notion of $t$-structure is imported to the setting of stable categories in [19, Sec. 1.2]: By definition, a $t$-structure on a stable category is a $t$-structure on its homotopy category, which canonically carries the structure of a triangulated category. It is shown that the various properties of $t$-structures lift to the stable setting, in particular, the existence of truncation functors, and the fact that the full subcategory on objects in the heart is abelian. [^4] Let us avoid a possible source of confusion. One might think that, insofar as stable categories generalize dg categories, the heart could be expected to have, in its hom spaces, whatever corresponds to the positive ext groups. This depends on whether or not the stable category is viewed as a usual $\infty$-category, or as an $\infty$-category enriched in spectra. Indeed, the positive ext groups (in cohomological grading conventions) correspond to negative homotopy groups, so are only manifest after the (canonical) enrichment in spectra. Here however the statement about the heart should be understood in terms of the not enriched $\infty$-category. For consistency with [3] (and in contrast to [19]), we adopt cohomological conventions and write $H^{i}$ instead of $\pi_{-i}$.
Remark 7.3.
When $C$ is a presentable stable category, then if either $C^{\geq 0}$ or $C^{\leq 0}$ is presentable, then so is the other, and all truncation functors are colimit preserving [19, 1.4.4.13]. In this case, the subcategory of compact objects $C^{c}$ is stable under the truncation functors and inherits a $t$-structure. Indeed, $\tau^{\geq 0}$ is left adjoint to the corresponding inclusion, assumed colimit preserving, hence $\tau^{\geq 0}$ preserves compact objects. Taking cones, so does $\tau^{\leq 0}$.
We will study sheaves of $t$-structures on sheaves of categories.
Definition 7.4.
Let $M$ be a topological space and $\mathcal{F}$ a sheaf of stable categories on $M$. We say a pair of sheaves of full subcategories $\mathcal{F}^{\leq 0}$ and $\mathcal{F}^{\geq 0}$ define a $t$-structure on $\mathcal{F}$ if $\mathcal{F}^{\leq 0}(U)$ and $\mathcal{F}^{\geq 0}(U)$ define a $t$-structure on $\mathcal{F}(U)$ for all $U$.
Remark 7.5.
It is easy to see that the property that $\mathcal{F}^{\leq 0}$ and $\mathcal{F}^{\geq 0}$ define a $t$-structure may be checked on sections on any base of open sets. Indeed, regarding condition (i) and (ii) of Definition 7.1, it is immediate from the sheaf condition that vanishing of Homs and containment of subcategories can be checked locally.
Regarding (iii), the key point is that for any candidate $t$-structure satisfying (i) and (ii), the space of fiber sequences $K^{\prime}\to K\to K^{\prime\prime}$ as requested in (iii) is either empty or contractible. Indeed, first recall that using property (ii) to apply (i) to shifts, we find that the following strengthening of (i): for any $K^{\prime}\in T^{\leq 0}$ and $K^{\prime\prime}\in T^{\geq 0}$, the negative exts, aka positive homotopy groups of the hom space $\operatorname{Hom}(K^{\prime},K^{\prime\prime}[-1])$, must vanish. Now given any $K^{\prime}\to K\to K^{\prime\prime}[-1]$ and $L^{\prime}\to K\to L^{\prime\prime}[-1]$ both satisfying (iii), we obtain a canonical null-homotopy of the composition $K^{\prime}\to K\to L^{\prime\prime}[-1]$ hence lift of $K^{\prime}\to K$ to $K^{\prime}\to L^{\prime}$, etc.
Having learned this contractibility, if (iii) holds locally, then we can canonically glue the local exact triangles to obtain (iii) globally.
Since pullbacks commute with limits, $\mathcal{F}^{\heartsuit}:=\mathcal{F}^{\leq 0}\cap\mathcal{F}^{\geq 0}$ defines a sheaf of $(\infty,1)$-categories. As hearts of $t$-structures, these categories are abelian, in particular, 1-categories.
We recall a comparison between various notions of sheaves of categories. For covers $U=\bigcup U_{i}$, an $\infty$-categorical sheaf of categories $\mathcal{C}$ satisfies the condition that
$\mathcal{C}(U)=\lim\left(\prod_{i\in I}\mathcal{C}(U_{i})\rightrightarrows\prod_{i,j\in I}\mathcal{C}(U_{i}\cap U_{j})\mathrel{\begin{subarray}{c}\textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\end{subarray}}\prod_{i,j,k\in I}\mathcal{C}(U_{i}\cap U_{j}\cap U_{k})\cdots\right)$
The corresponding definition of stacks in the classical literature instead involves a limit of 1-categories taken in the $(2,1)$-category of ordinary categories. (Note that for limits in $(2,1)$ categories, the compatibility condition on triple overlaps is already strict, hence the terms after triple overlaps in the above diagram are irrelevant to the limit. In the literature they are correspondingly not included in the descent condition.
The situation is analogous to the fact that the descent condition for sheaves of sets is usually given truncated after the second term.)
These notions agree: $1$-categories are $1$-truncated objects of $(\infty,1)$-categories, and, for any category $\mathcal{C}$, the inclusion of $k$-truncated objects $\tau_{\leq k}\mathcal{C}\hookrightarrow\mathcal{C}$ is limit-preserving [18, Proposition 5.5.6.5]. We summarize:
Lemma 7.6.
The heart of a $t$-structure on a sheaf of stable categories is a stack of abelian categories.
7.2. The perverse t-structure on constructible sheaves
We now review from [15, Sec. 10.3] the microlocal description of the perverse $t$-structure on constructible sheaves.
Let $M$ be a complex manifold. For a Lagrangian subset $\Lambda\subset T^{*}M$, we write $\Lambda^{cl}$ for the locus of smooth points of $\Lambda$ where the map $\Lambda\to M$ has locally constant rank. For a point $p\in\Lambda^{cl}$, we write $r(p)$ for this rank. Fix a $t$-structure $C^{\leq 0},C^{\geq 0}$ on our coefficient category $C$.
We write [15, Def. 10.3.7] using the notation of (10):
Definition 7.7.
[15, Def. 10.3.7] Consider the following full subcategories of $sh(M)_{\mathbb{C}-c}$.
${}^{\mu}sh(M)_{\mathbb{C}-c}^{\leq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{cl}\implies\omega_{p}^{-1}F[-2r(p)]\in C^{\leq 0}\}$
${}^{\mu}sh(M)_{\mathbb{C}-c}^{\geq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{cl}\implies\omega_{p}^{-1}F[-2r(p)])\in C^{\geq 0}\}$
It is proved in [15, Thm. 10.3.12] that the above pair agrees with the usual definition of the perverse $t$-structure on $M$, and so in particular is a $t$-structure. [^5] The result is stated in [15] for $C$ the bounded derived category of modules over a ring and and $t$ the standard $t$-structure. However, the arguments given there (or in [3]) for the existence of the perverse $t$-structure depend only on the general properties of the six functor formalism, and the comparison between Definition 7.7 and the usual stalk/costalk wise definition of the perverse $t$-structure depends only on standard properties of microsupports.
Lemma 7.8.
If $\Lambda\subset T^{*}M$ is (possibly singular) subanalytic complex Lagrangian then Definition 7.7 induces a $t$-structure on $sh_{\Lambda}(M)$. Moreover, $sh_{\Lambda}(M)^{\heartsuit}$ is closed under extensions inside $sh(M)$.
Proof.
We wish to apply Lemma 7.2. We should therefore show that the images of the truncation functors, when applied to objects $F\in sh_{\Lambda}(M)$, remain in $sh_{\Lambda}(M)$. Suppose not; consider any smooth Lagrangian point in $ss(\tau_{\leq 0}F)\setminus ss(F)$. (There must be such a point by the assumption that the microsupport is complex analytic Lagrangian.) But by construction, the truncation functors commute with $\omega_{p}$, whence it follows that $\omega_{p}^{-1}\tau_{\leq 0}F[-2r(p)]=\tau_{\leq 0}\omega_{p}^{-1}F[-2r(p)]=0$. A contradiction. The conclusion now follows from Lemma 7.2.
Now suppose given $F^{\prime},F^{\prime\prime}\in sh_{\Lambda}(M)^{\heartsuit}$ and some extension $0\to F^{\prime}\to F\to F^{\prime\prime}\to 0$. The microstalk functors are $t$-exact by construction so we get a corresponding extension of microstalks inside $C^{\heartsuit}$, which is closed under extensions. All microstalks of $F$ lie in $C^{\heartsuit}$, so $F\in sh_{\Lambda}(M)^{\heartsuit}$. The same argument also shows that $sh_{\Lambda}(M)^{\heartsuit}$ is closed under quotients and subobjects. ∎
We now review the results of Waschkies [27]. Consider $T^{\circ}M=T^{*}M\setminus M$, and the map $\pi:T^{\circ}M\to\mathbb{P}^{*}M$.
Lemma 7.9.
The natural map
(19) $\mu sh^{pre}(\pi^{-1}(m))\to(\mathbb{P}\mu sh)_{m}$
is an equivalence.
Proof.
We first check fully-faithfulness. It is equivalent to check that the natural map
$\mu sh^{pre}(\pi^{-1}(m))\to(\pi_{*}\mu sh)_{m}$
is fully-faithful. Given objects $K,L$ in the source, Waschkies [27, Sec. 2.4] shows that $\hom(K,L)$ in the source is computed by $(\pi_{*}\mu\operatorname{Hom}(K,L))_{m}$, which also computes $\hom(K,L)$ in the target.
We now check essential surjectivity. If $K\in(\mathbb{P}\mu sh)_{m}$, then it is represented by an object $\tilde{K}\in\mathbb{P}\mu sh(U),p\in U,$ which is representable by a sheaf at $p$. This means that (after possibly shrinking $U$) there is a complex-contact embedding $\chi:U\to V\subset\mathbb{P}^{*}N$ so that $\chi_{*}(K)=\chi_{*}(\tilde{K})\in(\pi_{*}\mu sh)_{\chi(m)}$ is in the image of $(\pi_{*}\mu sh^{pre})_{\chi(m)}\to(\pi_{*}\mu sh)_{\chi(m)}$. By standard quantization of sheaf kernels (see e.g. [27, Thm. 4.2.1]) we have
[…]
where the vertical arrows are (19). The claim follows. ∎
Theorem 7.10.
[27, Thm. 5.1.5] Assume $M$ is complex analytic, $\Lambda\subset T^{\circ}M$ is a $\mathbb{C}^{*}$ invariant (singular) Lagrangian, and $\Lambda/\mathbb{C}^{*}\to M$ is a finite map. Then the natural map of sheaves of categories
$sh_{\Lambda}/loc\to\mathbb{P}\mu sh_{\Lambda}$
is an equivalence.
Proof.
We check at stalks. There are natural maps:
$(sh_{\Lambda}/loc)_{m}=sh_{\Lambda,m}/loc_{m}\to\mu sh^{pre}(\pi^{-1}(\tilde{m}))\to(\mathbb{P}\mu sh_{\Lambda})_{\tilde{m}}$
where $m\in M$ and $\tilde{m}\in\Lambda$ its (possibly disconnected) preimage. In [27, Thm. 5.1.5], Waschkies constructs the inverse using the refined microlocal cutoff of D’Agnolo [6], and shows that it is fully faithful essentially by arguing that all categories in question have homs given by the $\mu hom$ functor, similarly to [15, Thm. 6.1.2]. ∎
Remark 7.11.
Note that one must sheafify quotients, so $(sh_{\Lambda}/loc)(U)$ is generally not $sh_{\Lambda}(U)/loc(U)$. However, both $sh_{\Lambda}$ and $loc$ are constructible, so for small enough open neighborhood $U$ of any $m\in M$, we have $(sh_{\Lambda}/loc)(U)=(sh_{\Lambda}/loc)_{m}=(sh_{\Lambda})_{m}/loc_{m}=sh_{\Lambda}(U)/loc(U)$.
Remark 7.12.
The analogue of Theorem 7.10 with complex analyticity hypotheses removed and with $\mathbb{C}^{*}$ replaced by $\mathbb{R}^{+}$ is also true (indeed easier: the cutoffs already in [15] are good enough and one does not need [6]); see [22, Lem 6.7, Prop. 6.9].
Corollary 7.13.
In the situation of Theorem 7.10, there is a $t$-structure on $\pi_{*}\mu sh_{\Lambda}$ characterized by the formulas in Definition 7.7.
Proof.
For any $U\subset M$, the $t$-structure descends to $sh_{\Lambda}(U)/loc(U)$ by Lemma 7.2 and Lemma 7.8. For $p\in T^{\circ}M$, the functor $\omega_{p}$ descends to $sh_{\Lambda}(U)/loc(U)$, and clearly the quotient $t$-structure on $sh_{\Lambda}(U)/loc(U)$ is characterized exactly as in Definition 7.7. ∎
Remark 7.14.
In [27, Thm. 6.1.5], this is extended to arbitrary $\Lambda\subset\mathbb{P}\mathbb{T}^{*}M$ by noting that complex contact transformation preserves the conditions of Definition 7.7 above, using the degree shift formula [15, prop. 7.4.6]. In our discussion below, we don’t explicitly use this complex contact invariance of Definition 7.7 to construct the microlocal perverse $t$-structure, but would need it to check that on cotangent bundles, we have defined the usual perverse $t$-structure.
7.3. Perverse microsheaves on complex contact and symplectic manifolds
Definition 7.15.
Fix a $t$-structure on the coefficient category $C$, e.g. the standard $t$-structure on the category of $R$-modules, for a commutative ring $R$. Let $V$ be a contact manifold of complex dimension $12n-$ 1. Fix any $(C,C^{\leq 0})$ orientation data $o$. (If $C=R-mod$, we may always choose the canonical orientation data.)
We define subcategories of $\mu sh_{\tilde{V},\mathbb{C}-c,o}$ by constraining the microstalks to respect the $t$-structure of $C$. In the notation of Definition 5.11:
$(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{\leq 0}:=(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{C^{\leq 0}[2n]}$
$(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{\geq 0}:=(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{C^{\geq 0}[2n]}$
We define similarly the corresponding notions for objects supported in some fixed (singular) Lagrangian, and define as always the corresponding notions on $V$ by pushforward.
The $[2n]$ is just a convention, set to match the usual conventions for perverse sheaves.
Theorem 7.16.
$(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\leq 0}$, $(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\geq 0}$ determine a $t$-structure on $\mathbb{P}\mu sh_{V,\mathbb{C}-c,o}$.
Proof.
Being a $t$-structure is an objectwise condition. So it suffices to show, for every fixed singular Lagrangian $\Lambda$, that objects supported in $\Lambda$ satisfy the corresponding conditions; it suffices therefore to show $(\mathbb{P}\mu sh_{\Lambda,o})^{\leq 0}$, $(\mathbb{P}\mu sh_{\Lambda,o})^{\geq 0}$ define a $t$-structure on $\mathbb{P}\mu sh_{\Lambda,o}$. We may check locally. Thus we may assume $\Lambda$ is an arbitrarily small and contractible neighborhood of some point $p\in\Lambda$.
By the Kashiwara–Kawai’s general position theorem (see [13, Sec. 1.6]), we may choose a complex manifold $M$ and an embedding $\iota:\Lambda\to\mathbb{P}T^{*}M$ such that $\Lambda\to M$ is finite, hence an embedding at generic points of $\Lambda$. Let $\eta$ be the Maslov datum used on $V$; recall that the grading part is canonical. Writing $\phi$ for the fiber Maslov datum on $T^{*}M$, we use the canonical identification of the ‘grading’ parts of $\eta$ and $\iota^{*}\phi$, and any arbitrary homotopy to identify the orientation data parts (which certainly exists since $\Lambda$ is contractible). Thus we have an identification $\iota^{*}\mu sh_{\iota({\Lambda}),o}\cong\mu sh_{\Lambda,o}$. Per 5.10 and Theorem 7.10, this identification intertwines the microstalk functors used to characterize the perverse $t$-structure on the left hand side with the functors we used in Definition 5.11, which therefore defines a $t$-structure as well. ∎
We immediately deduce Theorem 1.4 from the introduction.
Proof of Theorem 1.4.
As noted in Remark 6.4, $\mathbb{C}^{*}$ acts on $W\times\mathbb{C}$ by contactomorphism fixing $W\times\{0\}$ set-wise. Hence the inclusion $({\gamma_{\mathbb{C}}})_{*}\mathbb{P}\mu sh_{W\times\{0\}}\hookrightarrow({\gamma_{\mathbb{C}}})_{*}\pi_{*}\mu sh_{\widetilde{W\times\mathbb{C}}}$ is stable under this action. Taking $\mathbb{C}^{*}$-invariants, we get a fully faithful inclusion $\mu_{\mathbb{C}}sh_{W}=(({\gamma_{\mathbb{C}}})_{*}\mathbb{P}\mu sh_{W\times\{0\}})^{\mathbb{C}^{*}}\hookrightarrow(({\gamma_{\mathbb{C}}})_{*}\pi_{*}\mu sh_{\widetilde{W\times\mathbb{C}}})^{\mathbb{C}^{*}}=({\gamma_{\mathbb{C}}})_{*}\mu sh_{W}$. The $\mathbb{C}^{*}$-action manifestly preserves the subspaces $\mu sh_{W,\mathbb{C}-c}^{\leq 0}$ and $\mu sh_{W,\mathbb{C}-c}^{\geq 0}$, so the $t$-structure passes to the invariants (taking hearts is a pullback ($\mathcal{F}^{\heartsuit}:=\mathcal{F}^{\leq 0}\cap\mathcal{F}^{\geq 0}$) hence a limit, and hence commutes with taking $G$-invariants which is also a limit). ∎
As a consequence of Theorem 7.16 and Lemma 7.6, we conclude:
Corollary 7.17.
$(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\heartsuit}$ is a stack of abelian categories; in particular, the global sections $(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\heartsuit}(V)$ is an abelian category.
Remark 7.18.
A $t$-structure is said to be nondegenerate if $\bigcap C^{\leq 0}=0=\bigcap C^{\geq 0}$. By co-isotropicity of microsupport, the vanishing of all microstalks implies the vanishing of an object. We conclude that if the $t$-structure on $C$ is non-degenerate, then the $t$-structure on $\mathbb{P}\mu sh_{V,\mathbb{C}-c,o}$ is also non-degenerate.
Remark 7.19.
Choose a stable (not necessarily presentable) subcategory $D\subset C$ to which the $t$-structure restricts. Require $(C,C^{\leq 0},D)$-orientation data. Then it is evident from the definitions that the truncation functors preserve, hence define a $t$-structure on, the subcategory of objects with all microstalks in $D$, characterized by the same formulas save only with e.g. $C^{\leq 0}$ replaced by $C^{\leq 0}\cap D$. For instance, we can take various bounded categories e.g. $D=C^{+}=\bigcup C^{\geq n}$, or $D=C^{-}=\bigcap C^{\leq n}$, or $D=C^{b}=C^{+}\cap C^{-}$, or ask the microstalks to be compact objects $D=C^{c}$.
Remark 7.20.
Fix $D\subset C$ as above, and assume $D^{\heartsuit}$ is Artinian. (E.g., $C=k-mod$ for a field $k$ and $D=C^{c}$.) Then the full subcategory of $(\mathbb{P}\mu sh_{V,o})^{D,\heartsuit}$ on objects with finitely stratified (rather than just locally finite) support is also Artinian. Indeed, any descending chain must have eventually stabilizing microsupports; we may restrict attention to one each in the finitely many connected components of the smooth locus of the support, hence by some point, all have stabilized.
Remark 7.21.
Let $V$ be a complex contact manifold and let $L\subset V$ be a complex Legendrian. Then $\mathbb{P}\mu sh_{V,L}(-)$ is a sheaf of stable categories while $\mathbb{P}erv_{V,L}(-)$ is a sheaf of abelian categories. Be warned however that $D(\mathbb{P}erv_{V,L}(-))$ is only a presheaf of stable categories. In particular, the natural map $D(\mathbb{P}erv_{V,L}(-))\to\mathbb{P}\mu sh_{V,L}(-)$ may restrict to an equivalence on stalks without being an equivalence on global sections. A very special case: let $V=T^{*}S^{2}\times\mathbb{\mathbb{C}}$ and $L=0_{S^{2}}\times\{0\}$. Then $\mathbb{P}\mu sh_{V,L}(L)=loc(S^{2})\otimes loc(\mathbb{C}^{*})$ while $\mathbb{P}erv_{V,L}(L)=vect_{\mathbb{C}}\otimes loc(\mathbb{C}^{*})$, due to $S^{2}$ being simply connected. Similarly, $\mu_{\mathbb{C}}sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})=loc(S^{2})$ while $\mu_{\mathbb{C}}sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})^{\heartsuit}=vect_{\mathbb{C}}$.

### v3

@@
 Perverse sheaves are those $F$ with the following property:
-(1) $\dim_{\mathbb{C}}\{\,p\in M\,|\,H^{i}(p^{*}F)\neq 0\}\leq i\qquad\qquad\dim_{\mathbb{C}}\{\,p\in M\,|\,H^{i}(p^{!}F)\neq 0\}\leq i$
-Here $p^{*}$ and $p^{!}$ are the restriction functors to the point $p$.
+(1) $\dim_{\mathbb{C}}\{\,x\in M\,|\,H^{i}(\iota_{x}^{*}F)\neq 0\}\leq-i\qquad\qquad\dim_{\mathbb{C}}\{\,x\in M\,|\,H^{i}(\iota_{x}^{!}F)\neq 0\}\leq i$
+Here we denote by $\iota_{x}:\{x\}\hookrightarrow M$ the inclusion of the point $p$ to $M$ and by $\iota_{x}^{*}$ and $\iota_{x}^{!}$ the induced restriction functors.
 Perverse sheaves have played a pivotal role in many results in algebraic geometry and geometric representation theory.
-They turn to be natural both in terms of considerations of Frobenius eigenvalues in positive characteristic [3], and in terms of analytic considerations in characteristic zero, where they are the sheaves of solutions to regular holonomic differential equations, or more generally, D-modules [10, 21].
+They turn out to be natural both in terms of considerations of Frobenius eigenvalues in positive characteristic [3], and in terms of analytic considerations in characteristic zero, where they are the sheaves of solutions to regular holonomic differential equations, or more generally, D-modules [19, 23, 20, 34].
 This latter equivalence highlights a key feature, not immediately apparent from the definition: despite being a seemingly arbitrarily demarcated subcategory of a category of complexes, perverse sheaves form an abelian category.
@@
 More generally still, Kashiwara has constructed a sheaf of categories on any complex contact manifold, locally equivalent to the microlocalization of $D$-modules [21], see also [37].
-The variant appropriate to conic complex symplectic geometry has allowed the methods of geometric representation theory to be extended beyond cotangent bundles to more general symplectic resolutions and similar spaces [14, 4].
-The purpose of the present article is to construct categories of complex-constructible microsheaves on complex contact manifolds which naturally carry a perverse $t$-structure, globalizing the construction of Waschkies [27].
-In the sequel we establish a Riemann-Hilbert equivalence with the canonical stack of $\mathcal{E}$-modules defined by Kashiwara [11].
+A variant of this construction appropriate to conic complex symplectic geometry has allowed the methods of geometric representation theory to be extended beyond cotangent bundles to more general symplectic resolutions and similar spaces [24, 4].
+The purpose of the present article is to construct perverse $t$-structures on categories of complex-constructible microsheaves, globalizing the construction of Waschkies [42].
+In the sequel [29] we establish a Riemann-Hilbert equivalence with the canonical stack of $\mathcal{E}$-modules defined by Kashiwara [21].
 Our starting point is the globalization [39, 35] of the microlocal sheaf theory of Kashiwara and Schapira [25].
 We recall the relevant notions in Section 4.
-In brief, the theory takes as input a real contact or exact symplectic manifold $V$, a choice of symmetric monoidal stable presentable coefficient category $\mathcal{C}$, and certain topological ‘Maslov data’.
-Maslov data is by definition a choice of trivialization for a certain canonical obstruction $V\to B^{2}Pic(\mathcal{C})$, and in fact agrees with the choices made in defining Floer-theoretic invariants in the same target spaces.
-The output of the theory is a sheaf of stable categories $\mu sh_{V}$ on $V$.
+In brief, the theory takes as input a real contact or exact symplectic manifold $V$, a choice of symmetric monoidal stable compactly generated coefficient category $\mathcal{C}$, and a trivialization of a certain canonical obstruction $V\to B^{2}Pic(\mathcal{C})$.
+We refer to said trivialization as a Maslov datum; is is also what is required to define Floer-theoretic invariants in the same target spaces.
+The output of the theory is a sheaf of stable categories $\mu sh_{V}$ on $V$ [35, Thm.
+1.1].
 For a locally closed subset $X\subset V$, we write $\mu sh_{X}\subset\mu sh_{V}|_{X}$ for the subsheaf of full subcategories on objects locally supported in $X$.
-A Legendrian or conic Lagrangian $L\subset V$ determines an obstruction $L\to BPic(\mathcal{C})$, a choice of trivialization for which (‘secondary Maslov data’) yields an equivalence $\mu sh_{L}\cong Loc_{L}$ with the sheaf of categories of local systems along $L$.
-Here we show that in the complex case:
-Theorem 1.1 (see Section 3).
-Let $W$ be an exact complex symplectic manifold, and $\mathcal{C}$ the category of modules over a (discrete) commutative ring $R$.
-Then there is a canonical Maslov datum for $W$, with respect to which secondary Maslov data for a Lagrangian $L\subset W$ are identified with spin structures on $L$.
-From this and the existing theory [22], we deduce:
-Theorem 1.2 (see Section 6).
-Let $W$ be an exact complex symplectic manifold.
-For a discrete commutative ring $R$, there is a canonical sheaf of stable $R$-linear categories $\mu sh_{W}$ on $W$.
-For a complex Lagrangian $L\subset X$, a spin structure on $L$ determines an equivalence $\mu sh_{L}\cong Loc_{L}$.
-Taking $D$ to be a disk around a smooth Lagrangian point $x\in X$, we obtain a canonical (up to noncanonical automorphism) ‘microstalk’ functor $\mu_{x}:\mu sh_{X}(W)\to\mu sh_{X}(D)\xrightarrow{\sim}R-mod$.
-We write $\mu sh_{W,\mathbb{C}-c}\subset\mu sh_{W}$ for the sheaf of full subcategories on objects whose (micro)support is a complex analytic Lagrangian subset of $W$.
-We define $(\mu sh_{W,\mathbb{C}-c})^{\leq 0}\subset\mu sh_{W,\mathbb{C}-c}$ as the sheaves of full subcategories on those objects all of whose microstalks, as elements of $R-mod$, have cohomology concentrated in degrees $\leq 0$.
-We define similarly $(\mu sh_{W,\mathbb{C}-c})^{\geq 0}$.
+A Legendrian or conic Lagrangian $L\subset V$ determines an obstruction $L\to BPic(\mathcal{C})$, a choice of trivialization for which (a ‘secondary Maslov datum’) yields an equivalence
+$\mu sh_{L}\cong loc_{L}$
+with the sheaf of categories of local systems along $L$ [35, Thm.
+1.2].
+In particular, if $D\subset V$ is a smooth Legendrian disk containing a point $p$, a choice of secondary Maslov datum for $D$ determines an equivalence
+(2) $(\mu sh_{D})_{p}\xrightarrow{\sim}\mathcal{C}.$
+The space of secondary Maslov data for the disk $D$ is a torsor for $BPic(\mathcal{C})$, which acts on maps (2) in the evident way.
+In this paper, we will be interested in contact and conic symplectic manifolds which come from complex geometry.
+It turns out that such manifolds admit a canonical choice of Maslov datum, yielding a canonical notion of $\mu sh$.
+For simplicity, we state this in the symplectic setting.
+By an exact complex symplectic manifold, we mean a complex manifold $W$ along with a holomorphic $1$-form $\lambda$ such that $d\lambda$ is holomorphic symplectic.
+The underlying real manifold of $W$ carries the real exact symplectic structure $\operatorname{re}(\lambda)$, so that we can meaningfully discuss Maslov data and microlocal sheaves on $W$.
+Theorem 1.1.
+Let $W$ be an exact complex symplectic manifold and $\mathcal{C}$ the category of modules over a (discrete) commutative ring $R$.
+Then there is a canonical Maslov datum for $W$, with respect to which secondary Maslov data for a conic complex Lagrangian $L\subset W$ are identified with $R$-spin structures on $L$.
+By an $R$-spin structure on a real symplectic manifold $W$, we mean a null-homotopy of the composition $W\to BU\xrightarrow{w_{2}}B^{2}\mathbb{Z}^{\times}\to B^{2}R^{\times}$.
+When $R=\mathbb{Z}$, this is a spin structure in the ordinary sense; in general, if such structures exist, they form a torsor for $\mathrm{H}^{1}(X,R^{\times})$.
+The proof of Theorem 1.1 is in Section 4.6, where we also deduce, from the aforementioned general properties of microlocal sheaves:
+Corollary 1.2 (The canonical microsheaf category).
+Let $W,\mathcal{C}$ be as in Theorem 1.1.
+There is a canonical sheaf of stable $\mathcal{C}$-linear categories $\mu sh_{W}$ on $W$.
+For a complex conic Lagrangian $L\subset W$, an $R$-spin structure $\sigma$ on $L$ determines an equivalence $\mu sh_{L}\cong loc_{L}$.
+In particular, specializing to the neighborhood of a point gives:
+Corollary 1.3 (The microstalk functor).
+Let $W,\mathcal{C}$ be as in Theorem 1.1. If $X\subseteq W$ is a closed subset which locally around $p\in X$ is conic complex Lagrangian, there is a functor
+(3) $\omega_{p}^{-1}:(\mu sh_{X})_{p}\xrightarrow{\sim}\mathcal{C},$
+which is well-defined up to non-canonical invertible natural transformation. In particular, given $K\in\mu sh_{X}(W)$, the isomorphism class of $\omega_{p}^{-1}(K)$ is well defined.
+We refer to $\omega_{p}^{-1}(K)$ as the microstalk of $K$ at $p$. In Section 5.2, we will give a more explicit characterization of this microstalk.
+Definition 1.4.
+Write $\mu sh_{W,\mathbb{C}-c}\subset\mu sh_{W}$ for the sheaf of full subcategories on objects whose (micro)support is a complex analytic Lagrangian subset of $W$. We define
+(4) $(\mu sh_{W,\mathbb{C}-c})^{\leq 0}\subset\mu sh_{W,\mathbb{C}-c}\qquad\qquad(\mu sh_{W,\mathbb{C}-c})^{\geq 0}\subset\mu sh_{W,\mathbb{C}-c}$
+as the sheaves of full subcategories on those objects all of whose microstalks, as elements of RmodR-mod, have cohomology concentrated in degrees $\leq 0$, resp. $\geq 0$.
 We do not know in general whether $((\mu sh_{W,\mathbb{C}-c})^{\leq 0},(\mu sh_{W,\mathbb{C}-c})^{\geq 0})$ is a $t$-structure on $\mu sh_{W,\mathbb{C}-c}$.
 However, we will use these subcategories to construct $t$-structures on certain related categories.
-To introduce these other categories, we return to the basic definitions of microsheaves. For $M$ a manifold, we have the presheaf of categories $\mu sh^{pre}$ on $T^{*}M$ defined by
-$\mu sh^{pre}(U):=sh(M)/\{F\,|\,ss(F)\cap U=\emptyset\}.$
-We write $\mu sh$ for its sheafification.
-Let us write $T^{*}M_{\mathbb{R}}$ for the topological space whose open sets are the $\mathbb{R}^{>0}$-invariant open subsets of $T^{*}M$, and $\gamma_{\mathbb{R}}:T^{*}M\to T^{*}M_{\mathbb{R}}$ for the continous map given by the set theoretic identity.
-Let us write $\mu_{\mathbb{R}}sh$ for the sheafification of $(\gamma_{\mathbb{R}})_{*}\mu sh^{pre}$.
-Because microsupports are conic, there is an isomorphism of the sheaf pullback $\gamma_{\mathbb{R}}^{*}\mu_{\mathbb{R}}sh\xrightarrow{\sim}\mu sh$.
-In fact, because $\mathbb{R}^{>0}$ is contractible, the presheaf preimage $\gamma_{\mathbb{R}}^{*,pre}\mu_{\mathbb{R}}sh$ is already a sheaf, so already isomorphic to $\mu sh$.
-That is, nothing is gained or lost by considering $\mu_{\mathbb{R}}sh$ instead of $\mu sh$.
-By contrast, when $M$ is a complex manifold, consider similarly the topological space $T^{*}M_{[\mathbb{C}^{*}]}$ on $\mathbb{C}^{*}$-invariant open sets in $T^{*}M$, and define $\mu_{\mathbb{C}}sh$ etc.
-similarly.
-Now there are maps
-$\gamma_{\mathbb{C}}^{*,pre}\mu_{\mathbb{C}}sh\to\gamma_{\mathbb{C}}^{*}\mu_{\mathbb{C}}sh\to\mu sh$
-These maps are all fully faithful on sections [27, Sec. 2.4], but we do not know when or whether (or after imposing what microsupport conditions) they are essentially surjective.
-Consider the projectivized cotangent bundle, $\pi:T^{*}M\setminus 0_{M}\to\mathbb{P}^{*}M$. Then we have an inclusion of sheaves of full subcategories:
-(2) $\mathbb{P}\mu sh:=\pi_{*}\gamma_{\mathbb{C}^{*}}^{*,pre}\mu_{\mathbb{C}}sh\hookrightarrow\pi_{*}\mu sh$
-More generally, for $V$ a complex contact manifold and $\pi:\tilde{V}\to V$ its symplectization, we will identify a sheaf of full subcategories $\mathbb{P}\mu sh_{V}\subset\pi_{*}\mu sh_{\tilde{V}}$ which agrees with (2) in local charts (see Definition 5.14).
-We write $\mathbb{P}\mu sh_{V,\mathbb{C}-c}\subset\mathbb{P}\mu sh_{V}$ for the subsheaf of full subcategories on objects with complex Legendrian microsupport; note $\mathbb{P}\mu sh_{V,\mathbb{C}-c}=\mathbb{P}\mu sh_{V}\cap\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}$.
-For complex exact symplectic manifolds $(W,\lambda)$, we may use the contactization $(W\times\mathbb{C},\lambda+dz)$ to define a sheaf of categories $\mathbb{P}\mu sh_{W\times 0}$.
+Recall that for a complex contact manifold $V$, its symplectization is the $\mathbb{C}^{*}$ bundle $\pi:\widetilde{V}\to V$ whose (local) holomorphic sections give (local) complex contact forms; it carries canonically an exact complex symplectic form (as we review in more detail in Section 2) and hence we have the canonically defined $\mu sh_{\widetilde{V}}$.
+Theorem 1.5 (6.17).
+Then the pair $\left((\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c})^{\leq 0},(\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c})^{\geq 0}\right)$ determines a $t$-structure on $\pi_{*}\mu sh_{\widetilde{V},\mathbb{C}-c}$ (i.e., the sections of the above sheaves over any open set $U$ give a t-structure on $\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}(U)$).
+For an exact complex symplectic manifold $(W,\lambda)$, we may use the contactization $(W\times\mathbb{C},\lambda+dz)$ to define a sheaf of categories $\pi_{*}\mu sh_{\pi^{-1}(W\times\{0\})}$, i.e.
+the sheaf of full subcategories of $\pi_{*}\mu sh_{\widetilde{W\times\mathbb{C}}}$ on objects whose support is contained in $\pi^{-1}(W\times\{0\})$.
 We do not understand the relationship of this with $\mu sh_{W}$ in general.
 However, if the Liouville flow on $(W,\lambda)$ integrates to a weight-1 $\mathbb{C}^{*}$-action, then $\mathbb{C}^{*}$ naturally acts on $W\times\mathbb{C}$ by contactomorphism.
-Let $\gamma_{\mathbb{C}}$ be the set-theoretic identity on $W\times\mathbb{C}^{*}$ (resp.
+Let $\gamma_{\mathbb{C}}$ be the set-theoretic identity on $W\times\mathbb{C}$ (resp.
 on $W$) where the source carries the Euclidean topology but the target is endowed with the $\mathbb{C}^{*}$ invariant topology.
-In Theorem 6.3, we show that there is a natural $\mathbb{Z}[t,t^{-1}]$-linear structure on $\lambda_{*}\mathbb{P}\mu sh_{W\times\{0\}}$, and there are fully faithful embeddings
-(3) $(\gamma_{\mathbb{C}})_{*}\mathbb{P}\mu sh_{W\times\{0\}}\hookrightarrow(\gamma_{\mathbb{C}})_{*}\mu sh_{W}\otimes\mathbb{Z}[t,t^{-1}]$
-(4) $\mu_{\mathbb{C}}sh_{W}:=(\gamma_{\mathbb{C}})_{*}\mathbb{P}\mu sh_{W\times\{0\}}\otimes_{\mathbb{Z}[t,t^{-1}]}\mathbb{Z}\hookrightarrow(\gamma_{\mathbb{C}})_{*}\mu sh_{W}$
-We can now state our results on $t$-structures.
-Theorem 1.3 (7.16).
-Let $V$ be a complex contact manifold. Then the pair
-$\mathbb{P}\mu sh_{V,\mathbb{C}-c}^{\geq 0}:=\mathbb{P}\mu sh_{V}\cap\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}^{\geq 0}\qquad\qquad\qquad\mathbb{P}\mu sh_{V,\mathbb{C}-c}^{\leq 0}:=\mathbb{P}\mu sh_{V}\cap\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}^{\leq 0}$
-determines a $t$-structure on $\mathbb{P}\mu sh_{V,\mathbb{C}-c}$.
 Theorem 1.6.
-Let $W$ be a complex exact symplectic manifold whose Liouville vector field integrates to a weight-1 $\mathbb{C}^{*}$ action. Then the pair
-$\mu_{\mathbb{C}}sh_{W,\mathbb{C}-c}^{\geq 0}:=\mu_{\mathbb{C}}sh_{W}\cap\mu sh_{W,\mathbb{C}-c}^{\geq 0}\qquad\qquad\qquad\mu_{\mathbb{C}}sh_{W,\mathbb{C}-c}^{\leq 0}:=\mu_{\mathbb{C}}sh_{W}\cap\mu sh_{W,\mathbb{C}-c}^{\leq 0}$
-determines a $t$-structure on $\mu_{\mathbb{C}}sh_{W,\mathbb{C}-c}$. Moreover, the Hom sheaf of two objects in the heart is a ($\frac{1}{2}\dim W$-shifted) perverse sheaf.
-Example 1.5.
-Take $(W,\lambda)$ as in Theorem 1.4, and suppose in addition $(W,\operatorname{re}\lambda)$ is a Weinstein manifold admiting a complete collection of generalized cocores, so that the main theorem of [7] can be applied.
-Many examples coming from geometric representation theory – quiver varieties, hypertoric varieties, moduli of Higgs bundles, etc.
-satisfy this assumption; cf.
-[29].
-Let $L$ be the core of $W$.
-Then $\mu sh_{L}(W)$ is equivalent to (the module category of) $Fuk(W)$ [7], and so Theorem 1.4 asserts we have constructed a $t$-structure on the full subcategory
-$\mu_{\mathbb{C}}sh_{L}(W)\hookrightarrow\mu sh_{L}(W)\cong Fuk(W)$
-and correspondingly the perverse microsheaves determine a full abelian subcategory of $H^{0}Fuk(W)$.
-Let us conclude by explicitly advertising the question:
-Definition 1.6.
 Let $W$ be a complex exact symplectic manifold whose Liouville vector field integrates to a weight-1 $\mathbb{C}^{*}$ action.
-We say a complex conic Lagrangian $L\subset W$ has only complex global microsheaves if $\mu_{\mathbb{C}}sh_{L}(W)\hookrightarrow\mu sh_{L}(W)$ is essentially surjective.
-Question 1.7.
-Which conic complex Lagrangians have only complex global microsheaves?
-We do not know that the answer is not “all of them”.
-(In fact, in a previous draft of this article, we blithely assumed this was the case.) Example 1.5 provides one motivation for the question, since in the notation there, whenever $L$ has only complex global microsheaves, we have constructed a $t$-structure on the Fukaya category of $W$.
-Here is what little we know about 1.7.
-It is not difficult to see that $\mathbb{C}^{*}$-conic subsets of cotangent bundles are completely complex, and slightly more generally, that the same holds for Lagrangians which admit a cover by $\mathbb{C}^{*}$-conic subsets of cotangent bundles.
-(This implies the existence of $t$-structures on Fukaya categories of hypertoric varieties, ALE spaces, and a few other quiver varieties.) We do not presently know much more in general.
+Then the pair $((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c})^{\geq 0},((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c})^{\leq 0}$ determines a $t$-structure on $(\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c}$.
+Moreover, the Hom sheaf of two objects in the heart is a ($\frac{1}{2}\dim W$-shifted) perverse sheaf.
+Finally, we recall that there is a comparison theorem [11] between microsheaves and Fukaya categories; consequently, our results can be translated to give $t$-structures on certain Fukaya categories.
+We spell this out in Appendix B.
 Acknowledgements.
-We wish to thank Roman Bezrukavnikov, Sanath Devalapurkar, Peter Haine, Benjamin Gammage, Justin Hilburn, Michael McBreen, Semon Rezchikov, Marco Robalo, Pierre Schapira and Filip Živanović for enlightening conversations.
+We wish to thank Roman Bezrukavnikov, Sanath Devalapurkar, Peter Haine, Benjamin Gammage, Sam Gunningham, Justin Hilburn, Wenyuan Li, Michael McBreen, Semon Rezchikov, Marco Robalo, Pierre Schapira and Filip Živanović for enlightening conversations.
+We are particularly grateful to Sanath Devalapurkar for help with Section 3 and to Wenyuan Li for help with Proposition 5.9.
 LC was partially supported by Simons Foundation grant 385573 (Simons Collaboration on Homological Mirror Symmetry).
@@
 CK and VS were partially supported by: VILLUM FONDEN grant VILLUM Investigator 37814, and VS was supported in addition by Novo Nordisk Foundation grant NNF20OC0066298, Danish National Research Foundation grant DNRF157, and the USA NSF grant CAREER DMS-1654545.
+CK was supported by Max Planck Institude for Mathematics in Bonn during the revision of the paper.
 2. Complex contact and symplectic manifolds
@@
 Definition 2.1.
-A complex (or holomorphic) symplectic manifold is a complex manifold $X$ along with a closed holomorphic $2$-form $\omega\in H^{0}(X,\Omega^{2}_{X})$.
-A complex symplectic manifold $(X,\omega)$ determines a family, parameterized by $\hbar\in\mathbb{C}^{*}$, of real symplectic manifolds $(X,\operatorname{re}(\hbar\omega))$.
-By an exact complex symplectic manifold, we mean a pair $(X,\lambda)$ where $X$ is a complex manifold and $\lambda$ is a holomorphic 1-form such that $d\lambda=\partial\lambda$ is symplectic.
+A complex (or holomorphic) symplectic manifold is a complex manifold $W$ along with a closed holomorphic $2$-form $\omega\in H^{0}(W,\Omega^{2}_{W})$.
+A complex symplectic manifold $(W,\omega)$ determines a family, parameterized by $\hbar\in\mathbb{C}^{*}$, of real symplectic manifolds $(W,\operatorname{re}(\hbar\omega))$.
+By an exact complex symplectic manifold, we mean a pair $(W,\lambda)$ where $W$ is a complex manifold and $\lambda$ is a holomorphic 1-form such that $d\lambda=\partial\lambda$ is symplectic.
 Example 2.2.
-Let $X$ be a complex manifold.
-Then the holomorphic cotangent bundle $\Omega_{X}$ carries a canonical holomorphic $1$-form $\lambda_{can,\mathbb{C}}=ydx$, where $(x_{1},\dots,x_{n},y_{1},\dots,y_{n})$ are canonical holomorphic coordinates.
-Then $(\Omega_{X},d\lambda_{can,\mathbb{C}})$ is a complex symplectic manifold.
-Meanwhile, the real cotangent bundle $T^{*}M$ carries the canonical $1$-form $\lambda_{can}$. There is a natural identification $\Omega_{X}\to T^{*}M$ defined in local holomorphic coordinates by
+Let $M$ be a complex manifold.
+Then the holomorphic cotangent bundle $\Omega_{M}$ carries a canonical holomorphic $1$-form $\lambda_{can,\mathbb{C}}=ydx$, where $(x_{1},\dots,x_{n},y_{1},\dots,y_{n})$ are canonical holomorphic coordinates.
+Then $(\Omega_{M},d\lambda_{can,\mathbb{C}})$ is a complex symplectic manifold.
+Meanwhile, the real cotangent bundle $T^{*}M$ carries the canonical $1$-form $\lambda_{can}$. There is a natural identification $\Omega_{M}\to T^{*}M$ defined in local holomorphic coordinates by
 $(x,y)\mapsto(\operatorname{re}(x),\operatorname{im}(x),\operatorname{re}(y),-\operatorname{im}(y)).$
@@
 Definition 2.3.
-A complex (or holomorphic) contact manifold is a complex manifold $V$ of complex dimension $12n+$ 1 along with a holomorphic hyperplane field $\mathcal{H}\hookrightarrow\mathcal{T}_{V}$ which is maximally non-integrable.
+A complex (or holomorphic) contact manifold is a complex manifold $V$ of complex dimension $2n+1$ along with a holomorphic hyperplane field $\mathcal{H}\hookrightarrow\mathcal{T}_{V}$ which is maximally non-integrable.
 Concretely, this means that if $\alpha\in\Omega_{V}(U)$ is a holomorphic $1$-form for which $\mathcal{H}=\ker\alpha$ in some local chart $U\subset V$, then $\alpha\wedge(d\alpha)^{n}\neq 0$.
 Given a complex contact manifold $(V,\xi)$, there is a holomorphic line bundle $\mathcal{T}_{V}/\mathcal{H}\to V$.
-Contact forms are naturally sections of $(\mathcal{T}_{V}/\mathcal{H})^{\vee}$.
-Note the existence of holomorphic global sections is a rather more stringent condition than the corresponding condition in real smooth geometry; correspondingly complex contact manifolds typically do not have global contact forms.
+(Local) complex contact forms are nonvanishing (local) nonvanishing holomorphic sections of $(\mathcal{T}_{V}/\mathcal{H})^{\vee}$.
+Correspondingly, a global complex contact form is a global holomorphic trivialization of this line bundle (which does not typically exist).
 The bundle $(\mathcal{T}_{V}/\mathcal{H})^{\vee}$ is naturally a holomorphic sub-bundle of $\Omega_{V}$ (indeed, for $v\in V$, we have $(\mathcal{T}_{V}/\mathcal{H})^{\vee}_{v}=\{\alpha\in\Omega_{V,v}\mid\alpha(\xi)=0\}\subset\Omega_{V,v}$).
 We consider the $\mathbb{C}^{*}$-bundle
-$\pi:\tilde{V}:=(\mathcal{T}_{V}/\mathcal{H})^{\vee}\setminus 0_{V}\to V$
-which we call the complex symplectization of $V$ and let $\lambda_{\tilde{V}}$ denote the pullback of $\lambda_{can,\mathbb{C}}$ under the inclusion $\tilde{V}\hookrightarrow\Omega_{V}$.
-We also consider the projectivized $S^{1}$-bundle $p:\tilde{V}/\mathbb{R}_{+}\to V$ and set $\xi_{\hbar}:=\ker(\operatorname{re}(\hbar\lambda_{\tilde{V}}))$ for $\hbar\in\mathbb{C}^{*}$.
+$\pi:\widetilde{V}:=(\mathcal{T}_{V}/\mathcal{H})^{\vee}\setminus 0_{V}\to V$
+which we call the complex symplectization of $V$ and let $\lambda_{\widetilde{V}}$ denote the pullback of $\lambda_{can,\mathbb{C}}$ under the inclusion $\widetilde{V}\hookrightarrow\Omega_{V}$.
+We also consider the projectivized $S^{1}$-bundle $p:\widetilde{V}/\mathbb{R}_{+}\to V$ and set $)\xi_{\hbar}:=\ker(\operatorname{re}(\hbar\lambda_{\widetilde{V}}$)) for $\hbar\in\mathbb{C}^{*}$.
 The relation between these spaces is summarized by the following diagram:
@@
   1. (i)
-    $(\tilde{V},\hbar\lambda_{\tilde{V}})$ is an exact complex symplectic manifold
+[…]
   2. (ii)
-    $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$ is a real contact manifold
+    $(\widetilde{V}/\mathbb{R}_{+},\xi_{\hbar})$ is a real contact manifold
   3. (iii)
-    $(\tilde{V},\operatorname{re}(\hbar\lambda_{\tilde{V}}))$ is canonically isomorphic to the real symplectization of $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$.
-(This isomorphism intertwines the $1$-forms and the $\mathbb{R}_{+}$-bundle structure over $\tilde{V}/\mathbb{R}_{+}$).
+    $)(\widetilde{V},\operatorname{re}(\hbar\lambda_{\widetilde{V}}$)) is canonically isomorphic to the real symplectization of $(\widetilde{V}/\mathbb{R}_{+},\xi_{\hbar})$.
+(This isomorphism intertwines the $1$-forms and the $\mathbb{R}_{+}$-bundle structure over $\widetilde{V}$$\mathbb{R}_{+}$).
 Proof.
-We compute $\hbar\lambda_{\tilde{V}}$ locally by choosing a holomorphic contact $1$-form $\alpha$ defined on some open set $U\subset V$. Such a choice induces a holomorphic embedding
-$\iota_{\alpha}:\mathbb{C}^{*}\times U$ $\hookrightarrow\tilde{V}\subset\Omega_{V}$
+We compute $\hbar\lambda_{\widetilde{V}}$ locally by choosing a holomorphic contact $1$-form $\alpha$ defined on some open set $U\subset V$. Such a choice induces a holomorphic embedding
+$\iota_{\alpha}:\mathbb{C}^{*}\times U$ $\hookrightarrow\widetilde{V}\subset\Omega_{V}$
 $(z,x)$ $\mapsto z\alpha_{x}$
-To compute the pullback of $\hbar\lambda_{\tilde{V}}$ under $\iota_{\alpha}$, choose $(z,x)\in\mathbb{C}^{*}\times U$ and $d\iota_{\alpha}(v)\in T\tilde{V}_{z\alpha_{x}}\subset T\Omega_{V}$.
-Then we have $\hbar\lambda_{\tilde{V}}(d\iota_{\alpha}(v))=\hbar\lambda_{can,\mathbb{C}}(d\iota_{\alpha}(v))=z\hbar\alpha_{x}(d\pi\circ d\iota_{\alpha}(v))$.
+To compute the pullback of $\hbar\lambda_{\widetilde{V}}$ under $\iota_{\alpha}$, choose $(z,x)\in\mathbb{C}^{*}\times U$ and $d\iota_{\alpha}(v)\in T\widetilde{V}_{z\alpha_{x}}\subset T\Omega_{V}$.
+Then we have $\hbar\lambda_{\widetilde{V}}(d\iota_{\alpha}(v))=\hbar\lambda_{can,\mathbb{C}}(d\iota_{\alpha}(v))=z\hbar\alpha_{x}(d\pi\circ d\iota_{\alpha}(v))$.
 Hence
-(6) $\iota_{\alpha}^{*}\hbar\lambda_{\tilde{V}}=z\hbar\alpha.$
-Both (i) and (ii) can be checked immediately using (6). The proof of (iii) is an exercise in chasing definitions so is omitted.
+(6) $\iota_{\alpha}^{*}\hbar\lambda_{\widetilde{V}}=z\hbar\alpha.$
+Both (i) and (ii) can be checked immediately using (6). The proof of (iii) is an exercise in chasing definitions. ∎
 Example 2.5 (Example 2.2 continued).
-The complex projectivization $V:=(\Omega_{X}-0_{X})/\mathbb{C}^{*}$ is a complex contact manifold with respect to $\lambda_{can,\mathbb{C}}$ and we have $\tilde{V}=(\Omega_{X}-0_{X})$.
-The associated real projectivization $\tilde{V}/\mathbb{R}_{+}$ carries a circle of real contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.
+The complex projectivization $V:=(\Omega_{X}-0_{X})/\mathbb{C}^{*}$ is a complex contact manifold with respect to $\lambda_{can,\mathbb{C}}$ and we have $\widetilde{V}=(\Omega_{X}-0_{X})$.
+The associated real projectivization $\widetilde{V}$$\mathbb{R}_{+}$ carries a circle of real contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.
 Example 2.6.
@@
 Similarly, $\operatorname{re}(\hbar(\lambda+dz))$ defines a section of $\widetilde{X\times\mathbb{C}}/\mathbb{R}^{+}=X\times\mathbb{C}\times S^{1}$.
-Observe that there is a fiber-preserving $\mathbb{C}^{*}$-action on $\tilde{V}$: over some fiber $\tilde{V}_{v}\subset\Omega_{V,v}$, it sends $\alpha\mapsto z\alpha$ for $z\in\mathbb{C}^{*}$ and $v\in V$.
-For concreteness, we write $z=e^{t+i\theta}$ for $(t,\theta)\in\mathbb{R}_{+}\times S^{1}$ and let $\partial_{t},\partial_{\theta}$ denote the vector fields on $\tilde{V}$ generated by the $\mathbb{R}_{+}$ and $S^{1}$ actions.
-We have the following morphisms of bundles over $\tilde{V}$:
-(7) $\pi^{*}\xi\equiv\operatorname{ker}(\lambda_{\tilde{V}})/\langle\partial_{t},\partial_{\theta}\rangle\twoheadleftarrow\operatorname{ker}(\lambda_{\tilde{V}})\hookrightarrow\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\tilde{V}}))\hookrightarrow T\tilde{V}$
+Observe that there is a fiber-preserving $\mathbb{C}^{*}$-action on $\widetilde{V}$: over some fiber $\widetilde{V}_{v}\subset\Omega_{V,v}$, it sends $\alpha\mapsto z\alpha$ for $z\in\mathbb{C}^{*}$ and $v\in V$.
+For concreteness, we write $z=e^{t+i\theta}$ for $(t,\theta)\in\mathbb{R}_{+}\times S^{1}$ and let $\partial_{t},\partial_{\theta}$ denote the vector fields on $\widetilde{V}$ generated by the $\mathbb{R}_{+}$ and $S^{1}$ actions.
+We have the following morphisms of bundles over $\widetilde{V}$:
+(7) $\pi^{*}\xi\equiv\operatorname{ker}(\lambda_{\widetilde{V}})/\langle\partial_{t},\partial_{\theta}\rangle\twoheadleftarrow\operatorname{ker}(\lambda_{\widetilde{V}})\hookrightarrow\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\widetilde{V}}))\hookrightarrow T\widetilde{V}$
 Lemma 2.7.
-There is a splitting of vector bundles over $\tilde{V}$ (well-defined up to contractible choice)
-$\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\tilde{V}}))=\partial_{t}\oplus\langle\partial_{\theta},X\rangle\oplus\pi^{*}\xi,$
-where $X$ is a non-vanishing section of $\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\tilde{V}}))/\operatorname{ker}(\hbar\lambda_{\tilde{V}})$ and
+There is a splitting of vector bundles over $\widetilde{V}$ (well-defined up to contractible choice)
+$\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\widetilde{V}}))=\partial_{t}\oplus\langle\partial_{\theta},X\rangle\oplus\pi^{*}\xi,$
+[…]
   • (i)
-    $\langle\partial_{\theta},X\rangle$ is a trivial $2$-dimensional real symplectic vector bundle with respect to $d(\operatorname{re}(\hbar\lambda_{\tilde{V}}))$
+    $\langle\partial_{\theta},X\rangle$ is a trivial $2$-dimensional real symplectic vector bundle with respect to $)d(\operatorname{re}(\hbar\lambda_{\widetilde{V}}$))
   • (ii)
-    $\pi^{*}\xi$ is a complex symplectic vector bundle with respect to $d(\lambda_{\tilde{V}})$
+[…]
 Proof.
 The existence of this splitting is essentially a restatement of (7). The other statements can be checked locally using (6). ∎
-Given a complex contact manifold $(V,\mathcal{H})$ of complex dimension $12n+$ 1, a complex submanifold $L\subset V$ of complex dimension $n$ which is everywhere tangent to $\mathcal{H}$ is said to be a complex Legendrian.
+Given a complex contact manifold $(V,\mathcal{H})$ of complex dimension $2n+1$, a complex submanifold $L\subset V$ of complex dimension $n$ which is everywhere tangent to $\mathcal{H}$ is said to be a complex Legendrian.
 Lemma 2.8.
-If $L\subset(V,\xi)$ is complex Legendrian, then its preimage under $\tilde{V}\to V$ is denoted by $\tilde{L}$ and is complex exact Lagrangian.
-The quotient $\tilde{L}/\mathbb{R}_{+}\subset\tilde{V}/\mathbb{R}_{+}$ is a real Legendrian with respect to any of the $\mathbb{C}^{*}$ of contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.
+If $L\subset(V,\xi)$ is complex Legendrian, then its preimage under $\widetilde{V}\to V$ is denoted by $\widetilde{L}$ and is complex exact Lagrangian.
+The quotient $\widetilde{L}/\mathbb{R}_{+}\subset\widetilde{V}$$\mathbb{R}_{+}$ is a real Legendrian with respect to any of the $\mathbb{C}^{*}$ of contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.
 3. Grading and orientation data
-The material in this section is mostly algebraic topology.
-We will need it in Section 5 when we discuss how the condition of being a complex contact manifold/complex Legendrian interacts with orientations/gradings in microlocal sheaf theory.
 Consider the following inclusions of groups ($n/2$ ones defined only when $n$ even):
@@
 Here $\sqrt{SU}(n)$ is defined as the kernel of $U(n)\xrightarrow{det^{2}}U(1)$.
-Note that $U(n)$ is the maximal compact subgroup of both $Sp(2n,\mathbb{R})$ and $GL(n,\mathbb{C})$.
+Recall that $U(n)$ is the maximal compact subgroup of both $Sp(2n,\mathbb{R})$ and $GL(n,\mathbb{C})$, which retract to it; correspondingly, $BU(n)$ is canonically homotopy equivalent to the spaces classifying either complex or symplectic vector bundles.
 Meanwhile $U(n/2,\mathbb{H})$ is also known as the ‘compact symplectic group’ $Sp(n/2)=Sp(n,\mathbb{C})\cap U(n)$, the maximal compact subgroup of $Sp(n,\mathbb{C})$.
@@
 Consider a topological space $X$ carrying a hermitian bundle classified by a map $X\to BU(n)$. We give names to the following sorts of structures:
-  • A grading (or grading datum) is a lift to $B(\sqrt{SU}(n))$
+  • A grading is a lift to $B(\sqrt{SU}(n))$
   • A polarization is a lift to $BO(n,\mathbb{R})$
@@
 When $X$ is a symplectic manifold or contact manifold, by a polarization (etc.) on $X$, we always mean a polarization (etc.) on the symplectic tangent bundle of $X$ or the contact distribution.
+Concretely, a polarization of a symplectic vector bundle can be viewed as Lagrangian plane field, as follows by inspecting the fiber sequence $O(n)\to U(n)\to U(n)/O(n)=LGr(n)$.
+Similarly, a complex polarization of a complex symplectic vector bundle can be viewed as a complex Lagrangian plane field.
+We freely pass between both viewpoints in this paper.
 Lemma 3.2.
 A polarization induces a grading. A quaternionic structure induces a grading. A complex polarization induces both a quaternionic structure and a polarization, each of which induces the same grading. ∎
-Remark 3.3.
-We emphasize that the identifications in Definition 3.1 should be understood as holding up to coherent homotopy.
-So, for example, a polarization for $X\to BU(n)$ is a map $X\to BO(n)$ and a homotopy between the composition $X\to BO(n)\to BU(n)$ and the original structure map $X\to BU(n)$.
-Remark 3.4.
-Strictly speaking, a symplectic vector bundle $E$ on $X$ is classified by a map $X\to BSp(2n,\mathbb{R})$.
-However, $BU(n)\to BSp(2n,\mathbb{R})$ is a homotopy equivalence and we may always lift to some $X\to BU(n)$ uniquely up to contractible choice.
-(I.e.
-the space of Hermitian structures compatible with a given symplectic bundle is contractible.) Any Lagrangian distribution in $E$ will be totally real for any compatible Hermitian structure, hence define a reduction of structure to $O(n)$, i.e.
-what we have here called a polarization.
-Likewise, a complex symplectic bundle carries a quaternionic structure, uniquely up to contractible choices.
-We typically omit discussion of such issues in the sequel.
-In particular, for $(X,\omega)$ is complex symplectic, we simply say $TX$ carries a quaternionic structure.
-Remark 3.5.
+We will refer to gradings induced by any of the above structures as canonical gradings.
 Let us recall the relationship between null-homotopies and lifts.
@@
 Example 3.3.
-The sequence $O(n)\to U(n)\xrightarrow{det^{2}}S^{1}$ is a fiber sequence, which we may deloop to a fiber sequence $BO(n)\to BU(n)\xrightarrow{Bdet^{2}}BS^{1}$.
-So, a symplectic vector bundle $E$ admits a grading iff $02c_{1}(E)=$ 0.
-In particular, a symplectic manifold $X$ admits a grading iff $02c_{1}(X)=$ 0.
+A grading is a lift from $X\to BU(n)$ to $X\to B\sqrt{SU}(n)$, so equivalently, a null-homotopy of the composition $X\to BU(n)\xrightarrow{Bdet^{2}}BU(1)$.
+Recalling that homotopy class of $Bdet:BU(n)\to BU(1)=B^{2}\mathbb{Z}$ in $BU(n),B^{2}\mathbb{Z}]=\mathrm{H}^{2}(BU(n),\mathbb{Z})$ is the universal first Chern class, we see that the obstruction to the existence of a grading is twice the first Chern class.
 We consider the classical stablized compact groups
-$Sp=\lim_{n\to\infty}U(n,\mathbb{H})\qquad\qquad U=\lim_{n\to\infty}U(n,\mathbb{C})\qquad\qquad O=\lim_{n\to\infty}O(n,\mathbb{R})$
-and the Lagrangian Grassmannians
+$Sp=\lim_{n\to\infty}U(n,\mathbb{H})\qquad\qquad U=\lim_{n\to\infty}U(n,\mathbb{C})\qquad\qquad O=\lim_{n\to\infty}O(n,\mathbb{R}).$
+We write similarly $SU,\sqrt{SU}$.
+Note the inclusion $U(n,\mathbb{C})\to O(2n,\mathbb{R})$ lands in $SO(2n,\mathbb{R})$; correspondingly we have inclusions $U\subset SO\subset O$.
+We also have the Lagrangian Grassmannians
 $LGr_{\mathbb{C}}=Sp/U\qquad\qquad LGr=U/O$
 The natural inclusion $LGr_{\mathbb{C}}\to LGr$ is the limit of $U(n,\mathbb{H})/U(n,\mathbb{C})\hookrightarrow U(2n,\mathbb{C})/O(2n,\mathbb{R})$.
-There are evident stable analogues of the notions of Def. 3.1, and the stable analogue of Lemma 3.2 also holds.
+There are evident stable analogues of the notions of Definition 3.1, and the stable analogue of Lemma 3.2 also holds.
 Lemma 3.4.
-Let $(V,\xi)$ be a complex contact manifold and consider the real contact manifold $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$.
-Then the contact distribution $\xi_{\hbar}\to\tilde{V}/\mathbb{R}_{+}$ carries a stable quaternionic structure.
+Let $(V,\xi)$ be a complex contact manifold and consider the real contact manifold $(\widetilde{V}/\mathbb{R}_{+},\xi_{\hbar})$.
+Then the contact distribution $\xi_{\hbar}\to\widetilde{V}$$\mathbb{R}_{+}$ carries a stable quaternionic structure.
 Proof.
-It is equivalent to prove that $q^{*}\xi_{\hbar}$ carries a stable quaternionic structure, where $q:\tilde{V}\to\tilde{V}/\mathbb{R}_{+}$ is the quotient map. This follows from Lemma 2.7. ∎
-We write $w_{2}:BO\to B^{2}(\mathbb{Z}/2\mathbb{Z})$ for the map classified by the second Stiefel-Whitney class.
+It is equivalent to prove that $q^{*}\xi_{\hbar}$ carries a stable quaternionic structure, where $q:\widetilde{V}\to\widetilde{V}$$\mathbb{R}_{+}$ is the quotient map. This follows from Lemma 2.7. ∎
+One virtue of having passed to stabilizations is that $U/O$ is an infinite loop space.
+The map $det^{2}$ descends to an equivalence $det^{2}:\tau_{\leq 1}(U/O)\xrightarrow{\sim}B\mathbb{Z}$.
+Thus a grading on $X\to BU(n)$ is equivalently a null-homotopy of the composition $X\to BU(n)\to BU\to B(U/O)\to\tau_{\leq 2}B(U/O)$.
 Definition 3.5.
-Let $X$ be a topological space and consider a map $\nu:X\to BU$, i.e. a stable complex or symplectic vector bundle on $X$. Orientation data is a null-homotopy of the composite map:
+Grading/orientation data for $X\to BU(n)$ is a null-homotopy of the composition $X\to BU(n)\to\tau_{\leq 3}B(U/O)$.
+Note that a stable polarization for $X\to BU(n)$ is a null-homotopy of the composition $X\to BU\to B(U/O)$, so canonically provides grading/orientation data.
+We refer to such grading/orientation data as polarization grading/orientation data.
+Truncations give the fiber sequence (of infinite loop spaces)
+$B^{3}(\mathbb{Z}/2\mathbb{Z})\to\tau_{\leq 3}B(U/O)\to\tau_{\leq 2}B(U/O)=B\left(\tau_{\leq 1}(U/O)\right)\stackrel{{\scriptstyle Bdet^{2}}}{{=}}B^{2}\mathbb{Z}.$
+We refer to the space of grading/orientation data lifting a given grading as orientation data.
+Remark 3.6.
+It is presumably implicit in the construction of orientations on moduli spaces for Floer theory that there is a universal way to choose orientation data on all (real) symplectic manifolds.
+In fact there are two such ways; and their existence is somewhat subtle from our present viewpoint, as pointed out to us by Sanath Devalapurkar.
+The main source of the subtlety is the fact that the fiber sequence $B(\mathbb{Z}/2\mathbb{Z})\to\pi_{\leq 1}(U/O)\to\mathbb{Z}$ splits as a sequence of topological spaces, but not as a sequence of infinite loop spaces; in fact, $B^{3}(\mathbb{Z}/2\mathbb{Z})\to B^{2}\pi_{\leq 1}(U/O)\to B^{2}(\mathbb{Z})$ does not split as a sequence of spaces.
+We record his arguments in Appendix LABEL:sanath.
+This fact is not logically required for the present article, because the existence of orientation data for complex symplectic manifolds is less subtle, as will be shown presently.
+Lemma 3.7.
+The composition $Sp\to U\to U/O\to\tau_{\leq 2}(U/O)$ factors canonically through $\tau_{\leq 2}(Sp)=0$.
+Definition 3.8.
+Given a stable quaternionic bundle $X\to BSp$, the canonical grading/orientation datum of the induced complex bundle $X\to BSp\to BU$ is the null-homotopy induced from the null-homotopy of $Sp\to\tau_{\leq 2}(U/O)$ above.
+Similarly, the canonical grading is the null-homotopy induced from the similar null-homotopy of $Sp\rightarrow\tau_{\leq 1}(U/O)$.
+We note that the grading induced by the canonical grading/orientation datum agrees with the canonical grading of Lemma 3.2.
+Because $Sp\subset\sqrt{SU}$, the map $Sp/U\to U/O\xrightarrow{det^{2}}U(1)=\tau_{\leq 1}(U/O)$ is canonically null-homotopic.
+So the map $Sp/U\to U/O\to\tau_{\leq 2}(U/O)$ lifts to $Sp/U\to B^{2}(\mathbb{Z}/2\mathbb{Z})$.
+Lemma 3.9.
+This lift is the composition $Sp/U\to BU\xrightarrow{c_{1}}B^{2}\mathbb{Z}\to B^{2}(\mathbb{Z}/2\mathbb{Z})$, where $c_{1}$ is the first Chern class.
+Proof.
+Since $\tau_{\leq 2}Sp=0$, any map $Sp/U\to B^{2}(\mathbb{Z}/2\mathbb{Z})$ must factor through $BU$.
+There’s a unique nontrivial map in $BU,B^{2}(\mathbb{Z}/2\mathbb{Z})]$, and it’s the reduction mod 2 of the first Chern class.
+Finally, it is standard that $\tau_{\leq 2}(Sp/U\to U/O)$ is nontrivial: one has the map between fiber sequences
 […]
-The canonical orientation data is the null-homotopy arising from the canonical null-homotopy of $BU\to B(U/O)\to B^{2}O$.
-Evidently, the space of orientation data is identified (using the canonical orientation datum) with $Maps(X,B^{2}(\mathbb{Z}/2\mathbb{Z}))$; in particular, the choices up to homotopy are $H^{2}(X,\mathbb{Z}/2\mathbb{Z})$.
-[…]
-(8) $X\to BO\to BU\to B(U/O)\to B^{2}O\xrightarrow{Bw_{2}}B^{3}(\mathbb{Z}/2\mathbb{Z})$
-there are two evident null-homotopies, coming from the compositions $BO\to BU\to B(U/O)$ and $BU\to B(U/O)\to B^{2}O$.
-We term the first the polarization orientation data; the second is what we called above the canonical orientation data.
-Lemma 3.9.
-Fix a stable polarization $X\to BO$ of a given $X\to BU$. The space of homotopies between the canonical and polarization orientation data is equivalent to the space of lifts to $X\to BPin$.
+which induces a map between the long exact sequence of homotopy groups, which shows that $\pi_{2}(Sp/U)=\pi_{1}(U)=\mathbb{Z}$, $\pi_{2}(U/O)=\pi_{1}(O)=\mathbb{Z}/2$, and the map $\pi_{2}(Sp/U)\rightarrow\pi_{2}(U/O)$ is given by the quotient map $\mathbb{Z}\rightarrow\mathbb{Z}/2$.
+∎
+Proposition 3.10.
+Fix a stable complex polarization $X\to BU$ of a given $X\to BSp$.
+The space of homotopies between the canonical and polarization orientation data is equivalent to the space of null-homotopies of the composition $X\to BU\xrightarrow{c_{1}}B^{2}\mathbb{Z}\rightarrow B^{2}(\mathbb{Z}/2\mathbb{Z})$.
 Proof.
 First we recall some general facts about null-homotopies and exact triangles.
-Given two null-homotopies $n_{1},n_{2}$ of a given map $f:Q\to R$, we produce a pointed map $[n_{1},n_{2}]\in Q\times S^{1}\to R$ by taking $S^{1}=[-\pi,\pi]$, taking the map $f$ on $Q\times 0$ and applying the null-homotopy $n_{1}$ along $[-\pi,0]$ and the null-homotopy $n_{2}$ along $[0,\pi]$.
+Given two null-homotopies $n_{1},n_{2}$ of a given map $f:Q\to R$, we produce a pointed “comparison map” $n_{1},n_{2}]\in Q\times S^{1}\to R$ by taking $S^{1}=[-\pi,\pi]$, taking the map $f$ on $Q\times 0$ and applying the null-homotopy $n_{1}$ along $\pi,0]$ and the null-homotopy $n_{2}$ along $0,\pi]$.
 Note that $\operatorname{Hom}(Q\times S^{1},R)=\operatorname{Hom}(Q,\Omega R)$.
-A homotopy between $n_{1}$ and $n_{2}$ is a null-homotopy of $[n_{1},n_{2}]\in\operatorname{Hom}(Q,\Omega R)$.
+A homotopy between $n_{1}$ and $n_{2}$ is a null-homotopy of $n_{1},n_{2}]\in\operatorname{Hom}(Q,\Omega R)$.
 Suppose now given any exact triangle $P\to Q\to R\to BP$ in a stable category (for the application here, the stable category of spectra).
 The compositions $P\to Q\to R$ and $Q\to R\to BP$ give two null-homotopies of the composite map $P\to BP$.
-The definition of exact triangles [19, Def.
+The definition of exact triangles [30, Def.
 1.1.2.11] promises that the comparison of these two null-homotopies is identified with the identity of $\operatorname{Hom}(P,\Omega BP)=\operatorname{Hom}(P,P)$.
-[…]
-[…]
-given by $[s\circ n_{1}\circ p,s\circ n_{2}\circ p]\in\operatorname{Hom}(X,\Omega BS)$ is identified with $s\circ p\in\operatorname{Hom}(X,S)$. Here, $X$ need only be a space, not a spectrum.
-Now specializing to (8), we find that the space of homotopies between polarization and canonical orientation data is identified with the space of null-homotopies of $X\to BO\xrightarrow{w_{2}}B^{2}(\mathbb{Z}/2\mathbb{Z})$, i.e., the space of lifts to $X\to BPin$.
+Given maps $X\xrightarrow{p}P\xrightarrow{s}S$, we may compose to learn that the corresponding comparison between null-homotopies of
+$X\xrightarrow{p}P\to Q\to R\to BP\xrightarrow{Bs}BS$
+given by $s\circ n_{1}\circ p,s\circ n_{2}\circ p]\in\operatorname{Hom}(X,\Omega BS)$ is identified with $s\circ p\in\operatorname{Hom}(X,S)$. Here, $X$ need only be a space, not a spectrum.
+Now consider a stable quaternionic vector bundle with a stable complex polarization, i.e. we have a lift $X\to BU\to BSp$. Consider the composition
+(8) $X\to BU\to BSp\to B(Sp/U)\to B^{2}U\xrightarrow{Bc_{1}}B^{3}\mathbb{Z}\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$
+Now, the canonical orientation data factors through $BSp\to 0$, hence is induced by the null-homotopy of the sequence $BSp\to B(Sp/U)\to B^{2}U$.
+The polarization orientation data comes from the prescribed null-homotopy of $X\to B(U/O)$; since we have a complex polarization this factors through a null-homotopy of $X\to B(Sp/U)$, hence is induced from the null-homotopy of $BU\to BSp\to B(Sp/U)$.
+The result follows.
 ∎
+We recall that the mod 2 reduction of the first Chern class of a complex bundle is the second Stiefel-Whitney class $w_{2}$ of the underlying real bundle.
+Thus, null-homotopies of $X\to BU\xrightarrow{w_{2}}B^{2}(\mathbb{Z}/2\mathbb{Z})$ as above are the same as spin structures on the underlying real bundle classified by $X\to BU\to BSO$.
+Let $V$ be a symplectic (resp.
+contact) manifold.
+Given a Lagrangian (resp.
+Legendrian) submanifold, the Weinstein neighborhood theorem symplectomorphically (resp.
+contactomorphically) identifies a neighborhood of $L\subset V$ with a neighborhood of the zero section in $T^{*}L$ (resp.
+in $J^{1}L$).
+This identification is canonical up to contractible choice.
+We write $\phi_{L}$ for the fiber polarization of $T^{*}L$ (resp.
+$J^{1}L$).
 Definition 3.11.
-Let $V$ be a symplectic or contact manifold. Fix a polarization / grading / orientation data $m$ on $V$.
-Consider $L\subset V$ a Lagrangian or Legendrian.
-There is a canonical isomorphism of $T^{*}L$ with the restriction of the symplectic tangent bundle or contact distribution.
-We write $\phi_{L}$ for the fiber polarization.
-A relative polarization for $L$ is a homotopy between $m|_{L}$ and $\phi_{L}$.
-Relative grading / orientation data are similarly homotopies between the grading / orientation data induced by $m|_{L}$ and $\phi_{L}$.
-Note that the obstruction to existence of a relative grading is a class in $[L,B\mathbb{Z}]=H^{1}(L,\mathbb{Z})$, and the space of relative gradings is a torsor for $Map(L,\mathbb{Z})$, hence in particular, is discrete.
-Thus we simply ask whether relative gradings are equal, rather than discuss homotopies between them.
+Let $m$ be a grading (resp. grading/orientation datum) on $V$. Then by a secondary grading (resp. grading/orientation datum) for $L\subset V$, we mean a homotopy between $m|_{L}$ and $\phi_{L}$.
+Given grading/orientation data on $V$ and a secondary grading on $L$, then by secondary orientation data, we mean a lift of said secondary grading on $L$ to secondary grading/orientation data on $L$.
+Note that the obstruction to existence of a secondary grading is a class in $L,B\mathbb{Z}]=H^{1}(L,\mathbb{Z})$, and the space of secondary gradings is a torsor for $\operatorname{Map}(L,\mathbb{Z})$, hence in particular, is discrete.
+Thus we simply ask whether secondary gradings are equal, rather than discuss homotopies between them.
 Lemma 3.12.
-Let $V$ is a complex contact manifold, $\tilde{V}\to V$ the corresponding real contact manifold equipped with the canonical grading constructed in Lemma 3.2.
-Let $L\subset V$ be a smooth complex Legendrian.
-Then there is a canonical choice for the relative grading of $\tilde{L}\subset\tilde{V}$.
+Let $V$ be a complex symplectic manifold.
+Equip the underlying real symplectic manifold with the canonical grading constructed in Lemma 3.2.
+Let $L\subset V$ be a smooth complex Lagrangian.
+Then there is a canonical choice for the secondary grading of $L\subset V$.
 Proof.
@@
 Lemma 3.13.
-Fix the canonical orientation data on $V$.
-Then relative orientation data for $L\subset V$ is equivalent to a pin structure on $L$.
-(If $L$ is oriented, then a pin structure on $L$ is equivalent to a spin structure on $L$ compatible with the given orientation.)
+Let $L,V$ be as in Lemma 3.12.
+Fix the canonical grading/orientation data on $V$, and the canonical grading on $L$.
+Then secondary orientation data for $L\subset V$ is equivalent to a spin structure on $L$.
 Proof.
@@
 Let $M$ be a real manifold.
-Fix a symmetric monoidal stable presentable category , $C$.
-The reader will not lose much of the point of the paper taking throughout $C$ to be the derived category of dg modules over some commutative ring $R$.
-We write $sh(M)$ for the (stable) category of sheaves on $M$ with values in $C$.
+Fix a symmetric monoidal stable presentable compactly generated category, $\mathcal{C}$.
+The reader will not lose much of the point of the paper taking throughout $\mathcal{C}$ to be the derived category of dg modules over some commutative ring $R$.
+We write $sh(M)$ for the (stable) category of sheaves on $M$ with values in $\mathcal{C}$.
 In this subsection we review ideas from Kashiwara and Schapira [25].
@@
 The microsupport is easily seen to be conical and satisfy $ss(Cone(F\to G))\subset ss(F)\cup ss(G)$.
-A deep result of [15, Thm.
-6.5.4] is that the microsupport is coisotropic (also called involutive; see [15, Def.
+A deep result of [25, Thm.
+6.5.4] is that the microsupport is coisotropic (also called involutive; see [25, Def.
 6.5.1]).
@@
 For any subset $\Lambda\subset S^{*}M$, we write $sh_{\Lambda}(M):=sh_{\mathbb{R}_{+}\Lambda\cup 0_{M}}(M)$, with $0_{M}$ the zero section of the cotangent bundle.
+We write $T^{\circ}M:=T^{*}M-0_{M}$; following the usual convention, we will not distinguish between subsets of $S^{*}M$ and conic subsets of $T^{*}M-0_{M}$.
 The assignment $U\mapsto sh(U)$ defines a sheaf of categories on $M$; we denote it $sh$.
@@
 We consider the presheaf of stable categories on $T^{*}M$:
-(9) $\mu sh_{T^{*}M}^{pre}(U):=sh(M)/sh_{T^{*}M\setminus U}(M)$
+(9) $\mu sh_{T^{*}M}^{pre}(\Omega):=sh(M)/sh_{T^{*}M\setminus\Omega}(M)$
 Definition 4.4.
-Let $\mu sh_{T^{*}M}$ be the sheaf of categories on $T^{*}M$ defined by sheafifying the presheaf $sh_{T^{*}M}^{pre}(U)$ in (9).
+Let $\mu sh_{T^{*}M}$ be the sheaf of categories on $T^{*}M$ defined by sheafifying the presheaf $sh_{T^{*}M}^{pre}(\Omega)$ in (9).
 Similarly, let $\mu sh_{S^{*}M}$ be the presheaf of categories on $S^{*}M$ obtained by sheafifying $\mu sh^{pre}_{S^{*}M}$.
+In any sheaf of categories $\mathcal{X}$ on a topological space $T$, given $F,G\in\mathcal{X}(T)$, the assignment $U\mapsto\operatorname{Hom}_{\mathcal{X}(U)}(F|_{\Omega},G|_{\Omega})$ is a sheaf on $T$; let us denote it as $\mathcal{H}om_{\mathcal{X}}(F,G)$.
+A fundamental result of Kashiwara and Schapira computes this Hom sheaf via sheaf operations [25, Thm.
+6.1.2]: [^1] More precisely, [25] shows there is a morphism of this kind for $\mu sh^{pre}$ which is an isomorphism at stalks; the stated result follows upon sheafification.
+See [35] for some detailed discussions about the sheafification of $\mu sh^{pre}$.
+(10) $\mathcal{H}om_{\mu sh}(F,G)=\mu hom(F,G):=\mu_{\Delta}\operatorname{Hom}_{M\times M}(\pi_{1}^{*}F,\pi_{2}^{!}G).$
 For $F\in sh(M)$, one finds that the support of the image of $F$ in $\mu sh(\Omega)$ is $ss(F)\cap\Omega$.
@@
 This being a contractible $\mathbb{R}^{+}$, we may define a sheaf of categories $\mu sh_{S^{*}M}$ on $S^{*}M$ equivalently by pushforward or pullback along an arbitrary section of $T^{\circ}M\to S^{*}M$.
+As $ss(G)$ is conic for $G\in\mu sh_{S^{*}M}$, it uniquely determine a set $ss^{\infty}(G)\subseteq S^{*}M$.
 Definition 4.5.
@@
 Similarly, for $\Lambda\subset S^{*}M$, let $\mu sh_{\Lambda}\subset\mu sh_{S^{*}M}$ be the subsheaf of full subcategories on objects supported in $\Lambda$.
-The following is shown in [15, Prop. 6.6.1]:
 Proposition 4.6.
-Let $M$ be a manifold and let $N\subset M$ be a submanifold. Let $L=T^{*}_{N}M$ and fix a point $p\in L$. Then there is an equivalence of categories:
-(10) $\omega_{p}:C$ $\xrightarrow{\sim}(\mu sh_{L,\phi})_{p}$
+[25, Prop. 6.6.1] Let $M$ be a manifold and let $N\subset M$ be a submanifold. Let $L=T^{*}_{N}M$ and fix a point $p\in L$. Then there is an equivalence of categories:
+(11) $\mathcal{C}$ $\xrightarrow{\sim}(\mu sh_{L})_{p}$
 $A$ $\mapsto A_{N},$
 where $A_{N}$ is the image in $\mu sh$ of the constant sheaf on $N$ with value $A$. The corresponding result of course holds, and we use the same notations, in $S^{*}M$.
-Remark 4.7.
-Obviously $\omega_{p}$ depends only on the germ of $M$ near the image of $p$; for this remark we regard $M,N$ as germs.
-Consider another triple $M^{\prime},N^{\prime},L^{\prime}=T^{*}_{N^{\prime}}M^{\prime},p^{\prime}\in L^{\prime}$.
-If $\dim M=\dim M^{\prime}$ and $\dim N=\dim N^{\prime}$, then there is a diffeomorphism $(M,N)\to(M^{\prime},N^{\prime})$ whose canonical lift to the cotangent bundle sends $p\to p^{\prime}$.
-(Indeed, note that $p$ may be moved arbitrarily in $L\setminus 0$ by diffeomorphisms of $M$ fixing $N$.) Obviously the construction of $\omega_{p}$ is functorial in such diffeomorphisms.
-In any sheaf of categories $\mathcal{X}$ on a topological space $T$, given $F,G\in\mathcal{X}(T)$, the assignment $U\mapsto\operatorname{Hom}_{\mathcal{X}(U)}(F|_{U},G|_{U})$ is a sheaf on $T$; let us denote it as $\mathcal{H}om_{\mathcal{X}}(F,G)$.
-Per [15, Thm.
-6.1.2]: [^1] More precisely, [15] show there is a morphism of this kind for $\mu sh^{pre}$ which is an isomorphism at stalks; the stated result follows upon sheafification.
-See [22] for some detailed discussions about the sheafification of $\mu sh^{pre}$.
-(11) $\mathcal{H}om_{\mu sh}(F,G)=\mu hom(F,G)$
-where $\mu hom$ is a functor explicitly defined in terms of standard sheaf theoretic operations, namely the Fourier-Sato transformation along the diagonal of the external Hom:
-$\mu hom(F,G):=\mu_{\Delta}\operatorname{Hom}_{M\times M}(\pi_{1}^{!}F,\pi_{2}^{*}G)$
 Definition 4.7.
@@
 For $M$ complex analytic, we define $\mu sh_{\mathbb{C}-c}\subset\mu sh_{\mathbb{R}-c}$ as the subsheaf of full subcategories on objects whose support is complex analytic and the closure of its smooth Lagrangian locus.
-4.3. Microsheaves on polarized real contact manifolds
-A key tool of [15] is the functoriality of $\mu sh$ under (quantized) contact transformation [15, Sec. 7]. One readily obtains:
+Alternatively, one can also restrict to constructible sheaves all from the beginning and define
+(12) $\mu sh_{T^{*}M,\mathbb{R}-c}^{pre}(\Omega):=sh_{\mathbb{R}-c}(M)/sh_{T^{*}M\setminus\Omega,\mathbb{R}-c}(M),$
+and consider its sheafification. For a conic Lagrangian $\Lambda\subseteq T^{*}M$, define
+(13) $\mu sh_{\mathbb{R}-c,\Lambda}^{pre}(\Omega):=sh_{\mathbb{R}-c,\Lambda\cup(T^{*}M\setminus\Omega)}(M)/sh_{\mathbb{R}-c,T^{*}M\setminus\Omega}(M).$
+Unwrapping the definition, we see that $sh_{\mathbb{R}-c,\Lambda\cup(T^{*}M\setminus\Omega)}(M)=\{F\in sh_{\mathbb{R}-c}(M)|ss(F)\cap\Omega\subseteq\Lambda\}$.
+Similarly, one can define $\mu sh^{pre}_{T^{*}M,\mathbb{C}-c}$ when $M$ is complex, and $\mu sh^{pre}_{\Lambda,\mathbb{C}-c}$ when $\Lambda$ is complex.
+By microlocal cut-off, these notions agree with the ones given previously.
 Lemma 4.8.
-Suppose given a contact manifold $V$ and a contact embedding $\iota:V\times T^{*}D^{n}$, for some disk $D$.
-Then $\mu sh_{\iota(V)\times D^{n}}$ is locally constant along $D^{n}$, hence the pullback of a sheaf of categories along $V\times D^{n}\to V$.
+[42, Theorem 3.2.2] The canonical map $\mu sh_{T^{*}M,\mathbb{R}-c}^{pre}\rightarrow\mu sh_{T^{*}M,\mathbb{R}-c}$ induces an isomorphism upon sheafification.
+The similar statement holds for $\mu sh_{\mathbb{R}-c,\Lambda}$, $\mu sh_{T^{*}M,\mathbb{C}-c}$, and $\mu sh_{\mathbb{C}-c,\Lambda}$.
+We will need the following lemma in the next section.
+Let $i:N\hookrightarrow M$ be an inclusion of closed submanifolds.
+The map $i$ induces an inclusion $i_{\pi}:T^{*}M|_{N}\hookrightarrow T^{*}M$ and the transpose of its derivative induces a projection
+$(di)^{t}:T^{*}M|_{N}$ $\rightarrow T^{*}N$
+$(y,\xi)$ $\mapsto(y,\xi\circ di_{y}).$
+Lemma 4.9.
+The *-pushforward, $i_{*}:sh(N)\xrightarrow{\sim}sh_{N}(M)$, microlocalizes to an equivalence
+$(di)^{t}]^{*}\mu sh_{T^{*}N}\xrightarrow{\sim}\mu sh_{T^{*}M|_{N}}.$
+(Here $\mu sh_{T^{*}M|_{N}}\subset\mu sh_{T^{*}M}$ is the subsheaf of full subcategories on objects supported in $T^{*}M|_{N}\subset T^{*}M$; see Definition 4.5.)
+Proof.
+Since $i$ is a closed embedding, [25, Proposition 5.4.4] implies that $ss(i_{*}F)=({(di)^{t}})^{-1}ss(F)$ is contained in $T^{*}M|_{N}$ for $F\in sh(N)$. This implies that the assignment
+$\mu sh_{T^{*}N}^{pre}(\Omega)$ $\rightarrow\left(({(di)^{t}})_{*}{i_{\pi}}^{*}\mu sh_{T^{*}M|_{N}}^{pre}\right)(\Omega)$
+$F$ $\mapsto({i_{\pi}})_{*}F$
+is well-defined.
+To check the induced map $({(di)^{t}})^{*}\mu sh_{T^{*}N}\rightarrow{i_{\pi}}^{*}\mu sh_{T^{*}M}$ on sheaves is an equivalence, one can check at stalks.
+Fully faithfulness is then implied by [25, Proposition 4.4.7(ii)], as (10) shows that the Hom is computed by $\mu hom$.
+Essential surjectivity is implied by [25, Prop.
+6.6.1].
+∎
+Before we leave this section, we mention two common tools for studying microsheaves. The first is the contact transformation.
+Theorem 4.10 ([25, Corollary 7.2.2]).
+Let $\mathcal{U}\subseteq S^{*}M$, and $\mathcal{V}\subseteq S^{*}N$ be open sets, and $\chi:\mathcal{U}\xrightarrow{\mathcal{V}}$ be a contactomorphism.
+Then, for any given $p\in\mathcal{U}$, shrink $\mathcal{U}$ if needed, one can assume that there exists a sheaf $K\in sh(M\times N)$ such that the functor $\Phi_{K}:sh(M)\rightarrow sh(N)$ given by convolving with $K$ induces an equivalence
+$\Phi_{K}:\mu sh^{pre}_{S^{*}M}|_{\mathcal{U}}\xrightarrow{\sim}\chi^{*}\left(\mu sh^{pre}_{S^{*}N}|_{\mathcal{V}}\right).$
+Consequently, it induces an equivalence $\mu sh_{S^{*}M}|_{\mathcal{U}}\xrightarrow{\sim}\chi^{*}\mu sh_{S^{*}N}|_{\mathcal{V}}$ which commutes with the canonical map $\mu sh^{pre}\rightarrow\mu sh$.
+The other one is the doubling trick, which states that, if $\mathcal{F}\in\mu sh(\Omega)$, for some open set $\Omega\subseteq S^{*}M$, is a microsheaf whose support $ss^{\infty}(\mathcal{F})$ is contained in some closed Legendrian $\Lambda\subseteq S^{*}M$, then $\mathcal{F}$ can be represented by a genuine sheaf $F$ (with possibly larger microsupport).
+Theorem 4.11 ([35, Theorem 7.18], [28, Theorem 4.47] ).
+Let $\Lambda\subseteq S^{*}M$ be a closed Legendrian and $\Omega\subseteq S^{*}M$ an open set.
+Choose a positive isotopy $\Lambda_{t}$ of $\Lambda$ in $S^{*}M$ such that $\alpha(\partial_{t}\Lambda_{t})$ is nonzero on $\Omega$ and outside $\Omega$.
+Then, for some small enough $\epsilon$, the composition
+$sh_{\Lambda\cup\Lambda_{\epsilon}}(M)\rightarrow\mu sh_{\Lambda\cup\Lambda_{\epsilon}}(\Omega)\rightarrow\mu sh_{\Lambda}(\Omega)$
+is surjective. Here, we use the fact that $\Lambda$ and $\Lambda_{\epsilon}$ are disjoint in $\Omega$ for the second arrow.
+4.3. Microsheaves and Maslov index
+In this subsection, we describe following [6] the sheaf quantization of a rotation of $J^{1}\mathbb{R}^{n}$. This material will be used explicitly only in Section 5.2.
+Endow $T^{*}\mathbb{R}^{n}\times T^{*}\mathbb{R}$ with coordinates $(x,\xi,t,\tau)$.
+Identify the $1$-jet space $J^{1}\mathbb{R}^{n}$ with the hyperplane $\{\tau=1\}\subset T^{*}\mathbb{R}^{n}\times T^{*}\mathbb{R}$; the contact form restricted to $\{\tau=1\}$ is
+$\alpha\;=\;\xi\cdot dx+dt.$
+Fix $1\leq\ell\leq n$.
+As in [6], we consider a $1$-parameter family of contactomorphisms $\psi_{t}:J^{1}\mathbb{R}^{n}\to J^{1}\mathbb{R}^{n}$ $\psi_{t}:J^{1}\mathbb{R}^{n}\to J^{1}\mathbb{R}^{n}$ given by
+$\psi_{t}(x,\xi,\tau)\;=\;(x^{\prime},\xi^{\prime},\tau^{\prime})$
+with (for $i=1,\dots,\ell$)
+$x^{\prime}_{i}$ $=(\cos t)\,x_{i}-(\sin t)\,\xi_{i},$
+$\xi^{\prime}_{i}$ $=(\sin t)\,x_{i}+(\cos t)\,\xi_{i},$
+and for $i=\ell+1,\dots,n$
+$x^{\prime}_{i}=x_{i},\qquad\xi^{\prime}_{i}=\xi_{i}.$
+The $\tau$-coordinate transforms by
+$\tau^{\prime}\;=\;\tau+F_{t}(x,\xi),$
+where,
+(14) $F_{t}(x,\xi)=\sum_{i=1}^{\ell}\left(\frac{\sin t\cos t}{2}\,(x_{i}^{2}-\xi_{i}^{2})-2\sin^{2}t\cdot x_{i}\cdot\xi_{i}\right).$
+Note that $\psi_{t}$ is the lift of a $1$-parameter family of sympectomorphisms of $T^{*}\mathbb{R}^{n}$ which rotates the $(x_{i},\xi_{i})$ coordinates clockwise for $1\leq i\leq\ell$.
+These symplectomorphisms admit a sheaf quantization:
+Proposition 4.12 ([6, Proposition 3.10] [^2] As the Hamiltonian is not compactly supported, the quantization does not follow directly from the usual result of Guillermou, Kashiwara, and Schapira [15, Proposition A.6].).
+Consider the space $\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}_{t}\times\mathbb{R}_{s}$ where we view $s$ as the time direction.
+Then, there exists a sheaf kernel $\mathcal{S}\in sh(\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}_{t}\times\mathbb{R}_{s})$ such that
+  1. (1)
+    $\mathcal{S}|_{s=0}=1_{\Delta_{\mathbb{R}^{n}}}\boxtimes 1_{[0,\infty)}$.
+  2. (2)
+    $ss(\mathcal{S})\cap\{\sigma>0\}\subseteq\Lambda_{\Psi}$ where we use $\sigma$ to denote the cotangent coordinate for $\mathbb{R}_{s}$ and $\Lambda_{\Psi}$ is the Legendrian lift of the isotopy movie of the Hamiltonian rotation.
+  3. (3)
+    Over the open interval $s\in(0,\pi)$, we have the explicit formula,
+    $\mathcal{S}|_{(0,\pi)_{s}}=1_{\{(x,y,t,s)|t+F_{2s}(x,y)\geq 0\}}.$
+For a sheaf kernel $\mathcal{K}\in sh(M\times N\times\mathbb{R}_{t})$, one can consider the $\star$-convolution
+$sh(M\times\mathbb{R}_{t_{1}})$ $\rightarrow sh(N\times\mathbb{R}_{t})$
+$\mathcal{F}$ $\mapsto\mathcal{K}\star\mathcal{F}\coloneqq{(p_{N,+})}_{!}(p_{2}^{*}\mathcal{K}\otimes p_{M,1}^{*}\mathcal{F})$
+where $p_{M,1}$ and $p_{2}$ are the usual projections from $M\times N\times\mathbb{R}_{t_{1}}\times\mathbb{R}_{t_{2}}$ to $M\times\mathbb{R}_{t_{1}}$ and $M\times N\times\mathbb{R}_{t_{2}}$ but, when pushing forward, we use the addition for the $t$ direction by $p_{N,+}(x,y,t_{1},t_{2})=(y,t)$, $t=t_{1}+t_{2}$ instead.
+By (2) of the above Proposition 4.12, $\mathcal{S}\star(-)$ restricts to an equivalence on $\mu sh^{pre}|_{\{\sigma>0\}}$ and hence to the restriction of $\mu sh$ to $\{\sigma>0\}$.
+Let $K:=\{\xi=0,\ \tau=0\}=0_{\mathbb{R}^{n}}\times\{0\}\subset J^{1}\mathbb{R}^{n}$ be the zero section, and write $p=(0,0,0)$ for the origin.
+Set $K_{t}:=\psi_{t}(K)$ and note that $K_{\pi/2}$ is the conormal to $C_{k}\coloneqq\{x_{1}=\cdots=x_{k}=0\}\subset\mathbb{R}^{n}$.
+Lemma 4.13.
+We have the following diagram:
+(15) $\mathcal{C}$$(\mu sh_{K})_{p}$$\mathcal{C}$$(\mu sh_{K_{\pi/2}})_{p}$$\scriptstyle{A\mapsto A_{L}\boxtimes 1_{[0,\infty)}}$$\scriptstyle{[-k]}$$\scriptstyle{A\mapsto A_{C_{k}}\boxtimes 1_{[0,\infty)}}$
+where the right downward arrow is induced by the sheaf quantization of $\psi_{\pi/2}$ furnished by Proposition 4.12.
+Proof.
+By the Künneth formula, it is sufficient to prove the case for $l=1$.
+Since we are rotating the Lagrangian by $90^{o}$, we set $s=\pi/2$ and we have $\mathcal{S}|_{s=\pi/2}=1_{\{t-\frac{1}{2}x\xi\}}$ by (14).
+Thus, seeing the commutativity amounts to computing the object
+$\mathcal{S}|_{s=\pi/2}\star(A_{\mathbb{R}^{1}}\boxtimes 1_{[0,\infty)_{t}})=(p_{2,+})_{!}\left(A_{\{(x,y,t_{1},t_{2})|t_{1}\geq 0,\ t_{2}-\frac{1}{2}\geq 0\}}\right).$
+Now, the projection $p_{2,+}$ can be decomposed to the naive projection $q(x,y,t_{1},t_{2})=(y,t_{1},t_{2})$ follows by the addition $a(y,t_{1},t_{2})=(y,t_{1}+t_{2})$, we can thus compute the pushforward in two steps.
+The claim is that $q_{!}\left(A_{\{(x,y,t_{1},t_{2})|t_{1}\geq 0,\ t_{2}-\frac{1}{2}xy\geq 0\}}\right)=A_{\{0\}}[-1]\boxtimes 1_{[0,\infty)_{t_{1}}}\boxtimes 1_{[0,\infty)_{t_{2}}}$, and we will be done because the projection formula will imply that
+$a_{!}\left(A_{\{0\}}[-1]\boxtimes 1_{[0,\infty)_{t_{1}}}\boxtimes 1_{[0,\infty)_{t_{2}}}\right)=A_{\{0\}}[-1]\boxtimes(1_{[0,\infty)_{t_{1}}}\star 1_{[0,\infty)_{t_{2}})})=A_{\{0\}}[-1]\boxtimes 1_{[0,\infty)_{t}}$
+as we desired.
+To see the claim, again by the projection formula, we can ignore $t_{1}$ and we only have to show that, if $f(x,y,t)\coloneqq(y,t)$, then $f_{!}(A_{\{t-\frac{1}{2}xy\geq 0\}})=A_{(\{0,t)|t\geq 0\}}[-1]=A_{\{0\}}[-1]\boxtimes 1_{[0,\infty)}$.
+Lastly, we can check on stalks, and consider, for $(y,t)\in\mathbb{R}^{2}_{(y,t)}$,
+$\left(f_{!}(A_{\{t-\frac{1}{2}xy\geq 0\}})\right)_{(y,t)}=\Gamma_{c}(\{x\in\mathbb{R}|t-(1/2)xy\geq 0\};A)$
+the compactly supported sections on the set $\{x\in\mathbb{R}|t-\frac{1}{2}xy\geq 0\}$.
+We note that, when $y>0$, $t-\frac{1}{2}xy$ is equivalent to $x\leq\frac{2t}{y}$ so for any pair $(y,t)$ the set $\{x\in\mathbb{R}|t-\frac{1}{2}xy\geq 0\}$ is topologically $0,\infty)$ which has a vanishing compactly supported cohomology.
+The same situation holds when $y<0$.
+Thus, the only non-trivial stalks live over $y=0$, in which case $x$ has no condition.
+In other words, the stalk is given by $\Gamma_{c}(\mathbb{R};A)=A[-1]$ and is exactly what we wanted.
+∎
+Remark 4.14.
+Lemma 4.13 can presumably also be extracted from the discussion in [25, Sec. 7.5, Appendix A]. Similar computations have also appeared in e.g. [14, 17].
+4.4. Microsheaves on polarized (real) contact manifolds
+In the previous section, we discussed microlocal sheaves on cotangent/cosphere bundles.
+The high-codimensional embedding trick of [39] extends the definition of microlocal sheaves to arbitrary contact or exact symplectic manifolds equipped with Maslov data.
+In this section, we review how this works on polarized contact manifolds; we postpone the discussion of Maslov data to the next section.
+The starting point of [39] is the following lemma, which is a consequence of the functoriality of $\mu sh$ under (quantized) contact transformation [25, Sec. 7].
+Lemma 4.15.
+Suppose given a contact manifold $V$ and a contact embedding $\iota:V\times T^{*}D^{n}\hookrightarrow S^{*}M$, for some disk $D^{n}$ and some manifold $M$.
+Then $\mu sh_{\iota(V\times D^{n})}$ is locally constant along $D^{n}$, hence the pullback of a sheaf of categories along $V\times D^{n}\to V$.
+Proof.
+The statement is local, so we may assume $V$ is a Darboux ball.
+By the contact neighborhood theorem [13, Theorem 2.5.15], there is an open ball $U\subset S^{*}\mathbb{R}^{n+m}$ and a contactomorphisms $(V\times T^{*}D^{n},V\times\{0\})\simeq(U,U\cap S^{*}\mathbb{R}^{n})$, where $S^{*}\mathbb{R}^{m}\subset S^{*}\mathbb{R}^{n+m}$ is induced by the standard inclusion $\mathbb{R}^{m}=\mathbb{R}^{n}\times\{0\}^{n}\hookrightarrow\mathbb{R}^{m+n}$.
+By the functoriality of $\mu sh$ under (quantized) contact transformation [25, Sec.
+7], we may therefore assume without loss of generality that $V=U\cap S^{*}\mathbb{R}^{n}$.
+But, by Lemma 4.9, we conclude that $\mu sh_{S^{*}\mathbb{R}^{m}\times 0_{\mathbb{R}^{n}}}=q^{*}\mu sh_{S^{*}\mathbb{R}^{m}}$ where $q:S^{*}\mathbb{R}^{m}\times T^{*}\mathbb{R}^{n}\rightarrow S^{*}\mathbb{R}^{m}$ is the projection.
+∎
 The basic idea of [39] is to use Lemma 4.15 as a definition of $\mu sh_{V}$.
 Let $(V,\xi)$ be a contact manifold.
-Suppose given a (possibly positive codimensional) contact embedding $\iota:V\to S^{*}M$, and any Lagrangian distribution $\eta$ of the symplectic normal bundle to $\iota$.
-Let us write $V(\eta)$ for a thickening of $\iota(V)$ along $\eta$ in a tubular neighborhood.
-Then $\iota^{*}\mu sh_{V(\eta)}$ provides a sheaf of categories on $V$.
-By contact transformations, it is easy to see that this sheaf of categories is locally equivalent to microsheaves on a contact cosphere bundle of the same dimension.
+Suppose given a (possibly positive codimensional) contact embedding $\iota:V\to S^{*}M$ and any Lagrangian distribution $\eta$ of the symplectic normal bundle $N_{V}$ to $\iota$.
+The choice of $\eta$ comes with a coisotropic embedding $\tau_{\eta}:\operatorname{Op}(0_{V}\cap\eta)\to S^{*}M$ such that the differential satisfies $\operatorname{im}(d\tau_{\eta}:\{0\}\times\eta(p)\subset T_{p}V\times(N_{V})_{p}\to T_{\tau_{\eta}(p,0)}S^{*}M)=\eta(p),$ for all $p\in V$.
+Here $0_{V}$ is the zero section of the symplectic normal bundle $N_{V}$.
+We denote by $V(\eta)$ the image of $\tau_{\eta}$ and call it a thickening of $\iota(V)$ along $\eta$.
+Lemma 4.16.
+The sheaf of categories $\mu sh_{V(\eta)}$ is the pullback of certain a sheaf of categories on $V$. We will denote this sheaf as $\mu sh_{V,\iota,\eta}$.
+Proof.
+The Lagrangian bundle structure $\eta$ provides, near $V$, a contraction $r:V(\eta)\rightarrow V$, and thus defines a sheaf $\mu sh_{V,\iota,\eta}\coloneqq r_{*}\mu sh_{V(\eta)}$ by pushforward.
+However, the map $r$ is locally given by the projection $V\times D^{n}\rightarrow V$, and we’ve shown in Lemma 4.15 that the sheaf $\mu sh_{V(\eta)}$ is constant along the fiber direction, so $\mu sh_{V(\eta)}$ can be recovered from $\mu sh_{V,\iota,\eta}$ by
+$\mu sh_{V(\eta)}=r^{*}r_{*}\mu sh_{V(\eta)}=r^{*}\mu sh_{V,\iota,\eta}.$
+This completes the proof. ∎
 To eliminate the dependence on $\iota$ and $M$, one notes that Gromov’s $h$-principle for contact embeddings implies the existence of high codimension embeddings of $V$ into the standard contact $\mathbb{R}^{2n+1}$ for large enough $n$, the space of which moreover becomes arbitrarily connected as $n\to\infty$.
 There remains the (continuous) dependence on a choice of polarization of the stable symplectic normal bundle (“stable normal polarization”).
+That is:
 Theorem 4.17.
 [39] Given a contact manifold $(V,\xi)$ and a polarization $\eta$ of the stable symplectic normal bundle to $V$, there is a canonical sheaf of categories $\mu sh_{V,-\eta}$ on $V$, locally isomorphic to $\mu sh_{S^{*}M}$ for any $M$ of the appropriate dimension.
-Above, $-\eta$ can be regarded as just a notational choice.
+Above, the sign - on $\eta$ can be regarded as just a notational choice.
 To give it an actual meaning, note that $BO\to BU$ is a morphism of spectra, so given a stable symplectic vector bundle $E$ on some topological space $X$, say classified by some map $E:X\to BU$, and a polarization of $E$, i.e.
-lift to some $F:X\to BO$, then $-F$ gives a polarization of $-E$, where -- is the pullback by the canonical ‘inverse’ involution on $BO$ or $BU$.
-In fact [22, Sec.
+lift to some $F:X\to BO$, then F-F gives a polarization of E-E, where - is the pullback by the canonical ‘inverse’ involution on $BO$ or $BU$.
+In fact [35, Sec.
 10.2], it is always possible to build a contact manifold, $V^{T^{*}LGr(\xi)}$, which is a bundle over $V$ with fibers the cotangent bundles to the Lagrangian Grassmannians of the contact distribution $\xi$.
@@
 Thus we define $\mu sh_{V^{T^{*}LGr(\xi)}}$, and, restricting supports, $\mu sh_{V^{LGr(\xi)}}$.
+It is evident from the construction that $\mu sh_{V^{LGr(\xi)}}$ is locally constant in the Lagrangian Grassmannian direction.
+We may also stabilize $\xi\to\xi\oplus T^{*}\mathbb{R}^{n}$; taking $n\to\infty$, we have a sheaf of categories $\mu sh_{V^{LGr}}$ on the (stable) Lagrangian Grassmannian bundle $V^{LGr}$, locally constant along the Lagrangian Grassmannian direction.
 Let us explain how this recovers the previous notion.
 A polarization $\rho$ of the contact distribution is (by definition) a section $\rho:V\to V^{LGr(\xi)}$.
-Now, the normal bundle to $\rho(V)\subset V^{T^{*}LGr(\xi)}$ is $T^{*}LGr(\xi)$, this has the canonical polarization by the cotangent fiber.
+Now, the normal direction of a neighborhood to $\rho(V)\subset V^{T^{*}LGr(\xi)}$ can be identified with $T^{*}LGr(\xi)$.
+This carries the canonical polarization by the cotangent fiber.
 We may combine this with canonical polarization of the symplectic stable normal bundle of $V^{T^{*}LGr(\xi)}$ to obtain a polarization of the symplectic stable normal bundle of $V$.
-It’s an exercise to see that this polarization is canonically identified with $-\rho$.
+It’s an exercise to see that this polarization is canonically identified with $\rho$.
 We conclude:
@@
 We may choose the embedding of $V$ by first embedding $V^{LGr(\xi)}$.
-To obtain $\mu sh_{V,\rho}$, we must then thicken $V$ along a polarization of its normal bundle which stabilizes to $-\rho$.
+To obtain $\mu sh_{V,\rho}$, we must then thicken $V$ along a polarization of its normal bundle which stabilizes to $\rho$.
 We obtain such by thickening $V$ along $V^{LGr(\xi)}\subset V^{T^{*}LGr(\xi)}$ and then along the canonical polarization of the normal bundle to $V^{T^{*}LGr(\xi)}$.
@@
 Let $M$ be a manifold, and $\nu$ the stable normal bundle of $M$.
-Then the stable symplectic normal bundle to $S^{*}M$ is $\nu\oplus\nu^{*}$, and the stable normal polarization by $\nu$ is canonically identified with $-\phi$.
-Corollary 4.13.
-There is a canonical equivalence of categories $\mu sh_{S^{*}M}\cong\mu sh_{S^{*}M,\phi}$, where the LHS is defined by Equation 9 and the right hand side by Theorem 4.10.
+Then the stable symplectic normal bundle to $S^{*}M$ is $\nu\oplus\nu^{*}$, and the stable normal polarization by $\nu$ is canonically identified with $\phi$.
+Proposition 4.20.
+There is a canonical equivalence of categories $\mu sh_{S^{*}M}\simeq\mu sh_{S^{*}M,\phi}$, where the LHS is defined by (9) and the right hand side by Theorem 4.17.
 Proof.
@@
 Lemma 4.19 gives $\phi=-\nu$.
-Meanwhile, $i^{*}:sh(\mathbb{R}^{n})\to sh(M)$ microlocalizes to define a map $i^{*}:\sigma^{*}\mu sh_{S^{*}M(\nu)}\to\mu sh_{S^{*}M}$; using e.g. Lemma 4.9, one readily sees this is an isomorphism. ∎
+However, we’ve seen by the projectivized version of Lemma 4.9 that $i_{*}:sh(M)\to sh(\mathbb{R}^{n})$ microlocalizes to define an isomorphism $i_{*}:\mu sh_{S^{*}M}\xrightarrow{\sim}\sigma^{*}\mu sh_{S^{*}M(\nu)}$.
+∎
+Consider now a Legendrian $L\subset V$.
+Locally near $L$, we may choose a Weinstein neighborhood $D^{*}L\hookrightarrow V$, on which we may consider the polarization $\phi_{L}$ by cotangent fibers.
+Note that if $V=S^{*}M$, the polarization $\phi_{L}$ has nothing to do with the fiber polarization of $S^{*}M$.
+Lemma 4.21.
+One can uniformly fix trivializations $\mu sh_{L,\phi_{L}}\simeq loc_{L}$.
+Proof.
+We may assume without loss of generality that $V=T^{*}L$.
+To compute $\mu sh_{L,\phi_{L}}$, we note that the jet bundle $J^{1}L$ can be viewed as an open subset of the cosphere bundle $S^{*}(L\times\mathbb{R})$ by the contact embedding $J^{1}L\hookrightarrow S^{*}(L\times\mathbb{R}),(x,\xi,t)\mapsto(x,t,[\xi,1]).$ Since $\mathbb{R}$ is contractible, $\phi_{L}$ on $J^{1}L$ is the same as the restriction of $\phi_{L\times\mathbb{R}}$ on $S^{*}(L\times\mathbb{R})$ along this embedding.
+Thus, we have the identification between sheaves of categories $\mu sh_{L,\phi_{L}}=\mu sh_{L\times T^{*}_{0,>}\mathbb{R},\phi_{L\times\mathbb{R}}}$ on $S^{*}(L\times\mathbb{R})$.
+By Proposition 4.20, this realizes $\mu sh_{L,\phi_{L}}$ as $\mu sh_{S^{*}(L\times\mathbb{R});0_{L}\times T_{0}^{*}\mathbb{R}}$ in the classical sense recalled in (9).
+But the map
+$loc_{L}$ $\xrightarrow{\sim}\mu sh_{S^{*}(L\times\mathbb{R});0_{L}\times T_{0}^{*}\mathbb{R}}$
+$l$ $\mapsto l\boxtimes 1_{[0,\infty)}$
+identifies it with local systems on $L$. ∎
+In fact, we have made a choice in the above Lemma – the space of trivializations for a given $L$ is a $\mathrm{H}^{0}(L,\mathbb{Z})$ torsor.
+Our trivialization above is uniform in the sense that it is compatible with open embeddings and, in an appropriate sense, with stabilization.
+Even still, the space of uniform trivializations choices is still naturally only a $\mathbb{Z}$-torsor, of which we have chosen some particular element.
+For a discussion of this point, see [10, Remark 4.32].
+Before we leave this section, we mention common tools that are frequently used in the study of microsheaves.
+First, the notion of microsheaves is invariant under contact transform, which means the following:
+Theorem 4.22.
+Let $\mathcal{U}\subseteq S^{*}M$, $\mathcal{V}\subseteq S^{*}M$ be open subsets and $\chi:\mathcal{U}\xrightarrow{\sim}\mathcal{V}$ be contactomorphism.
 4.5. Maslov data
-It is evident from the construction that $\mu sh_{V^{LGr(\xi)}}$ is locally constant in the Lagrangian Grassmannian direction.
-We stabilize $\xi\to\xi\oplus T^{*}\mathbb{R}^{n}$; taking $n\to\infty$, we have a sheaf of categories $\mu sh_{V^{LGr}}$ on the (stable) Lagrangian Grassmannian bundle $V^{LGr}$, locally constant along the Lagrangian Grassmannian direction.
-In fact:
+Because $\mu sh_{V^{LGr}}$ is locally constant along the Lagrangian Grassmannian direction, one may expect that its descendability from $V^{LGr}$ to $V$ depends only on the ‘monodromy’ in this direction.
+Indeed this is the case, as was established in [35]; we will recall the setup here.
+Recall that for a symmetric monoidal category $\mathcal{C}$, the group of invertible objects is denoted $\operatorname{Pic}(\mathcal{C})$.
+Theorem 4.23 ([35, Sec. 11]).
+There is a map of infinite loop spaces $\mu_{\mathcal{C}}:LGr\to B\operatorname{Pic}(\mathcal{C})$ such that the sheaf of categories $\mu sh_{V^{LGr}}$ descends to the $B\operatorname{Pic}(\mathcal{C})$ bundle over $V$ classified by the map
+$V\xrightarrow{\xi}BU\to BLGr\xrightarrow{B\mu_{\mathcal{C}}}B^{2}\operatorname{Pic}(\mathcal{C})$
+Definition 4.24.
+By $\mathcal{C}$-Maslov data for $V$, we mean a null-homotopy of the map $V\to B^{2}Pic(\mathcal{C})$.
+By a $\mathcal{C}$-grading, we mean a null-homotopy of $V\xrightarrow{B\mu_{\mathcal{C}}}B^{2}\operatorname{Pic}(\mathcal{C})\to B^{2}\pi_{0}\operatorname{Pic}(\mathcal{C})$.
+We refer to the space of $\mathcal{C}$-Maslov data lifting a given $\mathcal{C}$-grading as $\mathcal{C}$-orientation data.
+In the case when $\mathcal{C}=R-mod$ for a commutative ring (spectrum) $R$, we often simplify the notation by writing $R$ in place of RmodR-mod, e.g.
+$\mu_{R}:=\mu_{R-mod}$, $\operatorname{Pic}(R)\coloneqq\operatorname{Pic}(R-mod)$, etc.
+A polarization $\rho$ provides a null-homotopy of the map $V\xrightarrow{\xi}BU\to BLGr$, so $B\mu_{C}\circ\rho$ is $\mathcal{C}$-Maslov data.
+Lemma 4.18 implies that $\mu sh_{V,\rho}=\mu sh_{V,B\mu_{C}\circ\rho}$, where the left hand side is defined as in Theorem 4.17 and the right hand side is understood in the sense above.
+For this reason, given a polarization $\rho$ we will also just write $\rho$ for the Maslov data $B\mu_{C}\circ\rho$.
+Definition 4.25.
+For $V$ a contact manifold, and a $\mathcal{C}$-Maslov datum $\eta$ for $V$, we write write $\mu sh_{V,\eta}$ for the sheaf of categories on $V$ obtained by the pullback along the zero section of the $B\operatorname{Pic}(\mathcal{C})$ bundle, trivialized by $\eta$, to which Theorem 4.23 asserts that $\mu sh_{V^{LGr}}$ descends.
+For a subset $X\subset V$ we write $\mu sh_{X,\eta}\subset\mu sh_{V,\eta}$ for the sheaf of full subcategories on objects supported in $X$.
+For an exact (real) symplectic manifold $(W,\omega=d\lambda)$, we always use implicitly the embedding in the contactization $W=W\times\{0\}\subset W\times\mathbb{R}$, and hence write $\mu sh_{W,\eta}:=\mu sh_{W\times\{0\},\eta}|_{W}$.
+Thus $\mu sh_{V,\,\cdot\,}$ is a map from the space $Mas(V)$ of Maslov data for $V$ to the category $sh(V,\mathcal{C}-cat)$ of sheaves of $\mathcal{C}$-linear categories on $V$.
+In particular, a homotopy of $\mathcal{C}$-Maslov data $h:\mu\sim\nu$ induces an equivalence of sheaves of categories $\psi(h):\mu sh_{V,\mu}\cong\mu sh_{V,\nu}$, and a homotopy $g:h_{1}\approx h_{2}$ induces an invertible natural transformation between the equivalences $\psi_{h_{1}},\psi_{h_{2}}:\mu sh_{V,\mu}\cong\mu sh_{V,\nu}$.
+Taking based loops at some Maslov datum $\eta$, we get a map $\Omega_{\eta}Mas(V)\to Aut_{C}(\mu sh_{V,\eta})$.
+Now, the space of Maslov data is a torsor for $\operatorname{Map}(V,B\operatorname{Pic}(\mathcal{C}))$; if this is nonempty, it follows that $\Omega_{\eta}Mas(V)=\operatorname{Map}(V,Pic(\mathcal{C}))$.
+It follows from the construction in [35, Definition 11.18] that the map $\Omega_{\eta}Mas(V)\to Aut(\mu sh_{V,\eta})$ is the natural map $\operatorname{Map}(V,Pic(\mathcal{C}))\to Aut_{C}(\mu sh_{V,\eta})$.
+In particular, as $\operatorname{Map}(V,Pic(\mathcal{C}))$ classifies invertible local system on $V$, any homotopy $g:h_{1}\approx h_{2}$ as above corresponds to some invertible $l(g)\in loc(V)$ and the equivalences are related by
+$\psi_{h_{2}}(-)=l(g)\otimes\psi_{h_{1}}(-):\mu sh_{V,\mu}\cong\mu sh_{V,\nu}.$
+Corollary 4.26.
+Let $V$ be a contractible contact manifold, and $\mu,\nu$ any two choices of $\mathcal{C}$-Maslov data for $V$ inducing the same $\mathcal{C}$-grading.
+Then there is an isomorphism, canonical up to non-canonical natural transformation, $\mu sh_{V,\mu}\cong\mu sh_{V,\nu}$
+Proof.
+Similar to Maslov data, the space of gradings is a torsor for $\operatorname{Map}\left(V,\tau_{\leq 1}B\operatorname{Pic}(\mathcal{C})\right)$; here the hypothesis of ‘inducing the same $C$ grading’ should be understood as meaning that we are given a choice of path between the gradings associated to the given Maslov data.
+Since $V$ is assumed contractible, we may lift the path to a path of Maslov data and obtain the desired isomorphism.
+Two different paths differ by an loop which is trivial in $\operatorname{Map}(V,\pi_{0}(Pic(\mathcal{C})))$ hence by an automorphism which is the identity in $\pi_{0}(Aut_{C}(\mu sh_{V,\eta}))$.
+∎
+In the discussion thus far, we have been agnostic as far as the choice of the category $\mathcal{C}$, and we have also not needed to compute the map $\mu_{\mathcal{C}}:U/O\to B\operatorname{Pic}(\mathcal{C})$.
+We now turn to this question.
+Note first that given a map of symmetric monoidal stable categories $\mathcal{C}\to\mathcal{D}$, it follows from the construction that $\mu_{\mathcal{D}}$ is the composition of $\mu_{\mathcal{C}}$ with the natural map $Pic(\mathcal{C})\to Pic(\mathcal{D})$.
+In particular, when $R$ is a discrete commutative ring ($R=\pi_{0}(R)$), the map $\mu_{R-mod}$ factors through $\mu_{\mathbb{Z}}$.
+The map $\tau_{\leq 0}\Omega\mu_{\mathbb{Z}}:\Omega LGr\to\mathbb{Z}$ was shown to be the Maslov index by Kashiwara and Schapira [25, Thm.
+7.5.11], and was later fully characterized by Guillermou [14].
+More generally, any symmetric monoidal stable category $C$ admits a symmetric monoidal functor from the category of spectra (aka modules over the sphere spectrum $\mathbb{S}$).
+The map $\mu_{\mathbb{S}}:\Omega LGr\to\operatorname{Pic}(\mathbb{S})$ was shown by Jin [18] to agree with the J-homomorphism.
+By truncation one recovers Guillermou’s result in a more convenient (for us) formulation.
 Theorem 4.27.
-[22, Sec. 11] There is a map of infinite loop spaces $\mu:LGr\to BPic(C)$ such that the sheaf of categories $\mu sh_{V^{LGr}}$ descends to the $BPic(C)$ bundle over $V$ classified by the map
+[14, 18] The map $\Omega\mu_{\mathbb{Z}}:\Omega LGr\to\operatorname{Pic}(\mathbb{Z})$ is the following composition:
+$\Omega LGr\to\tau_{\leq 1}(\Omega LGr)\xrightarrow{\tau_{\leq 1}(J)}\tau_{\leq 1}\operatorname{Pic}(\mathbb{S})=\operatorname{Pic}(\mathbb{Z}).$
+It follows in particular that $\mathbb{Z}$ Maslov data is precisely grading/orientation data, and hence that grading/orientation data provide $R$ Maslov data for any commutative ring $R$ (although not all $R$ Maslov data need arise in this way).
+4.6. Secondary Maslov data
+If $L\subset V$ is a Lagrangian (resp.
+Legendrian) in a symplectic (resp.
+contact) manifold, as in Definition 3.11, the Weinstein neighborhood theorem provides a polarization $\phi_{L}$ near $L$.
+Assume further that $V$ is equipped with a Maslov datum $\eta$.
+Recall from [35] that we say a homotopy $\phi_{L}\sim\eta|_{L}$ is a secondary Maslov datum for $L$.
+As noted in [35, Remark 11.20], a choice of secondary Maslov data identifies microsheaves on $L$ with local systems.
+Indeed, a secondary Maslov datum induces an equivalence $\mu sh_{L,\eta}\simeq\mu sh_{L,\phi_{L}}$ which we can further compose with the equivalence $\mu sh_{L,\phi_{L}}\simeq loc_{L}$ from Lemma 4.21.
+Taking stalks at a smooth point $p\in L$, we obtain:
+Corollary 4.28.
+Fix a contact manifold (resp.
+exact symplectic) $V$ and a Legendrian (resp.
+conical Lagrangian) $L\subset V$, and $p\in L$.
+Fix a Maslov datum on $V$ and a secondary Maslov datum on $L$.
+Then is an equivalence
+(16) $\mathcal{C}\xrightarrow{\sim}(\mu sh_{L})|_{p}.$
+Any choice of such an isomorphism is termed a microstalk functor.
+The microstalk functor (16) depends on the choice of secondary Maslov datum $\eta\sim\phi_{L}$, and the ambiguity is a torsor for $\operatorname{Pic}(\mathcal{C})$.
+However, if $L$ is endowed with the additional datum of a secondary $\mathcal{C}$-grading, we can cut down the ambiguity to a smaller group.
+To explain this, let $\eta,\eta^{\prime}$ be $\mathcal{C}$-Maslov data on a contact manifold $V$ (henceforth we restrict our attention to the contact case, leaving the symplectic analogue to the reader).
+As noted in the proof of Proposition 3.10, a homotopy of Maslov data $\eta\sim\eta^{\prime}$ is the same thing as null-homotopy of the difference $\eta,\eta^{\prime}]:V\to\Omega B^{2}Pic(\mathcal{C})$.
+Similarly, if $\eta,\eta^{\prime}$ are $\mathcal{C}$-gradings (Definition 4.24), then a homotopy of $\mathcal{C}$-gradings is a homotopy of the corresponding map $V\to\Omega B^{2}\pi_{0}\operatorname{Pic}(\mathcal{C})$.
+Suppose now that $L\subset V$ is a Legendrian, and let $\phi_{L}$ denote the (Maslov datum induced by the) canonical fiber polarization near $L$.
+Consider the difference $\eta,\phi_{L}]:Op(L)\to\Omega B^{2}\operatorname{Pic}(\mathcal{C})$ and we let $\overline{[\eta,\phi_{L}]}:Op(L)\to\Omega B^{2}\pi_{0}\operatorname{Pic}(\mathcal{C})$ denote the composition of $\eta,\phi_{L}]$ with the natural map $\Omega B^{2}Pic(\mathcal{C})\to\Omega B^{2}\pi_{0}Pic(\mathcal{C})$.
+A secondary $\mathcal{C}$-grading $h$ shall mean a null-homotopy of $\overline{[\eta,\phi_{L}]}.$
+Definition 4.29.
+Given $V,L,h$ as above, a secondary $\mathcal{C}$-orientation datum for $L$ is a is a null-homotopy of $\eta,\phi_{L}]:Op(L)\to\Omega B^{2}\pi_{0}\operatorname{Pic}(\mathcal{C})$ lifting the null-homotopy $h$.
+The difference between any secondary $\mathcal{C}$-orientation data lifting $h$ is a map $V\to\Omega^{2}B^{2}\operatorname{Pic}(\mathcal{C})=\operatorname{Pic}(\mathcal{C})$ such that the composition $V\to\Omega^{2}B^{2}\operatorname{Pic}(\mathcal{C})\to\Omega^{2}B^{2}\pi_{0}\operatorname{Pic}(\mathcal{C})$ is null, i.e.
+a map $V\to\Omega^{2}B^{2}\operatorname{Pic}_{0}(\mathcal{C})=\operatorname{Pic}_{0}(\mathcal{C})$.
+Corollary 4.30.
+In the situation of Corollary 4.28, assume given a secondary $\mathcal{C}$-grading $h$ on $L$.
+Given two choices of secondary orientation data lifting $h$, the corresponding microstalk functors (16) are a torsor for $\operatorname{Pic}_{0}(\mathcal{C})$.
+Since the space of secondary orientation data is certainly nonempty when $L$ is contractible, (16) is canonical up to an action of $\operatorname{Pic}_{0}(\mathcal{C})$ by natural transformation.
+In particular, (16) is unambiguous at the level of objects.
+4.7. Constrained Maslov data
+In this subsection, we define a notion of constrained Maslov data.
+The purpose of the discussion here is to have a framework for perverse t-structures that come from exotic t-structures on the coefficient category $\mathcal{C}$ that are different from the standard one one RmodR-mod as considered in Definition 6.16.
+None of this material is needed for proving the main results of this paper stated in the introduction and readers who care primarily about the canonical t-structure can safely skip it.
+Let $\{\mathcal{D}_{i}\}$ be a collection of subcategories $\mathcal{D}_{i}\subseteq\mathcal{C}$. One denote by
+$\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}\coloneqq\{x\in\operatorname{Pic}(\mathcal{C})|x\otimes y_{i}\in D_{i},\ \forall y_{i}\in D_{i}\forall i\}$
+the submonoid of $\operatorname{Pic}(\mathcal{C})$ which preserves each $\mathcal{D}_{i}$.
+When the collection consists of only one subcategory $\mathcal{D}$, we simplify the notation and denote it by $\operatorname{Pic}(\mathcal{C})_{\mathcal{D}}$.
+Beware that in general, elements of $\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$ need not induce an automorphism on each $\mathcal{D}_{i}$.
+Example 4.31.
+Suppose $\mathcal{C}=R-mod$ for a discrete ring $R$ and $\mathcal{C}^{\geq 0}$ is the subcategory of objects supported in non-negative degrees.
+Then tensoring with $R[-1]\otimes M=M[-1]$ has the effect of shifting the cohomology degree up by one so $R[-1]\in\operatorname{Pic}(R)_{\mathcal{C}^{\geq 0}}$.
+But clearly, its inverse $R[1]\not\in\operatorname{Pic}(R)_{\mathcal{C}^{\geq 0}}$ so $\operatorname{Pic}(R)_{\mathcal{C}^{\geq 0}}$ is only a monoid instead of a group.
+Definition 4.32.
+We say $\{\mathcal{D}_{i}\}$ is an anchored collection if $\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$ is a subgroup.
+Our reason for considering anchored collections is that the situation encountered in Example 4.31 (where the functor induced by tensoring with R[1]R[-1] fixes $\{\mathcal{C}^{\geq 0}\}$ but is not surjective) cannot happen:
+Lemma 4.33.
+Let $\{\mathcal{D}_{i}\}$ be an anchored collection.
+Then, for any $x\in\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$, the functor $x\otimes(-):\mathcal{C}\xrightarrow{\sim}\mathcal{C}$ restricts to an equivalence $x\otimes(-):\mathcal{D}_{i}\xhookrightarrow{\sim}\mathcal{D}_{i}$.
+Proof.
+Fully-faithfulness is automatic so we show that if $x\in\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$, then for any $a\in\mathcal{D}_{i}$ there exists $b\in\mathcal{D}_{i}$ such that $x\otimes b=a$ .
+But we know $b$ exists in $\mathcal{C}$, and it must satisfy $b=x^{-1}\otimes a$.
+The fact that $b\in\mathcal{D}_{i}$ then follows from the anchored assumption since $x^{-1}$ is also in $\operatorname{Pic}(\mathcal{C})_{\mathcal{D}_{i}}$.
+∎
+From now on, we always assume $\{\mathcal{D}_{i}\}$ to be anchored.
+Clearly, the identity component $\operatorname{Pic}(\mathcal{C})_{0}\subseteq\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$ is contained in any such subgroup.
+Thus, the cofiber $\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}\rightarrow\operatorname{Pic}(\mathcal{C})$ is a discrete group, as it is a quotient of $\pi_{0}(\operatorname{Pic}(\mathcal{C}))$.
+We denote it by $\pi_{0}\left(\operatorname{Pic}(\mathcal{C});\{\mathcal{D}_{i}\}\right)$.
+Definition 4.34.
+A $(\mathcal{C},\{\mathcal{D}_{i}\})$-grading is a null-homotopy of the map
+$V\rightarrow B^{2}\operatorname{Pic}(\mathcal{C})\rightarrow\pi_{0}\left(\operatorname{Pic}(\mathcal{C});\{\mathcal{D}_{i}\}\right).$
+A $(\mathcal{C},\{\mathcal{D}_{i}\})$-orientation of a given $(\mathcal{C},\{\mathcal{D}_{i}\})$-grading is a lift to a $\mathcal{C}$-Maslov datum.
+Example 4.35.
+Let $\mathcal{C}=\operatorname{Coh}(\mathbb{P}^{1})$ be the category of coherent sheaves on $\mathbb{P}^{1}$.
+In this case, invertible objects are of the form $\mathcal{O}(n)[m]$ for $n,m\in\mathbb{Z}$, so $\pi_{0}(\operatorname{Pic}(\mathbb{P}^{1}))=\mathbb{Z}\times\mathbb{Z}$.
+Let $\mathcal{D}_{1}$ be the full subcategory whose objects are honest coherent sheaves (viewed as complexes supported in degree zero).
+Then the group $\operatorname{Pic}(\mathbb{P}^{1})_{\{\mathcal{D}_{1}\}}$ contains only $\mathcal{O}(n)$ for $n\in\mathbb{Z}$.
+Thus $\pi_{0}\left(\operatorname{Pic}(\mathcal{C});\{\mathcal{D}_{1}\}\right)=\mathbb{Z}$ only remembers the homological degree shift.
+As illustrated in the diagram below, a usual $C$-grading induces a $(\mathcal{C},\{\mathcal{D}_{i}\})$-grading, for which a $\mathcal{C}$-orientation induces a $(\mathcal{C},\{\mathcal{D}_{i}\})$-orientation.
+However, the latter is more general.
 […]
+We can give “constrained” analogues of the constructions in Section 4.6.
+Namely, suppose now that $L\subset V$ is a Legendrian.
+If $\{\mathcal{D}_{i}\}$ is a collection of subcategories, as in Definition 4.34, a secondary orientation on $L$ constrained by $\{\mathcal{D}_{i}\}$ is a homotopy between the polarization $(\mathcal{C},\{\mathcal{D}_{i}\})$-grading and the $(\mathcal{C},\{\mathcal{D}_{i}\})$-grading induced from $\eta$.
+A secondary Maslov datum for $L$ lifting a given secondary polarization constrained by $\{\mathcal{D}_{i}\}$ shall be called a secondary orientation datum for $L$ constrained by $\{\mathcal{D}_{i}\}$.
+Corollary 4.36 (cf. Corollary 4.30).
+Let $\eta$ be a Maslov datum and $L$ be an Legendrian endowed with a secondary grading constrained by $\{\mathcal{D}_{i}\}$.
+Then the ambiguity of the equivalence $\mathcal{C}\cong(\mu sh_{L,\eta})_{p}$ from Corollary 4.28 can be reduced to $\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$.
+In particular, per Corollary 4.36, the statement that “the microstalk of $\mathcal{F}\in(\mu sh_{L,\eta})|_{p}$ is contained in $\mathcal{D}_{i}$” is well-defined.
+Example 4.37.
+Suppose that $(\mathcal{C},\mathcal{D})$ are as in Example 4.35.
+Then the microstalk of $\mathcal{F}\in(\mu sh_{L,\eta})|_{p}$ is an object in $\operatorname{Coh}(\mathbb{P}^{1})$ which is well-defined up to tensoring with $\mathcal{O}(n)$.
+Hence, it is meaningful to ask whether the microstalk of $\mathcal{F}$ belongs to the heart $\operatorname{Coh}(\mathbb{P}^{1})^{\heartsuit}$ (i.e.
+whether it is represented by an honest coherent sheaf).
+5. Microsheaves in the complex setting
+5.1. Microsheaves on complex cotangent bundles
+We now review the results of Waschkies [42].
+Denote by $\pi:T^{\circ}M\rightarrow\mathbb{P}^{*}M$ the projection.
+The perverse microsheaves on $\mathbb{P}^{*}M$ [42, Definition 6.1.2] is defined as a subsheaf of the following sheaf.
 Definition 5.1.
-By $C$-Maslov data for $V$, we mean a null-homotopy of the map $V\to B^{2}Pic(C)$.
-A choice $\eta$ of $C$-Maslov data gives a trivialization of the $BPic(C)$ bundle to which Theorem 4.14 asserts that $\mu sh_{V^{LGr}}$ descends.
-We write $\mu sh_{V,\eta}$ for the sheaf of categories on $V$ obtained by the pullback along the zero section of this trivial bundle (of groups).
+We define the presheaf $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}^{pre}$ on $\mathbb{P}^{*}M$ by
+$\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}^{pre}(\Omega)\coloneqq\mu sh^{pre}(\pi^{-1}(\Omega))$
+for an open set $\Omega\subseteq\mathbb{P}^{*}M$, and denote its sheafification by $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}$.
+If $\Lambda\subseteq\Omega$ is a closed complex Legendrian, we can similarly define a presheaf on $\Omega$, using the notation in (13), by
+$\operatorname{\mathbb{P}\mu sh}_{\Lambda}^{pre}(\Omega^{\prime})\coloneqq\mu sh_{\pi^{-1}(\Lambda)}^{pre}(\pi^{-1}(\Omega)),$
+and denote its sheafification by $\operatorname{\mathbb{P}\mu sh}_{\Lambda}$.
+Per Lemma 4.8, the canonical map $\operatorname{\mathbb{P}\mu sh}_{\Lambda}\hookrightarrow\operatorname{\mathbb{P}\mu sh}$ is fully faithful and its image consists of those objects supported in $\Lambda$.
+Lastly, we define $\operatorname{\mathbb{P}\mu sh}_{\mathbb{C}-c}$ to be the subsheaf consisting of objects supported on complex Legendrians, which can also be obtained as the shaefification of $\operatorname{\mathbb{P}\mu sh}_{\mathbb{C}-c}^{pre}(\Omega)=\mu sh_{\mathbb{C}-c}^{pre}(\pi^{-1}(\Omega))$
+The following Theorem 5.3 is one of the main theorems in [42], which gives a simple description of $\operatorname{\mathbb{P}\mu sh}_{\Lambda}$ at stalks when $\Lambda$ is in generic position:
+Definition 5.2.
+Suppose that $M$ is complex analytic and $\Lambda\subset P^{*}M$ is a (singular) complex Legendrian.
+Let $r:P^{*}M\to M$ be the projection.
+We say that $\Lambda$ is in generic position if that $m\in M$ has the property that $\Lambda\cap r^{-1}(m)=\{p_{1},\dots,p_{k}\}$ is a finite set.
+Theorem 5.3 ([42, Thm. 5.1.5]).
+In the situation of Definition 5.2, let $p\in r^{-1}(m)\cap\Lambda$. There is a fully faithful functor
+$(\operatorname{\mathbb{P}\mu sh}_{\Lambda})_{p}=\mu sh^{pre}_{\pi^{-1}(\Lambda)}(\mathbb{C}^{*}\cdot p)\to(sh_{\mathbb{C}-c}/loc)_{m}=(sh_{\mathbb{C}-c})_{m}/loc_{m}.$
+Its essential image is the full subcategory of $(sh_{\mathbb{C}-c}/loc)_{m}$ on objects whose microsupport is contained in $\Lambda\cap Op(\mathbb{C}^{*}\cdot p)$ for some small open neighborhood $Op(\mathbb{C}^{*}\cdot p)$.
+Proof.
+The fully faithful functor is furnished by Waschkies [42, Thm.
+5.1.5] using the microlocal cutoff.
+Let us temporarily denote by $\mathcal{A}$ the full subcategory of $(sh_{\mathbb{C}-c}/loc)_{m}$ on objects whose microsupport is entirely contained in $\pi^{-1}(\Lambda)\cap Op(\mathbb{C}^{*}\cdot p)$.
+That Waschkies’ functor lands inside $\mathcal{A}$ is a consequence of [42, Thm.
+5.1.5(3)].
+We should prove that every $F\in\mathcal{A}$ lies in the image of Wacshkies’ functor.
+We may assume $M$ is a ball, so that we have coordinates $T^{\circ}M=M\times T^{\circ}_{m}M$.
+By assumption, there is some neighborhood $U\subset M$ of $m$ such that $SS(F)\cap\dot{\pi}^{-1}(U)\subset U\times\gamma$, where $\gamma$ is a neighborhood of $\mathbb{C}^{*}\cdot p$.
+Then by construction, Waschkies’ map sends (the object in $\mu sh^{pre}_{\Lambda}(\mathbb{C}^{*}\cdot p)$ represented by) $F$ to $\Phi_{U,\gamma}(\mathcal{F})$, where $\Phi_{U,\gamma}(-)$ is the microlocal cutoff.
+There is always a map $\alpha:\Phi_{U}(F)\to F$, and we need to show that the cone is a local system on $U$.
+Up to shrinking $M$, assume $U=M$.
+Then $cone(\alpha)$ has no microsupport in $\dot{\pi}^{-1}(M)\setminus(\gamma\times V)$ (because neither $F$ nor $\Phi_{U,\gamma}F$ have microsupport there).
+But $cone(\alpha)$ also has no microsupport in $U$ because $\Phi_{U,\gamma}$ induces an isomorphism in $\mu sh^{pre}(\gamma\times M)$ [42, Def.
+2.3.1(iii)].
+Hence the microsupport of $cone(\alpha)$ is contained in the zero section, as claimed.
+∎
+Remark 5.4.
+The analogue of Theorem 5.3 with complex analyticity hypotheses removed and with $\mathbb{C}^{*}$ replaced by $\mathbb{R}^{+}$ is also true (indeed easier: the cutoffs already in [25] are good enough and one does not need [8]); see [35, Lem 6.7, Prop.
+6.9].
+Now, we compare the two notions of microsheaves $\mu sh_{T^{*}M}$ and $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}$. The canonical map
+$\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}^{pre}(\Omega)\coloneqq\mu sh^{pre}(\pi^{-1}(\Omega))\rightarrow\mu sh(\pi^{-1}(\Omega))$
+defines a map $\operatorname{\mathbb{P}\mu sh}_{P^{*}M}\rightarrow\pi_{*}(\mu sh|_{T^{\circ}M})$.
+Lemma 5.5.
+The map $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}\hookrightarrow\pi_{*}\left(\mu sh|_{T^{\circ}M}\right)$ is fully-faithful.
+Proof.
+It’s enough to check on stalks.
+Take $p\in\mathbb{P}^{*}M$, the map $(\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M})_{p}\rightarrow\left[\pi_{*}\left(\mu sh|_{T^{\circ}M}\right)\right]_{p}$ is given by
+$\mu sh^{pre}(\pi^{-1}(p))\rightarrow\mu sh(\pi^{-1}(p)).$
+As mentioned in (10), the Hom sheaf on the right-hand side $\mathcal{H}om_{\mu sh}$ is computed by $\mu hom$.
+In particular, $\operatorname{Hom}_{\mu sh(\pi^{-1}(p))}$ is computed by $\Gamma(\pi^{-1}(p);\mu hom(-,-)).$ But [42, Proposition 2.4.4] shows that it is also the case for the category on the left-hand side so we are done.
+∎
+Since $\mu sh^{pre}(\pi^{-1}(p))=sh(M)/sh_{T^{*}M\setminus\pi^{-1}(p)}(M)$ is a quotient, objects in $\mu sh^{pre}(\pi^{-1}(p))$ are presented by sheaves.
+Thus, the proof of the above lemma shows characterizes the image as the following:
+Corollary 5.6.
+The subsheaf $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}$ in $\pi_{*}(\mu sh_{T^{\circ}M})$ is equivalent to the subsheaf stalkwisely on $P^{*}M$ presented by a sheaf
+$\{\mathcal{F}\in\pi_{*}(\mu sh|_{T^{\circ}M})|\mathcal{F}|_{p}\in\operatorname{im}(sh(M)\rightarrow\mu sh(\pi^{-1}(p)),\forall p\in\mathbb{P}^{*}M\}.$
+Similar to the real situation Theorem 4.10, there is a complex version of the contact transform.
+Theorem 5.7 ([25, (11.4.8)]).
+Let $\mathcal{U}\subseteq\mathbb{P}^{*}M$, and $\mathcal{V}\subseteq\mathbb{P}^{*}N$ be open sets, and $\chi:\mathcal{U}\xrightarrow{\mathcal{V}}$ be a complex contactomorphism.
+Then, for any given $p\in\mathcal{U}$, shrink $\mathcal{U}$ if needed, one can assume that there exists a sheaf $K\in sh(M\times N)$ such that the functor $\Phi_{K}:sh(M)\rightarrow sh(N)$ given by convolving with $K$ induces an equivalence, often referred as contact transformation,
 […]
-For a subset $X\subset V$ we write $\mu sh_{X,\eta}\subset\mu sh_{V,\eta}$ for the sheaf of full subcategories on objects supported in $X$.
-For any smooth Legendrian $L\subset V$, Theorem 4.14 plus the Weinstein neighborhood theorem imply $\mu sh_{L,\eta}\subset\mu sh_{V,\eta}$ depends only on $\eta|_{L}$ (and not on $V$).
+where $\widetilde{\mathcal{U}}$, $\widetilde{\mathcal{V}}$ and $\tilde{\chi}$ are the corresponding symplectic lifts.
+Consequently, it induces an equivalence $\mu sh_{T^{\circ}M}|_{\mathcal{U}}\xrightarrow{\sim}\tilde{\chi}^{*}\mu sh_{T^{\circ}N}|_{\mathcal{V}}$ which commutes with the canonical map $\mu sh^{pre}\rightarrow\mu sh$.
+Corollary 5.8.
+With the notation as above, $\Phi_{K}$ induces an equivalence ${\pi_{M}}_{*}\mu sh_{\widetilde{\mathcal{U}}}\xrightarrow{\sim}\chi^{*}{\pi_{N}}_{*}\mu sh_{\widetilde{\mathcal{V}}}$ which restricts to $\operatorname{\mathbb{P}\mu sh}_{\mathcal{U}}\xrightarrow{\sim}\chi^{*}\operatorname{\mathbb{P}\mu sh}_{\mathcal{V}}$.
+Proof.
+The first equivalence is tautological.
+The second equivalence follows from the characterization of $\operatorname{\mathbb{P}\mu sh}$ in Corollary 5.6 as locally presentable by sheaves and the fact that contact transformation commutes with the canonical map $\mu sh^{pre}\rightarrow\mu sh$ as mentioned in Theorem 5.7.
+∎
+In general, we do not know if the inclusion $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}\hookrightarrow\pi_{*}(\mu sh_{T^{\circ}M})$ is an equivalence.
+However, it is the case when we restrict to complex constructible objects.
+Proposition 5.9.
+The inclusion $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M,\mathbb{C}-c}\xhookrightarrow{\sim}\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})$ is an equivalence.
+Proof.
+Let $p\in\mathbb{P}^{*}M$ and consider $\mathcal{F}\in[\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})]_{p}=\mu sh_{T^{\circ}M,\mathbb{C}-c}(\pi^{-1}(p))$.
+Since objects in $\mu sh_{T^{\circ}M,\mathbb{C}-c}(\pi^{-1}(p))$ are germs of microsheaves near $\pi^{-1}(p)$, we can pick some $\Omega\subseteq\mathbb{P}^{*}M$ containing $p$ and realize $\mathcal{F}$ as a microsheaf on $\pi^{-1}(\Omega)$ with $ss(F)\subseteq\pi^{-1}(\Omega)$ being a complex Lagrangian.
+Denote by $\Lambda\subseteq\Omega$ the corresponding complex Legendrian.
+By the previous Corollary 5.8, being in $\operatorname{\mathbb{P}\mu sh}$ is invariant under contact transform, so we can apply the Kashiwara–Kawai general position theorem (see [23, Sec.
+1.6]) and assume that the composition $\Lambda\subseteq\mathbb{P}^{*}M\xrightarrow{r}M$ is finite to one near $p$.
+Shrinking $\Omega$ and $M$ if needed, we may assume that $\Lambda$ admits the standard form [26, (2.5)]: There exists local coordinates $(z,\xi)$ such that $p=dz_{n}$ and $\Lambda$ is of the form $\mathbb{P}^{*}_{S}M$, where $S=\{f=0\}$ is the zero locus of some holomorphic function $f=z_{n}^{k}+g(z)$ for some $k\in\mathbb{N}$ and for some $g(z)\in(z_{1},\cdots,z_{n})^{k+1}$.
+But this implies that $\mathcal{F}\in\mu sh_{\mathbb{P}_{S}^{*}M}(\Omega)$ so we can apply Theorem 4.11 and conclude that there exists an $F\in sh(M)$ which projects to $\mathcal{F}$, which in particular implies that $\mathcal{F}\in\operatorname{\mathbb{P}\mu sh}_{P^{*}M,\mathbb{C}-c}$.
+∎
+5.2. Canonical microsheaves and microstalks
+We return to considering an exact symplectic or contact manifold $W$ with corresponding structure morphism $W\rightarrow BU$.
+Suppose given moreover a lift to stable quaternionic structure $W\rightarrow BSp$, e.g.
+arising from an underlying complex symplectic or contact structure as in Section 2.
+Per Definition 3.8, $W$ carries a canonical orientation/grading datum.
+Per Theorem 4.27, this determines a canonical $R$-Maslov datum, for $R$ a commutative ring $R$.
+Definition 4.25 thus furnishes a sheaf $\mu sh_{W}$ of $R$-linear categories on $W$.
+More generally, fix a coefficient category $\mathcal{C}$ and an anchored collection of subcategories $\{\mathcal{D}_{i}\}$.
+Endow $W$ with the $(\mathcal{C},\{\mathcal{D}_{i}\})$ grading induced by the canonical $\mathbb{S}$-grading (see Definition 3.8 and (17)), and sssume given a $(\mathcal{C},\{\mathcal{D}_{i}\})$ orientation $o$.
+Then Definition 4.25 furnishes a sheaf of $\mathcal{C}$-linear categories on $W$ which we denote by $\mu sh_{W,o}$.
+Proof of Theorem 1.1.
+Let us first discuss the case $R=\mathbb{Z}$.
+We take the canonical $\mathbb{Z}$-Maslov datum.
+Assume $L$ is a complex Lagrangian or Legendrian.
+Then, the polarization $\phi_{L}:L\rightarrow BO$ admits a lift to a complex polarization $L\rightarrow BU\rightarrow BO$.
+To compare it to the fiber polarization $\phi_{L}$ near L, we consider the secondary Maslov data.
+By Theorem 4.27 secondary Maslov data are in one-to-one correspondence with secondary orientation/grading.
+The theorem thus reduces to Lemma 3.13.
+For general $R$, note that Lemma 3.13 is itself deduced from Proposition 3.10.
+Tracing the argument there, the only modification needed is to further compose $B^{2}\mathbb{Z}\rightarrow B^{2}(\mathbb{Z}/2\mathbb{Z})$ with $B^{2}(\mathbb{Z}/2\mathbb{Z})\rightarrow B^{2}R^{\times}$ (the latter map is the twice delooping of $\mathbb{Z}/2\mathbb{Z}=\mathbb{Z}^{\times}\rightarrow R^{\times}$).
+∎
+By Theorem 1.1, $R$-secondary Maslov data for the canonical Maslov data are given by $R$-spin structures.
+Proof of Corollary 1.2.
+From Theorem 1.1 we have a canonical Maslov datum for $W$, giving the sheaf of stable categories $\mu sh_{W}$ via Definition 4.25.
+Also from Theorem 1.1, an $R$-spin structure $\sigma$ on $L$ determines a secondary Maslov datum which, per Lemma 4.21, determines an equivalence $\mu sh_{L,\eta}\simeq\mu sh_{L,\phi_{L}}\simeq loc_{L}$.
+∎
+Proof of Corollary 1.3.
+Without loss of generality, we can assume that $X=L$ is a (conical) complex Lagrangian disk.
+Endow $W$ with the canonical $R$-Maslov datum.
+According to Lemma 3.12, there is a canonical choice of secondary grading.
+So any choice of secondary orientation induces a secondary Maslov datum, and hence, by Corollary 4.28, a microstalk functor $\omega_{p}^{-1}:(\mu sh_{L,o})_{p}\to R-mod$.
+By Corollary 4.28, $\omega_{p}^{-1}$ acts unambiguously on the set of objects.
+∎
+We can also generalized Corollary 1.3 to the “constrained” setting (see Section 4.7).
+Namely, assume that $W$ is an exact complex symplectic manifold and let $L\subset W$ be a (conical, smooth) complex Lagrangian.
+Fix a coefficient category $\mathcal{C}$ and an anchored collection of subcategories $\{\mathcal{D}_{i}\}$.
+Endow $W$ with the canonical $(\mathcal{C},\{\mathcal{D}_{i}\})$ grading and assume given a $(\mathcal{C},\{\mathcal{D}_{i}\})$ orientation $o$.
+According to Lemma 3.12, there is a canonical choice of secondary grading.
+So any choice of secondary orientation induces a secondary Maslov datum, and hence, by Corollary 4.28, a microstalk functor $\omega_{p}^{-1}:(\mu sh_{L,o})_{p}\to\mathcal{C}$.
+By Corollary 4.36, the ambiguity of $\omega_{p}^{-1}$ is an element of of $\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$.
+Hence, given one of the $\mathcal{D}_{i}$, it is meaningful to ask for the image of $\omega_{p}^{-1}$ to be contained in $\mathcal{D}_{i}$.
+(If $\{\mathcal{D}_{i}\}$ is the total partition of $\mathcal{C}$ into $1$-object subcategories), $\omega_{p}^{-1}:(\mu sh_{L,o})_{p}\to R-mod$ is well-defined as an object, so we recover Corollary 1.3.)
 Remark 5.10.
-There is a canonical choice of Maslov data on (or in a neighborhood of) a real Legendrian $L$.
-Indeed, canonically $\xi|_{L}\cong T^{*}L$, and taking the fiber polarization $\phi_{L}$ gives a canonical null-homotopy of $L\to V\to BLGr$.
-By composition we obtain a canonical null-homotopy $\phi_{L}$ of $L\to B^{2}Pic(C)$.
-Given Maslov data $\eta$ for $V$, we obtain another such null-homotopy $\eta|_{L}$, we say that a homotopy $\phi_{L}\sim\eta|_{L}$ is secondary Maslov data for $L$.
-Evidently, secondary Maslov data induces an isomorphism $\mu sh_{L,\eta|_{L}}\cong\mu sh_{L,\phi_{L}}$.
-It is possible to further fix an isomorphism $\mu sh_{L,\phi_{L}}\cong Loc_{L}$, the right hand side being the sheaf of categories of local systems on $L$.
-To do so, consider the embedding $L=L\times 0\subset L\times\mathbb{R}$.
-Let $L^{+}\subset S^{*}(L\times\mathbb{R})$ be the positive conormal to $L$.
-Observe that the fiber polarization of $S^{*}(L\times\mathbb{R})$ restricts to $\phi_{L^{+}}$.
-Taking the pushforward along the inclusion and then microlocalizing along the positive conormal gives a map $Loc_{L}\to\mu sh_{L}\cong\mu sh_{L,\phi_{L}}$, which we can check locally is an equivalence (e.g.
-using [15, Proposition 6.6.1]).
-Remark 4.17.
-The category of co-orientable contact manifolds is equivalent to the category of symplectic manifolds with a free transitive $\mathbb{R}_{>0}$ action scaling the symplectic form: given a contact manifold $V$, we may form the corresponding symplectic manifold $\mathbb{R}_{>0}V$ by taking the locus in $T^{*}V$ of all graphs of contact 1-forms respecting the co-orientation.
-An embedding $V\hookrightarrow S^{*}M$ always lifts to a $\mathbb{R}_{>0}$-equivariant embedding $\mathbb{R}_{>0}V\to T^{\circ}M=\mathbb{R}_{>0}S^{*}M$, uniquely up to shift by a function $f:V\to\mathbb{R}_{>0}$, and a thickening of $V$ canonically lifts to a thickening of $\mathbb{R}_{>0}V$, etc.
-Stable polarizations and Maslov data being topological, these notions are equivalent on $V$ and $\mathbb{R}_{>0}V$.
-Thus by embedding, thickening, descending, etc., we find that from Maslov data $\eta$ on $V$, we obtain an $\mathbb{R}_{>0}$-equivariant sheaf of categories $\mu sh_{\mathbb{R}_{>0}V,\eta}$ on $\mathbb{R}_{>0}V$.
-It is of course canonically isomorphic to the pullback of $\mu sh_{V,\eta}$ along the projection $\mathbb{R}_{>0}V\to V$.
-In the discussion thus far, we have been agnostic as far as the choice of the category $C$, and we have also not needed to compute the map $\mu_{C}:U/O\to BPic(C)$.
-We now turn to this question.
-Note first that given a map of symmetric monoidal stable categories $C\to D$, it is clear that $\mu_{D}$ is the composition of $\mu_{C}$ with the natural map $Pic(C)\to Pic(D)$.
-In particular, when $R$ is an ordinary commutative ring ($R=\pi_{0}(R)$), the map $\mu_{R-mod}$ factors through $\mu_{\mathbb{Z}-mod}$.
-We have $Pic(\mathbb{Z}-mod)=\mathbb{Z}\oplus B(\mathbb{Z}/2\mathbb{Z})$, where $\mathbb{Z}/2\mathbb{Z}$ is really the group of invertible integers, {±$\{\pm 1\}$.
-The map $\mu_{\mathbb{Z}-mod}$ has been calculated:
-Theorem 4.18.
-[8] The map $\mu_{\mathbb{Z}-mod}$ of Theorem 4.14 is $LGr\xrightarrow{\det^{2}\oplus w_{2}}S^{1}\oplus B^{2}(\mathbb{Z}/2\mathbb{Z})$.
-That is, $\mathbb{Z}-mod$ Maslov data is precisely grading and orientation data.
-It follows that grading and orientation data provide $R-mod$ Maslov data for any commutative ring $R$, although of course there may be $R-mod$ Maslov data which does not arise in this way.
-Remark 4.19.
-Let $C=\mathbb{Z}-mod$.
-Let $M$ be a manifold.
-Let $m_{can}$ be the Maslov datum whose grading comes from the fiber polarization, and whose orientation data is the canonical orientation data of Def.
-3.8.
-By Lemma 3.9, Corollary 4.13, and Theorem 4.18, a spin structure on $M$ gives an equivalence of sheaves of categories on $T^{\circ}M$: $\mu sh_{S^{*}M}\cong\mu sh_{S^{*}M,m_{can}}$ where the right hand side is defined by Equation (9) and the LHS by Theorem 4.14.
-If $M=N\times\mathbb{R}$ is not spin, then in fact $\mu sh_{S^{*}M}\not\cong\mu sh_{S^{*}M,m_{can}}$.
-Indeed, in this case the conormal Legendrian to $N\times 0$ can be seen to support a microsheaf in $\mu sh_{S^{*}M}$ but not in $\mu sh_{S^{*}M,m_{can}}$.
-This corresponds to the observation of Kragh that $Fuk(T^{*}M)\neq Loc(M)$ when $M$ is not spin [17, Remark 1.3].
-Remark 4.20.
-We return to the case of general $C$.
-Consider first the universal case $C=\mathbb{S}-mod$ of spectra.
-The map $\mu_{\mathbb{S}-mod}$ was calculated in [9]; the result is that it is a delooping of the $J$-homomorphism [9].
-The map $\mathbb{S}\to\pi_{0}(\mathbb{S})=\mathbb{Z}$ induces $Pic(\mathbb{S}-mod)\to\pi_{\leq 1}Pic(\mathbb{S}-mod)=Pic(\mathbb{Z}-mod)$.
-In particular, the inclusion of the shift $\mathbb{Z}\to Pic(\mathbb{S}-mod)$ splits and we can write $Pic(\mathbb{S}-mod)=\mathbb{Z}\oplus Pic(\mathbb{S})_{0}$, where $Pic(\mathbb{S})_{0}$ is the identity connected component.
-(It is known that $Pic(\mathbb{S})_{0}=B\mathbb{S}^{\times}$ for some $\mathbb{S}^{\times}$ with a natural interpretation as “the units of $\mathbb{S}$”.) We write $w_{\mathbb{S}-mod}$ for the composition $LGr\to BPic(\mathbb{S}-mod)\to B^{2}\mathbb{S}^{\times}$, and define $\mathbb{S}-mod$ orientation data to be a null-homotopy of the appropriate composition with $Bw_{\mathbb{S}-mod}$.
-Tautologically, then, $\mathbb{S}$-mod Maslov data is the same as a grading and $\mathbb{S}-mod$ orientation data.
-For more general $C$, let us write $Pic(C)_{0}$ for the identity connected component.
-Then the canonical map $\mathbb{S}-mod\to C$ induces a map $Pic(\mathbb{S})_{0}\to Pic(C)_{0}$, and we write $w_{C}$ for the composition of $w_{\mathbb{S}-mod}$ with the delooping of this map.
-We define $C$-orientation data to be a null-homotopy of the appropriate composition with $Bw_{C}$.
-Evidently a grading and $C$-orientation datum yield a $C$-Maslov datum, though not all $C$-Maslov data need arise in this way.
-Still more generally, suppose given any partition $\Pi$ of the set of isomorphism classes of objects of $C$.
-We write $Pic(C)_{\Pi}\subset Pic(C)$ for the full subcategory on invertible elements which preserve the partition $\Pi$.
-By $(C,\Pi)$-orientation data, we mean a null-homotopy of the composite map to $BPic(C)_{\Pi}$.
-($C$-orientation data being the case of the total partition).
-In particular, for $D\subset C$ a subcategory, we consider the partition into objects in versus not in $D$, and term the corresponding orientation data as $(C,D)$-orientation data.
-5. Microsheaves on complex contact manifolds
+Integrability plays no role in the above arguments.
+That is, for the purpose of defining the canonical microstalk functor $\omega_{p}^{-1}$, it would be enough to assume that $W$ is an exact symplectic manifold endowed with a map $X\to BSp$ lifting the classifying map, and that $L$ is a (conical, smooth) Lagrangian endowed with a complex polarization $L\to LGr_{\mathbb{C}}$.
+Our next task is to compute $\omega_{p}^{-1}$ on cotangent bundles of complex manifolds, endowed with the Maslov datum induced by the complex polarization.
+Concretely, we are interested in the case when $L\subset S^{*}M$ is the conormal to a complex submanifold $N\subset M$ of complex codimension $n$-$\ell$, for $1\leq\ell\leq n$.
+By Darboux, we may assume without loss of generality that $U=J^{1}\mathbb{R}^{2n}$ and $L$ is the conormal to $\mathbb{R}^{2\ell}\times\{0\}^{2n-2\ell}$.
+Let $K\subset J^{1}(\mathbb{R}^{2n})$ be the zero section, so that $\psi_{\pi/2}(K)=L$, where $\psi_{\pi/2}(-)$ is defined as in Section 4.6.
+Henceforth we write $\psi=\psi_{\pi/2}$.
+Assume, as above, that $T^{*}M$ is endowed with the canonical $(\mathcal{C},\{\mathcal{D}_{i}\})$ grading and a fixed $(\mathcal{C},\{\mathcal{D}_{i}\})$ orientation $o$.
+We denote by $\eta$ the corresponding Maslov datum.
+We let $\phi_{K},\phi_{L}$ denote the canonical polarizations transverse to $K,L$.
+We denote by $\psi_{*}\eta:=\eta\circ\psi^{-1}$ and $\psi_{*}\phi_{K}:=d\psi(\phi_{K})$ the pushforwards of $\eta$ and $\phi_{K}$ under $\psi$.
+Observe that $\psi_{*}\phi_{K}:=d\psi(\phi_{K})=\phi_{L}$, by definition of $\psi$.
+(18) $\mathcal{C}$$(\mu sh_{K})_{p}$$(\mu sh_{K,\phi_{K}})_{p}$$(\mu sh_{K,\phi_{K}})_{p}=\mathcal{C}$$(\mu sh_{L,\phi_{L}})_{p}$$(\mu sh_{L,\phi_{L}})_{p}=\mathcal{C}$$\mathcal{C}$$(\mu sh_{L})_{p}$$(\mu sh_{L,\phi_{K}})_{p}$$(\mu sh_{L,\phi_{L}})_{p}=\mathcal{C}$$\scriptstyle{\simeq}$$\scriptstyle{[-2\ell]}\scriptstyle{\simeq}$$\scriptstyle{(\psi)_{*}}\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$$\scriptstyle{\omega_{p}^{-1}}$$\scriptstyle{=}\scriptstyle{\simeq}$$\scriptstyle{=}$$\scriptstyle{[-\ell]}\scriptstyle{\simeq}\scriptstyle$$\simeq$$\scriptstyle{\omega_{p}^{-1}}\scriptstyle{\simeq}$
+The dotted arrow means that the rightmost square commutes up to a noncanonical natural transformation (so its effect on objects is unambiguous).
+We first explain the meaning of the arrows in (18).
+The leftmost square is defined as in Lemma 4.13.
+The map $\mu sh_{K,\phi_{K}}\to\mu sh_{L,\phi_{K}}$ is induced from the contact transformation $\psi_{*}:\mu sh_{T^{*}K,\phi_{K}}\to\mu sh_{T^{*}K,\phi_{K}}$, after stabilizing with $(T^{*}\mathbb{R}^{N},0_{\mathbb{R}^{N}}),N\gg 1$.
+The identification $\mu sh_{K,\phi_{K}}=\mu sh_{\psi(K),\psi_{*}\phi_{K}}$ and the corresponding top-right commutative square are induced by the contactomorphism $\psi$.
+To explain the remaining arrows, recall that $\phi\mapsto(\mu sh_{L,\phi})_{p}$ forms a local system of categories over $LGr=U/O$, which is classified by $U/O\xrightarrow{BJ}BPic(\mathcal{C})$.
+The postcomposition $U/O\to BPic(\mathcal{C})\to B\mathbb{Z}=U(1)$ is precisely $det^{2}(-)$; hence up to the action of $Pic_{0}(\mathcal{C})$, the monodromy automorphismm of any loop is precisely shifting by the degree of the loop under $det^{2}(-)$.
+The canonical microstalk functor $\omega_{p}^{-1}:(\mu sh_{L,\phi_{K}})_{p}\to(\mu sh_{L,\phi_{L}})_{p}=\mathcal{C}$ is realized by choosing a homotopy of polarizations $\phi_{K}\rightsquigarrow\phi_{L}$ through complex Lagrangians, and parallel transporting.
+Similarly, the arrow $(\mu sh_{L,\phi_{L}})_{p}\to(\mu sh_{L,\phi_{K}})_{p}$ in induced by the homotopy of polarizations $\psi_{t}^{-1}:\psi_{*}\phi_{K}=\phi_{L}\rightsquigarrow\phi_{K}$.
+Hence the dotted arrow is induced by the loop of polarizations $\phi_{L}\rightsquigarrow\phi_{K}\rightsquigarrow\phi_{L}$, which defines an automorphism $\mathcal{C}=(\mu sh_{L,\phi_{L}})_{p}$.
+.
+The first homotopy lies in the kernel of $det^{2}(-)$, while the image of the path $\psi_{t}^{-1}:\psi_{*}\phi_{K}\rightsquigarrow\phi_{K}$ under $det^{2}(-)$ is a loop of degree -$\ell$.
+Corollary 5.11.
+Let $M,\eta$ be as above.
+Suppose that $L$ is a smooth Lagrangian disk contained in the the conormal of a complex submanifold $N\subset M$ of complex codimension $\ell$, for $1\leq\ell\leq n$.
+Let $p\in L$ be a smooth point.
+Given $A\in\mathcal{C}$, $\omega_{p}^{-1}(A_{N})=A[\ell]$.
+Proof.
+Without loss of generality, we can assume that $M=\mathbb{R}^{2n}$ and $N=\mathbb{R}^{2\ell}\times\{0\}^{2n-2\ell}$.
+Let $K=0_{\mathbb{R}^{2n}}$.
+Then $\omega_{p}^{-1}(A_{K})=A$; by (18), $\omega_{p}^{-1}(A_{N}[-2\ell])=\omega_{p}^{-1}(A_{K})[-\ell]$, which proves the claim.
+∎
+5.3. Microsheaves on complex contact manifolds and symplectic manifolds
 Let $V$ be a complex contact manifold. Recall from Section 2 that we have maps
 […]
-where $\tilde{V}$ is an exact complex symplectic manifold with $1$-form $\lambda_{\tilde{V}}$, and $\tilde{V}/\mathbb{R}_{+}$ carries a $\mathbb{C}^{*}$ of real contact forms $\xi_{\hbar}:=\operatorname{re}(\hbar\lambda_{\tilde{V}})$.
-Evidently $(\tilde{V},\operatorname{re}(\hbar\lambda_{\tilde{V}}))$ is the real symplectization of $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$.
-For specificity we fix $\hbar=0$ (but see Remark 5.7).
-For a complex Legendrian $L\subset V$, we similarly write $\tilde{L}:=\pi^{-1}(L)$ and $\tilde{L}/\mathbb{R}_{+}:=p^{-1}(L)$.
+where $\widetilde{V}$ is an exact complex symplectic manifold with holomorphic $1$-form $\lambda_{\widetilde{V}}$.
+For any subset (typically complex Legendrian) $L\subset V$, we similarly write $\widetilde{L}:=\pi^{-1}(L)$ and $\widetilde{L}/\mathbb{R}_{+}:=p^{-1}(L)$.
+Letting $\hbar\in\mathbb{C}^{*}$ act by the $\mathbb{C}^{*}$ principal bundle structure of $\widetilde{V}$ over $V$, we have $\hbar^{*}\lambda_{\widetilde{V}}=\hbar\lambda_{\widetilde{V}}$ for $\hbar\in\mathbb{C}^{*}$.
+In particular, multiplication by $\hbar$ descends to a real contactomorphism $(\widetilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\lambda)\cong(\widetilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\hbar\lambda)$.
+Per Definition 3.8, $(\widetilde{V},\operatorname{re}\lambda)$ carries canonical Maslov data $\eta_{can}$. We write
+$\mu sh_{\widetilde{V}}\coloneqq\mu sh_{(\widetilde{V},\operatorname{re}\lambda),\eta_{can}}.$
+Here we have, and will henceforth, set $\hbar=1$. This is no loss of generality:
+Lemma 5.12.
+There is a canonical isomorphism $\hbar^{*}\mu sh_{(\widetilde{V},\operatorname{re}\hbar\lambda)}\cong\mu sh_{(\widetilde{V},\operatorname{re}\lambda)}$.
+Proof.
+The canonical Maslov datum is pulled back from $V$, so the action of the contactomorphism given by multiplication by $\hbar$ is also canonically trivial on it. ∎
+More generally, we can consider any Maslov data $\eta$ on $\widetilde{V}$ and the associated sheaf $\mu sh_{\widetilde{V},\eta}$.
+Remark 5.13.
+The sheaf of categories $\mu sh_{\widetilde{V}}$ is the pullback of the corresponding object on the real contact manifold $\widetilde{V}/\mathbb{R}_{>0}$; we adopt our present formulation solely to avoid the minor cognitive dissonance of passing constantly to a non-complex manifold.
+A closed and complex analytic subset $\widetilde{\Lambda}\subset\widetilde{V}$ is $\mathbb{R}_{>0}$-invariant iff it is $\mathbb{C}^{*}$-invariant, and hence the preimage of some $\Lambda\subset V$.
+We will however also write $\widetilde{\Lambda}$ for complex and $\mathbb{R}_{>0}$-invariant (but not necessarily closed or $\mathbb{C}^{*}$-invariant) subsets of $\widetilde{V}$.
+We say such a subset is Lagrangian if it is contained in the closure of its smooth and Lagrangian locus.
 Definition 5.14.
-Let $V$ be a complex contact manifold and $\tilde{V}$ its symplectization.
-Per Lemma 3.2 and Remark 3.4, there is a canonical grading for $\tilde{V}$.
-Choose $C$-orientation data $o$ on $V$ and pull it back to $\tilde{V}$; combining this with with the canonical grading, we obtain Maslov data.
-We also write $o$ for this Maslov data, and correspondingly $\mu sh_{\tilde{V},o}$ for the corresponding sheaf of categories on $\tilde{V}$.
-In case $C=R-mod$ for a commutative ring $R$, then we have available the canonical orientation data; when we make this choice, we omit $o$ from the notation entirely, writing simply $\mu sh_{\tilde{V}}$.
-For $\tilde{\Lambda}\subset\tilde{V}$ a $\mathbb{R}_{>0}$-invariant complex-analytic subset which is the closure of its smooth Lagrangian locus, we write $\mu sh_{\tilde{\Lambda}}\subset\mu sh_{\tilde{V}}$ for the subsheaf of full subcategories on objects supported in $\tilde{\Lambda}$.
-We write $\mu sh_{\tilde{V},\mathbb{C}-c,o}$ for the union of all such $\mu sh_{\tilde{\Lambda},o}$.
+We write $\mu sh_{\widetilde{V},\mathbb{C}-c,\eta}\subset\mu sh_{\widetilde{V},\eta}$ for the sheaf of full subcategories on objects with complex Lagrangian microsupport.
+By definition, for $\mathbb{R}_{>0}$-conic subsets $\Omega\subset\widetilde{V}$, the space of sections $\mu sh_{\widetilde{V},\mathbb{C}-c,o}$ is the union of the categories $\mu sh_{\widetilde{\Lambda},o}(\Omega)$, where $\widetilde{\Lambda}$ varies over $\mathbb{R}_{>0}$-conic complex Lagrangian subsets $\widetilde{\Lambda}$ which are closed in $\Omega$.
 Remark 5.15.
@@
 The situation for the microsheaf categories is entirely analogous.
-The distinction is rarely relevant: for example, in the present article, while we state theorems for $\mu sh_{\tilde{V},\mathbb{C}-c,o}$, their proofs quickly reduce to statements about $\mu sh_{\tilde{\Lambda},o}$.
-Remark 5.3.
-A closed and complex analytic subset $\tilde{\Lambda}\subset\tilde{V}$ is $\mathbb{R}_{>0}$-invariant iff it is $\mathbb{C}^{*}$-invariant.
-We use the formulation above because we may want to consider open sets which are $\mathbb{R}_{>0}$-invariant but not $\mathbb{C}^{*}$-invariant, and so need a corresponding local notion.
-Remark 5.4.
-Per Remark 4.17, we could equivalently have defined a category of microsheaves on the real contact manifold $\tilde{V}/\mathbb{R}_{>0}$ and pulled it back along $q:\tilde{V}\to\tilde{V}/\mathbb{R}_{+}$; our choice of formulation above is solely to avoid the minor cognitive dissonance of passing constantly to a non-complex manifold.
-Remark 5.5.
-One could also consider the subsheaf of full subcategories of $\mu sh_{\tilde{V}}$ on objects whose microsupport is complex-analytic, but this would e.g. not be closed under taking summands.
-Lemma 5.6.
-Let $f:U\to V$ be a complex contact embedding, and $\tilde{f}:\tilde{U}\to\tilde{V}$ its lift to the complex symplectization.
-Then there is a canonical isomorphism $\mu sh_{U,f^{*}o}=f^{*}\mu sh_{V,o}$.
-In particular, when $o$ is the canonical orientation datum, there is a canonical isomorphism $\mu sh_{U}=f^{*}\mu sh_{V}$.
-Proof.
-It suffices to argue that the Maslov data of $V$ pulls back to the Maslov data of $U$, up to canonical homotopy.
-The orientation data is respected by definition, while the grading is canonically induced by the complex structure, hence behaves naturally under the above pullback.
-∎
-Remark 5.7.
-$\tilde{V}$ is a $\mathbb{C}^{*}$-principal bundle over $V$, and the natural $\mathbb{C}^{*}$ action evidently satisfies $\hbar^{*}\lambda_{\tilde{V}}=\hbar^{-1}\lambda_{\tilde{V}}$ for $\hbar\in\mathbb{C}^{*}$.
-In particular, multiplication by $\hbar$ descends to a real contactomorphism $(\tilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\lambda)\cong(\tilde{V}/\mathbb{R}_{>0},\ker\operatorname{re}\lambda)$.
-The action of $\hbar$ is also canonically trivial on Maslov data, as this is pulled back from $V$.
-Thus by similar reasoning as in Lemma 5.6, there is a canonical isomorphism $\hbar^{*}\mu sh_{(\tilde{V},\hbar\lambda)}\cong\mu sh_{(\tilde{V},\lambda)}$.
-Lemma 5.8.
-Let $M$ be a complex manifold.
-Consider the complex contact manifold $V=\mathbb{P}T^{*}M$, and the corresponding $\tilde{V}=T^{\circ}M$, with canonical orientation data.
-A spin structure on $M$ determines an equivalence of sheaves of categories on $T^{\circ}M$: $\mu sh\cong\mu sh_{T^{\circ}M}$ where the left hand side is as in (9) and the right hand side is defined as in Definition 5.1.
-Proof.
-Follows from Corollary 4.13 by Lemma 3.11 and Lemma 3.12. ∎
-Even if we had used some arbitrary choice of Maslov data $\mu$ on $V$, not necessarily coming from orientation data and the canonical complex grading, the (complex) Darboux theorem and general functoriality of microsheaves would nevertheless imply that $\mathbb{P}\mu sh_{V,\mu}$ is locally isomorphic to any $\mathbb{P}\mu sh_{\mathbb{P}T^{*}M}$ of the correct dimension, since the space of Maslov data, being a torsor for $Maps(V,BPic(C))$, is certainly connected when $V$ is contractible.
-However, the resulting isomorphism is not canonical, even locally.
-What we gain from using the canonical grading is:
-Lemma 5.9.
-Let $U$ be a contractible complex contact manifold, and $u:U\to V$ be any complex contact embedding.
-Then for any choices of $C$-orientation data $o,o^{\prime}$, there is an isomorphism $\mu sh_{U,o}\cong u^{*}\mu sh_{V,o^{\prime}}$ which is canonical up to a non-canonical natural transformation.
-Proof.
-Using Lemma 5.6, a homotopy of orientation data $u^{*}o^{\prime}\sim o$ will determine such an isomorphism.
-This is defined up to the action of loops in the space of orientation data; as $U$ is contractible, this is just (by definition of $C$-orientation data, see Remark 4.20) the connected component of the identity $Pic(C)_{0}\subset Pic(C)$, which acts by natural transformations.
-∎
-If $\Lambda\subset V$ is a Legendrian in a real contact manifold equipped with appropriate orientation and grading data, it follows, ultimately from Proposition 4.6, that there is a non-canonical isomorphism $(\mu sh_{\Lambda})_{p}\xrightarrow{\sim}C$ at any smooth point $p\in\Lambda$.
-Any choice of such an isomorphism is termed a microstalk functor; there is a $Pic(C)$ worth of choices.
-In (10), we saw how to define a distinguished microstalk functor for Legendrians which occur as conormals to embedded submanifolds, where the orientation data is induced by the fiber polarization.
-However, leaving this class of submanifolds, the torsor of microstalk functors has some monodromy (classified by the map $\mu_{C}$ from Theorem 4.14), typically involving nontrivial shifts.
-A central observation of this paper is that the situation is better in the complex world:
-Definition-Proposition 5.10.
-Let $V$ be a complex contact manifold equipped with some orientation data $o$. Let $L\subset V$ be a complex Legendrian, and $p\in L$ a smooth point with lift $\tilde{p}\in\tilde{L}$.
-Choose a codimension zero embedding $\iota:Nbd(p)\hookrightarrow\mathbb{P}T^{*}M$ for a complex manifold $M$, such that the projection $L\cap Nbd(p)\to M$ is an embedding onto its image.
-Consider the composition of isomorphisms:
-$C\xrightarrow{\omega_{\iota(\tilde{p})}}(\mu sh_{\iota(\tilde{L})})_{\iota(\tilde{p})}=(\mu sh_{\iota(\tilde{L}),\phi})_{\tilde{p}}\cong(\mu sh_{\tilde{L},o})_{\tilde{p}}$
-here, the first arrow is (10), the second is (the $T^{\circ}M$ version of) Corollary 4.13, and the last is Lemma 5.9.
-We have seen the first two maps are independent of choices, and the final map depends on choices only through (noncanonical) natural transformation.
-We denote the composition by $\omega_{\tilde{p}}$, and, henceforth, when we refer in this context to microstalk functors, we mean only those of the form $\omega_{\tilde{p}}^{-1}$.
-In particular, given $F\in(\mu sh_{\tilde{L},o})_{\tilde{p}}$, the object $\omega_{\tilde{p}}^{-1}(F)\in C$ is well defined (i.e.
-independent of the embedding $\iota$).
-We term this object the microstalk at $\tilde{p}$.
-Definition 5.11.
-Fix a subcategory $D$ of our coefficient category $C$, and fix $C$-orientation data $o$ (or more generally, $(C,D)$-orientation data).
-Let $V$ be a complex contact manifold.
-We write $(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{D}\subset\mu sh_{\tilde{V},\mathbb{C}-c,o}$ for the full subcategory on objects $F$ such that at every smooth point $p\in supp(F)$, one has $\omega_{\mathbb{C}}^{-1}(F_{\tilde{p}})\in D$.
-Remark 5.12.
-By the same argument as in Lemma 5.9, the isomorphism class of $\omega_{\mathbb{C}}^{-1}(F_{\tilde{p}})$ is constant in the smooth locus of the microsupport.
-So we need only check $\omega_{\mathbb{C}}^{-1}(F_{\tilde{p}})\in D$ at one point in each connected component of the smooth locus.
-Legendrians give objects:
-Proposition 5.13.
-Let $V$ be a complex contact manifold and $L\subset V$ a complex Legendrian.
-Then relative orientation data on $L$ (e.g.
-a spin structure if we use the canonical orientation datum on $V$) determines an equivalence of categories $\mu sh_{\tilde{V},\mathbb{C}}|_{\tilde{L}}\cong Loc_{\tilde{L}}$ where $Loc_{\tilde{L}}$ is the sheaf of categories of local systems.
-Proof.
-As described in Remark 4.16, it suffices to provide secondary Maslov data.
-The grading part comes from Lemma 3.11.
-Lemma 3.12 explains why, for the canonical orientation datum, the corresponding notion of relative orientation data reduces is identified with spin structure.
-∎
-We now introduce sheaves of categories which will carry the microlocal perverse $t$-structure.
-Definition 5.14.
-Let $V$ be a complex-contact manifold.
-Given an object $K\in\pi_{*}\mu sh_{\tilde{V}}(U),U\subset V,$ and a point $p\in ss(K)\cap U$, we say that $K$ is representable at $p$ by a sheaf if there exists a contact embedding $\iota:Nbd(p)\hookrightarrow\mathbb{P}^{*}M$ such that (the image of) $K$ in $(\pi_{*}\mu sh_{T^{*}M-0_{M}})_{p}$ is represented by a sheaf $\mathcal{F}_{K}\in sh(M)$.222i.e.
-the image of $K$ in $(\pi_{*}\mu sh_{T^{*}M-0_{M}})_{p}$ coincides with the image of $\mathcal{F}_{K}$ under the quotient $sh(M)\to(\pi_{*}\mu sh_{T^{*}M-0_{M}})_{p}$.
-Here $M$ is any complex manifold and we recall that $\pi$ is the projection $T^{*}M-0_{M}=\widetilde{\mathbb{P}^{*}M}\to\mathbb{P}^{*}M$ – not projection to the base.
-We say that $K$ is pointwise representable by a sheaf if it is so at all points in its microsupport.
-Remark 5.15.
-By the discussion around [15, (11.4.8)] this definition is independent of the choice of the Darboux chart.
-Definition 5.16.
-Let $V$ be a contact manifold, $o$ orientation data, and $\Lambda$ a complex (singular) Legendrian. We denote by:
-$\mathbb{P}\mu sh_{V,o}\subset\pi_{*}\mu sh_{\tilde{V},o}\qquad\qquad\mathbb{P}\mu sh_{\Lambda,o}\subset\pi_{*}\mu sh_{\pi^{-1}(\Lambda),o}\qquad\qquad\mathbb{P}\mu sh_{V,\mathbb{C}-c,o}\subset\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c,o}$
-the sheaves of full subcategories of $\pi_{*}\mu sh_{\tilde{V},o}$ (resp.
-$\pi_{*}\mu sh_{\pi^{-1}(\Lambda),o}$, $\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c,o}$) on objects which are pointwise representable by sheaves.
-Remark 5.17.
-It is a tautology that $\mathbb{P}\mu sh_{V,\mathbb{C}-c}$ is a union over all complex (singular) legendrians $\Lambda$ of $\mathbb{P}\mu sh_{\Lambda}$.
-Remark 5.18.
-We could equivalently have pushed forward the corresponding sheaves of categories defined on the circle bundle $\tilde{V}/\mathbb{R}_{>0}$.
-6. Microsheaves on complex exact symplectic manifolds
-Let $X=(X,\lambda)$ be an exact complex symplectic manifold. Consider the contact thickening $(X\times\mathbb{C},\lambda+dz)$ and the associated symplectization
-(13) $(X\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz)).$
-Let $V_{\mathbb{R}}$ denote the Liouville vector field for $(X,\lambda)$ and let $I$ denote the almost-complex structure induced by complex multiplication.
-Integrating the flow of $Z$ in the first component and the flow of $V_{S^{1}}:=IV_{\mathbb{R}}$ in second, we obtain an action $A:\mathbb{R}\times\mathbb{R}\times X\to X$ which satisfies $A_{(t,\theta)}^{*}\lambda=e^{t+2\pi i\theta}\lambda$.
-If $U\subset\mathbb{C}^{*}$ is contractible, then we can define a action $U\times X\to X$ by lifting $U$ to $\mathbb{R}\times\mathbb{R}$ by the covering map $(t,\theta)\mapsto e^{t+2\pi i\theta}$.
+The distinction is rarely relevant: for example, in the present article, while we state theorems for $\mu sh_{\widetilde{V},\mathbb{C}-c,o}$, their proofs quickly reduce to statements about $\mu sh_{\widetilde{\Lambda},o}$.
+We now discuss complex symplectic manifolds.
+Let $W=(W,\lambda)$ be an exact complex symplectic manifold, consider the contact thickening $(W\times\mathbb{C},\lambda+dz)$ and the associated symplectization
+(20) $(W\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz)).$
+Let $V_{\mathbb{R}}$ denote the Liouville vector field for $(W,\lambda)$ and let $I$ denote the almost-complex structure induced by complex multiplication.
+Integrating the flow of $V_{\mathbb{R}}$ in the first component and the flow of $V_{S^{1}}:=IV_{\mathbb{R}}$ in second, we obtain an action $A:\mathbb{R}\times\mathbb{R}\times W\to W$ which satisfies $A_{(t,\theta)}^{*}\lambda=e^{t+2\pi i\theta}\lambda$.
+If $U\subset\mathbb{C}^{*}$ is contractible, then we can define an action $U\times W\to W$ by lifting $U$ to $\mathbb{R}\times\mathbb{R}$ by the covering map $(t,\theta)\mapsto e^{t+2\pi i\theta}$.
 In many situations, the flow in the $IZ$ direction factors through $k\mathbb{Z}\subset\mathbb{R},k\geq 1$; in this case, we can lift the action via the covering map $(t,\theta)\mapsto e^{k(t+2\pi i\theta)}$ to define a weight $k$ action of $\mathbb{C}^{*}$.
@@
 We record the following convenient change of variables:
-(14) $(X\times\mathbb{C}\times U,\lambda+wdz)$ $\xrightarrow{\simeq}(X\times\mathbb{C}\times U,w(\lambda+dz))$
+(21) $(W\times\mathbb{C}\times U,\lambda+wdz)$ $\xrightarrow{\simeq}(W\times\mathbb{C}\times U,w(\lambda+dz))$
 $(x,z,w)$ $\mapsto(A_{w^{-1}}(x),z,w)$
-We let $\pi:X\times\mathbb{C}\times\mathbb{C}^{*}\to X\times\mathbb{C}$ be the projection which forgets the last component.
-Recall that in 5.10, we show that, for a complex contact manifold $V$ and a complex Legendrian $L\subseteq V$, we can fix a canonical identification $(\mu sh_{\tilde{L}})_{\tilde{p}}\simeq C$ where $\tilde{L}$ is its Lagrangian lift and $\tilde{p}$ a smooth point.
-The above change of coordinates implies that a similar choice can be made for the symplectic case.
-Lemma 6.1.
-Suppose now that $L\subset X$ is a (possibly singular) conic complex-Lagrangian. Given $p\in L$, we have
-$(\mu sh_{L})_{p}=\left(\mu sh_{\pi^{-1}(L\times\{0\})}\right)_{(p,0,1)}$
-where the left hand side is defined as in [22, Section 9.1], for any exact symplectic manifolds, and the right hand side is defined in Definition 5.1, only for symplectizations.
+Given an exact complex-symplectic manifold $(W,\lambda)$, consider the contactization $(W\times\mathbb{C},\lambda+dz)$.
+This is a complex-contact manifold, and so we can consider $\mu sh_{W\times\{0\}}\subset\mu sh_{W\times\mathbb{C}}$, which is the full subcategory on objects microsupported on $W\times\{0\}$.
+We also write $\mu sh_{W\times\{0\},\mathbb{C}-c}$ for the full subcategory of $\mu sh_{W\times\{0\}}$ on objects with complex constructible support.
+We now have two sheaves of categories on $W$, namely $\mu sh_{W}\coloneqq\mu sh_{(W,\operatorname{re}\lambda)}$ from Definition 4.25 and $\mu sh_{W\times\{0\}}$.
+They are not the same: for example, if $W$ is a point, then $\mu sh_{W}=C$ while $\mu sh_{W\times\{0\}}$ is the category of local systems on $\mathbb{C}^{*}$.
+Theorem 5.16.
+Suppose that the Liouville vector field of $(W,\lambda)$ integrates to a weight-1 $\mathbb{C}^{*}$-action.
+Let $\gamma_{\mathbb{C}}:W\to W$ be the set-theoretic identity, where the source is endowed with the Euclidean topology and the target with the $\mathbb{C}^{*}$-invariant topology.
+Then there is a natural $\mathbb{Z}=\Omega S^{1}$-linear structure on $(\gamma_{\mathbb{C}})_{*}\mu sh_{W\times\{0\}}$ and an equivalence
+(22) $(\gamma_{\mathbb{C}})_{*}\left(\mu sh_{W\times\{0\}}\otimes_{\Omega S^{1}}\bullet\right)\simeq(\gamma_{\mathbb{C}})_{*}\mu sh_{W}.$
+Furthermore, this equivalence respects complex constructibility.
 Proof.
-We have $\pi^{-1}(L\times\{0\})=L\times T^{\circ}_{0}\mathbb{C}=L\times\{0\}\times\mathbb{C}^{*}$. As a consequence of the local splitting (14), we have
-(15) $\left(\mu sh_{\pi^{-1}(L\times\{0\})}\right)_{(p,0,1)}\simeq(\mu sh_{L})_{p}\otimes(\mu sh_{T^{\circ}_{0}\mathbb{C}})_{(0,1)}\simeq(\mu sh_{L})_{p}$
-We now define the microstalk functor for exact complex-symplectic manifolds.
-Definition 6.2.
-Let $(X,\lambda)$ be an exact complex-symplectic manifold equipped with some orientation data $o$.
-Let $L\subset X$ a (germ of a) conical complex Lagrangian submanifold.
-For $p\in L$ a smooth point, we have a microstalk funtor defined as the composition
-$C\to\left(\mu sh_{\pi^{-1}(L\times\{0\}),o}\right)_{(p,0,1)}=(\mu sh_{L,o})_{p}$
-where the first arrow is the microstalk functor introduced in 5.10 and the equality is given by (15) above.
-Proof of Theorem 1.2.
-Per Lemma 3.2 and Remark 3.4, there is a canonical grading for $\tilde{V}$.
-Choose the canonical orientation data.
-Combining these, we obtain Maslov canonical data and hence a canonical sheaf of categories $\mu sh_{X}$.
-Note that here we are viewing $X=(X,\operatorname{re}\lambda)$ as a real exact symplectic manifold, so that one can define $\mu sh_{X}$ as in [22, Section 9.1].
-More precisely, we identify $X$ with the subset $X\times\{0\}$ of its contactization $X\times\mathbb{R}$, and we then define $\mu sh_{X}$ as the restriction $\mu sh_{X\times\{0\}}|_{X\times\{0\}}$.
-Finally, if $L$ is a smooth smooth conic complex spin Lagrangian, the canonical equivalence $\mu sh_{W,L}(-)\simeq loc(L)$ is established exactly like Proposition 5.13. ∎
-Given an exact complex-symplectic manifold $(X,\lambda)$, consider the contactization $(X\times\mathbb{C},\lambda+dz)$.
-This is a complex-contact manifold, and so we can consider $\mathbb{P}\mu sh_{X\times\{0\}}\subset\mathbb{P}\mu sh_{X\times\mathbb{C}}$, which is the full subcategory on objects microsupported on $X\times\{0\}$.
-We now have two sheaves of categories on $X$, namely $\mu sh_{X}$ and $\mathbb{P}\mu sh_{X\times\{0\}}$.
-They are not the same: for example, if $X$ is a point, then $\mu sh_{X}=C$ while $\mathbb{P}\mu sh_{X\times\{0\}}$ is the category of local systems on $\mathbb{C}^{*}$.
-In general, we do not know how to relate $\mu sh_{X}$ and $\mathbb{P}\mu sh_{X\times\{0\}}$.
-However:
-Theorem 6.3.
-When the Liouville vector field of $(X,\lambda)$ integrates to a weight-1 $\mathbb{C}^{*}$-action, there is a natural $\mathbb{Z}=\Omega S^{1}$-linear structure on $\mathbb{P}\mu sh_{X\times\{0\}}$, and there are fully faithful embeddings
-(16) $\mathbb{P}\mu sh_{X\times\{0\}}\hookrightarrow\mu sh_{X}\otimes\Omega S^{1}$
-(17) $\mu_{\mathbb{C}}sh_{X}:=\mathbb{P}\mu sh_{X\times\{0\}}\otimes_{\Omega S^{1}}\bullet\hookrightarrow\mu sh_{X}$
-Of course, when working over a discrete ring, we may replace $\Omega S^{1}$ and $\bullet$ by their linearizations $\mathbb{Z}[\mathbb{Z}]=\mathbb{Z}[t,t^{-1}]$ and $\mathbb{Z}$.
-Proof.
-The change of variables (14) is global:
-(18) $(\widetilde{X\times\mathbb{C}}=X\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz))$ $\xrightarrow{\simeq}(X\times\mathbb{C}\times\mathbb{C}^{*},\lambda+wdz)$
+Up to replacing $W$ by a $\mathbb{C}^{*}$-invariant open, it is enough to exhibit the $\Omega S^{1}$-linear structure on $\mu sh_{W\times\{0\}}(W\times\{0\})$, and to prove the equivalence (22) on global sections.
+By our assumption on Liouville vector field, the change of variables (21) is global:
+(23) $(\widetilde{W\times\mathbb{C}}=W\times\mathbb{C}\times\mathbb{C}^{*},w(\lambda+dz))$ $\xrightarrow{\simeq}(W\times\mathbb{C}\times\mathbb{C}^{*},\lambda+wdz).$
 The Liouville structure on the right hand side is a product; we have the Künneth isomorphism:
-$\mu sh_{X\times 0\times\mathbb{C}^{*}}=\mu sh_{X}\boxtimes\mu sh_{\mathbb{C}^{*}\subset T^{*}\mathbb{C}^{*}}.$
-Pushing forward and composing with the embedding $\mathbb{P}\mu sh_{X\times 0}\hookrightarrow\pi_{*}\mu sh_{X\times 0\times\mathbb{C}^{*}}$ gives the first assertion, from which the second follows formally.
-∎
+(24) $\mu sh_{W\times 0\times\mathbb{C}^{*}}=\mu sh_{W}\boxtimes\mu sh_{\mathbb{C}^{*}\subset T^{*}\mathbb{C}^{*}}.$
+By definition
+[…]
+where $\mu sh_{\pi^{-1}(W\times\{0\})}$ denotes the sheaf of full subcategories of $\mu sh_{\widetilde{W\times\mathbb{C}}}$ on objects whose support is contained in $\pi^{-1}(W\times\{0\})$; see (19).
+But by (23) and (24), we have
+$\left(\pi_{*}\mu sh_{\pi^{-1}(W\times\{0\})}\right)(\widetilde{W\times\mathbb{C}})=\mu sh_{W}(W)\otimes loc(\mathbb{C}^{*}).$
 Remark 5.17.
-(17) amounts to taking invariants of a topological $\mathbb{C}^{*}$-action; see [20, Sec.
+Let $\mathbb{C}^{*}\times W\to W,(\theta,z)\mapsto\theta\cdot z$ be the weight-11 $\mathbb{C}^{*}$ action on $W$.
+Then $(\theta;x,z)\mapsto(\theta\cdot x,\theta z)$ defines a $\mathbb{C}^{*}$ action on $W\times\mathbb{C}$ by contactomorphism, which fixes $W\times\{0\}$ set-wise.
+(22) amounts to taking invariants of this action; see [33, Sec.
 6].
-With respect to the split symplectic form (i.e.
-the right hand side of (18)), this action is $\theta\cdot(x,z,w)\mapsto(x,\theta z,\theta^{-1}w)$.
-With respect to the coupled symplectic form (i.e.
-the left hand side of (18)), this corresponds to the diagonal $\mathbb{C}^{*}$ action $\theta(x,z,w)=(\theta\cdot x,\theta z,\theta^{-1}w)$.
-[^3] Note the slight abuse of notation: $\theta\cdot x$ refers to the image of $x\in X$ under the time $\theta$ flow of the Liouville vector field, which integrates by assumption to a $\mathbb{C}^{*}$ action.
-Observe that this is just the lift to the symplectization of the $\mathbb{C}^{*}$-action $\theta\cdot(x,z)=(\theta\cdot x,\theta z)$ on $X\times\mathbb{C}$ by contactomorphism.
-Remark 6.5.
-If $(X,\omega)$ admits a weight $\mathbb{C}^{*}$-action of weight $k>1$, we expect a related statement, except that $\mathbb{P}\mu sh_{X}$ should be a full subcategory of $(\mu sh_{X})^{\mathbb{Z}/k}$.
 6. The perverse t-structure
@@
 1.3.6].
-The prototypical example is when $\mathcal{T}$ is a derived category of chain complexes, and $\mathcal{T}^{\leq 0}$ consists of the complexes whose cohomology is concentrated in degrees $\leq 0$, etc.
+The prototypical example is when $\mathcal{T}$ is a derived category of chain complexes, $\mathcal{T}^{\leq 0}$ (resp.
+$\mathcal{T}^{\geq 0}$) consists of the complexes whose cohomology is concentrated in degrees $\leq 0$ (resp.
+$\geq 0$).
 Let us recall a result about when $t$-structures pass to quotient categories.
@@
     if the equivalent assertions of (1) hold, $Q(\mathcal{T}^{\leq 0}),Q(\mathcal{T}^{\geq 0})$ determine a $t$-structure if and only if $\mathcal{I}\cap\mathcal{T}^{\heartsuit}\subset\mathcal{T}^{\heartsuit}$ is a “Serre subcategory” (meaning it is closed under extensions, quotients and sub-objects).
-The notion of $t$-structure is imported to the setting of stable categories in [19, Sec.
+The notion of $t$-structure is imported to the setting of stable categories in [30, Sec.
 1.2]: By definition, a $t$-structure on a stable category is a $t$-structure on its homotopy category, which canonically carries the structure of a triangulated category.
@@
 Remark 6.3.
-When $C$ is a presentable stable category, then if either $C^{\geq 0}$ or $C^{\leq 0}$ is presentable, then so is the other, and all truncation functors are colimit preserving [19, 1.4.4.13].
-In this case, the subcategory of compact objects $C^{c}$ is stable under the truncation functors and inherits a $t$-structure.
+When $\mathcal{C}$ is a presentable stable category, then if either $\mathcal{C}^{\geq 0}$ or $\mathcal{C}^{\leq 0}$ is presentable, then so is the other, and all truncation functors are colimit preserving [30, 1.4.4.13].
+In this case, the subcategory of compact objects $\mathcal{C}^{c}$ is stable under the truncation functors and inherits a $t$-structure.
 Indeed, $\tau^{\geq 0}$ is left adjoint to the corresponding inclusion, assumed colimit preserving, hence $\tau^{\geq 0}$ preserves compact objects.
@@
 We say a pair of sheaves of full subcategories $\mathcal{F}^{\leq 0}$ and $\mathcal{F}^{\geq 0}$ define a $t$-structure on $\mathcal{F}$ if $\mathcal{F}^{\leq 0}(U)$ and $\mathcal{F}^{\geq 0}(U)$ define a $t$-structure on $\mathcal{F}(U)$ for all $U$.
-Remark 7.5.
-It is easy to see that the property that $\mathcal{F}^{\leq 0}$ and $\mathcal{F}^{\geq 0}$ define a $t$-structure may be checked on sections on any base of open sets.
+Lemma 6.5.
+The property that $\mathcal{F}^{\leq 0}$ and $\mathcal{F}^{\geq 0}$ define a $t$-structure may be checked on sections on any base of open sets.
+Proof.
 Indeed, regarding condition (i) and (ii) of Definition 6.1, it is immediate from the sheaf condition that vanishing of Homs and containment of subcategories can be checked locally.
 Regarding (iii), the key point is that for any candidate $t$-structure satisfying (i) and (ii), the space of fiber sequences $K^{\prime}\to K\to K^{\prime\prime}$ as requested in (iii) is either empty or contractible.
-Indeed, first recall that using property (ii) to apply (i) to shifts, we find that the following strengthening of (i): for any $K^{\prime}\in T^{\leq 0}$ and $K^{\prime\prime}\in T^{\geq 0}$, the negative exts, aka positive homotopy groups of the hom space $\operatorname{Hom}(K^{\prime},K^{\prime\prime}[-1])$, must vanish.
+Indeed, first recall that using property (ii) to apply (i) to shifts, we find the following strengthening of (i): for any $K^{\prime}\in T^{\leq 0}$ and $K^{\prime\prime}\in T^{\geq 0}$, the negative exts, aka positive homotopy groups of the hom space $\operatorname{Hom}(K^{\prime},K^{\prime\prime}[-1])$, must vanish.
 Now given any $K^{\prime}\to K\to K^{\prime\prime}[-1]$ and $L^{\prime}\to K\to L^{\prime\prime}[-1]$ both satisfying (iii), we obtain a canonical null-homotopy of the composition $K^{\prime}\to K\to L^{\prime\prime}[-1]$ hence lift of $K^{\prime}\to K$ to $K^{\prime}\to L^{\prime}$, etc.
-Having learned this contractibility, if (iii) holds locally, then we can canonically glue the local exact triangles to obtain (iii) globally.
+Having learned this contractibility, if (iii) holds locally, then we can canonically glue the local exact triangles to obtain (iii) globally. ∎
 Since pullbacks commute with limits, $\mathcal{F}^{\heartsuit}:=\mathcal{F}^{\leq 0}\cap\mathcal{F}^{\geq 0}$ defines a sheaf of $(\infty,1)$-categories.
 As hearts of $t$-structures, these categories are abelian, in particular, 1-categories.
-We recall a comparison between various notions of sheaves of categories. For covers $U=\bigcup U_{i}$, an $\infty$-categorical sheaf of categories $\mathcal{C}$ satisfies the condition that
-$\mathcal{C}(U)=\lim\left(\prod_{i\in I}\mathcal{C}(U_{i})\rightrightarrows\prod_{i,j\in I}\mathcal{C}(U_{i}\cap U_{j})\mathrel{\begin{subarray}{c}\textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\end{subarray}}\prod_{i,j,k\in I}\mathcal{C}(U_{i}\cap U_{j}\cap U_{k})\cdots\right)$
-The corresponding definition of stacks in the classical literature instead involves a limit of 1-categories taken in the $(2,1)$-category of ordinary categories.
-(Note that for limits in $(2,1)$ categories, the compatibility condition on triple overlaps is already strict, hence the terms after triple overlaps in the above diagram are irrelevant to the limit.
-In the literature they are correspondingly not included in the descent condition.
-The situation is analogous to the fact that the descent condition for sheaves of sets is usually given truncated after the second term.)
-These notions agree: $1$-categories are $1$-truncated objects of $(\infty,1)$-categories, and, for any category $\mathcal{C}$, the inclusion of $k$-truncated objects $\tau_{\leq k}\mathcal{C}\hookrightarrow\mathcal{C}$ is limit-preserving [18, Proposition 5.5.6.5].
-We summarize:
+Remark 6.6.
+In the classical literature, the sheaf condition for sheaves of 1-categories (such sheaves are sometimes called stacks) is formulated as a limit of 1-categories taken in the $(2,1)$-category of ordinary categories.
+In this formulation, the compatibility condition on triple overlaps is strict.
+By contrast, the notion of $\infty$-categorical sheaf of $(\infty,1)$-categories requires that for such a sheaf $\mathcal{C}$ and covers $U=\bigcup U_{i}$, one has
+$\mathcal{C}(U)=\lim\left(\prod_{i\in I}\mathcal{C}(U_{i})\rightrightarrows\prod_{i,j\in I}\mathcal{C}(U_{i}\cap U_{j})\mathrel{\begin{subarray}{c}\textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\\[-2.58334pt] \textstyle\rightarrow\end{subarray}}\prod_{i,j,k\in I}\mathcal{C}(U_{i}\cap U_{j}\cap U_{k})\cdots\right).$
+These notions are equivalent for $1$-categories: one passes from the $\infty$-categorical notion to the $1$-categorical notion by truncation, and for the reverse direction one need only note that $1$-categories are $1$-truncated objects of $(\infty,1)$-categories, and the inclusion of $k$-truncated objects is limit-preserving [32, Proposition 5.5.6.5].
+6.2. The perverse t-structure on constructible sheaves
+We now review from [25, Sec. 10.3] the microlocal description of the perverse $t$-structure on constructible sheaves.
+Let $M$ be a complex manifold.
+For a Lagrangian subset $\Lambda\subset T^{*}M$, we write $\Lambda^{\circ}$ for the locus of smooth points of $\Lambda$ where the map $\Lambda\to M$ has locally constant rank.
+Fix a $t$-structure $\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}$ on our coefficient category $\mathcal{C}$, with corresponding truncation functors $\tau^{\leq 0},\tau^{\geq 0}$.
+It is proved in [25, Theorem 10.3.12] that the following prescription characterizes the perverse $t$-structure on $sh(M)_{\mathbb{C}-c}$.
+[^4] The results stated in [25] are for $\mathcal{C}$ the bounded derived category of modules over a ring and and $t$ the standard $t$-structure.
+However, the arguments given there (or in [3]) for the existence of the perverse $t$-structure depend only on the general properties of the six functor formalism, and the comparison between Definition 6.7 and the usual stalk/costalk wise definition of the perverse $t$-structure depends only on standard properties of microsupports.
+Definition 6.7 ([25, (10.3.7) and Definition 10.3.7]).
+Let $^{\mu}sh(M)_{\mathbb{C}-c}^{\leq 0}$ (resp.
+$^{\mu}sh(M)_{\mathbb{C}-c}^{\geq 0}$) be the full subcategory of $sh(M)_{\mathbb{C}-c}$ on objects $F$ with the property that, for every $p\in ss(F)^{\circ}$ such that $\pi:SS(F)\to M$ has constant rank on a neighborhood of $p$, there exists a submanifold $N$ and $L\in\mathcal{C}$ such that $F\simeq L_{Y}[dim(N)]\in(\mu sh_{T^{*}M})_{p}$ and $\tau^{\geq 1}F\simeq 0$ (resp.
+$\tau^{\leq-1}F\simeq 0$).
+Definition 6.7 can be equivalently expressed using the microstalk functor of Corollary 1.3.
+Definition 6.8.
+Consider the following full subcategories of $sh(M)_{\mathbb{C}-c}$.
+$^{\mu}sh(M)_{\mathbb{C}-c}^{\leq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{\circ}\implies\omega_{p}^{-1}F[-n]\in\mathcal{C}^{\leq 0}\}$
+$^{\mu}sh(M)_{\mathbb{C}-c}^{\geq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{\circ}\implies\omega_{p}^{-1}F[-n]\in\mathcal{C}^{\geq 0}\}$
+Proposition 6.9.
+Proof.
+Follows immediately from Corollary 5.11. ∎
 Lemma 6.10.
-The heart of a $t$-structure on a sheaf of stable categories is a stack of abelian categories.
-7.2. The perverse t-structure on constructible sheaves
-We now review from [15, Sec. 10.3] the microlocal description of the perverse $t$-structure on constructible sheaves.
-Let $M$ be a complex manifold.
-For a Lagrangian subset $\Lambda\subset T^{*}M$, we write $\Lambda^{cl}$ for the locus of smooth points of $\Lambda$ where the map $\Lambda\to M$ has locally constant rank.
-For a point $p\in\Lambda^{cl}$, we write $r(p)$ for this rank.
-Fix a $t$-structure $C^{\leq 0},C^{\geq 0}$ on our coefficient category $C$.
-We write [15, Def. 10.3.7] using the notation of (10):
-Definition 7.7.
-[15, Def. 10.3.7] Consider the following full subcategories of $sh(M)_{\mathbb{C}-c}$.
-${}^{\mu}sh(M)_{\mathbb{C}-c}^{\leq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{cl}\implies\omega_{p}^{-1}F[-2r(p)]\in C^{\leq 0}\}$
-${}^{\mu}sh(M)_{\mathbb{C}-c}^{\geq 0}:=\{F\in sh(M)_{\mathbb{C}-c}\,|\,p\in ss(F)^{cl}\implies\omega_{p}^{-1}F[-2r(p)])\in C^{\geq 0}\}$
-It is proved in [15, Thm.
-10.3.12] that the above pair agrees with the usual definition of the perverse $t$-structure on $M$, and so in particular is a $t$-structure.
-[^5] The result is stated in [15] for $C$ the bounded derived category of modules over a ring and and $t$ the standard $t$-structure.
-However, the arguments given there (or in [3]) for the existence of the perverse $t$-structure depend only on the general properties of the six functor formalism, and the comparison between Definition 7.7 and the usual stalk/costalk wise definition of the perverse $t$-structure depends only on standard properties of microsupports.
-Lemma 7.8.
-If $\Lambda\subset T^{*}M$ is (possibly singular) subanalytic complex Lagrangian then Definition 7.7 induces a $t$-structure on $sh_{\Lambda}(M)$.
-Moreover, $sh_{\Lambda}(M)^{\heartsuit}$ is closed under extensions inside $sh(M)$.
+If $\Lambda\subset T^{*}M$ is (possibly singular) subanalytic complex Lagrangian and $\Omega\subseteq\mathbb{P}^{*}M$ is an open set, then Definition 6.7 induces a $t$-structure on
+$sh_{\Lambda\cup\pi^{-1}(\Omega^{c}),\mathbb{C}-c}(M)=\{F\in sh_{\mathbb{C}-c}(M)|ss(F)\cap\pi^{-1}(\Omega)\subseteq\Lambda\}.$
+Moreover, $sh_{\Lambda\cup\pi^{-1}(\Omega^{c}),\mathbb{C}-c}(M)^{\heartsuit}$ is closed under extensions inside $sh(M)$.
 Proof.
-We wish to apply Lemma 7.2.
-We should therefore show that the images of the truncation functors, when applied to objects $F\in sh_{\Lambda}(M)$, remain in $sh_{\Lambda}(M)$.
-Suppose not; consider any smooth Lagrangian point in $ss(\tau_{\leq 0}F)\setminus ss(F)$.
-(There must be such a point by the assumption that the microsupport is complex analytic Lagrangian.) But by construction, the truncation functors commute with $\omega_{p}$, whence it follows that $\omega_{p}^{-1}\tau_{\leq 0}F[-2r(p)]=\tau_{\leq 0}\omega_{p}^{-1}F[-2r(p)]=0$.
+By (1) of Lemma 6.2, we only need to check that $\tau_{\leq k}F$ is contained in the subcategory if $F$ is.
+Suppose not; then for any neighborhood $U$ of $\mathbb{C}^{*}\cdot p$, $ss(\tau_{\leq k}F)-\Lambda$ is non-empty.
+Since $\tau_{\leq k}F$ is constructible, $(ss(\tau_{\leq k}F)-\Lambda)\cap U$ must have a smooth Legendrian point $q$.
+But then the microstalk of $\tau_{\leq k}F$ at this point $q$ is the truncation of the microstalk of $F$, which is zero.
 A contradiction.
-The conclusion now follows from Lemma 7.2.
 Now suppose given $F^{\prime},F^{\prime\prime}\in sh_{\Lambda}(M)^{\heartsuit}$ and some extension $0\to F^{\prime}\to F\to F^{\prime\prime}\to 0$.
@@
 ∎
-We now review the results of Waschkies [27]. Consider $T^{\circ}M=T^{*}M\setminus M$, and the map $\pi:T^{\circ}M\to\mathbb{P}^{*}M$.
+Theorem 6.11.
+[42] The microlocal perverse t-structure in Definition 6.8 induces a perverse $t$-structure on $\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})$.
+Furthermore, for $\mathcal{F},\mathcal{G}\in[\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})]^{\heartsuit}$, the hom-sheaf
+$\mathcal{H}om_{\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})}(\mathcal{G},\mathcal{F})[\dim M]$
+is a perverse sheaf on the $T^{\circ}M$.
+Proof.
+By Proposition 5.9, the inclusion $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M,\mathbb{C}-c}\xhookrightarrow{\sim}\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})$ is an equivalence so we can work with $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M,\mathbb{C}-c}$.
+Any $\mathcal{F}\in\operatorname{\mathbb{P}\mu sh}_{P^{*}M,\mathbb{C}-c}(\Omega)$ tautologically belongs to $\operatorname{\mathbb{P}\mu sh}_{\Lambda}(\Omega)$ where $\Lambda\coloneqq\operatorname{supp}(\mathcal{F})$, so we could fix a support condition $\Lambda$.
+By Lemma 6.5, it’s enough to check it on an open cover.
+But since $\operatorname{\mathbb{P}\mu sh}_{\Lambda}$ is constructible, there exits a cover $\mathcal{U}$ such that each $\Omega$ satisfies $\operatorname{\mathbb{P}\mu sh}_{\Lambda}(\Omega)=(\operatorname{\mathbb{P}\mu sh}_{\Lambda})_{p}$ for some $p\in\Omega$, so we can check on stalks.
+By Corollary 5.8 and Proposition 6.9, but $(\operatorname{\mathbb{P}\mu sh}_{\Lambda})_{p}$ and the notion of microstalks are invariant under contact transform, so we can assume $\Lambda$ is in general position.
+In this case, by Theorem 5.3, $(\operatorname{\mathbb{P}\mu sh}_{\Lambda})_{p}\hookrightarrow(sh_{\Lambda})_{m}/loc_{m}$ has image given by the category in the previous Lemma 6.10, which we’ve to have a t-structure, and it clear induces a $t$-structure on the quotient by applying Lemma 6.2 to locloc.
+For the statement regarding the hom sheaf, let $\mathcal{F},\mathcal{G}\in(\pi_{*}\mu sh_{T^{\circ}M,\mathbb{C}-c})^{\heartsuit}$.
+Since perversity can be checked locally, we may pick some sheaves $F$ and $G$ representing $\mathcal{F}$ and $\mathcal{G}$.
+But in this case, by (10), we have the identification
+$\mathcal{H}om_{\pi_{*}\mu sh_{T^{\circ}M,\mathbb{C}-c}}(\mathcal{G},\mathcal{F})=\mathcal{H}om_{\pi_{*}\mu sh_{T^{\circ}M,\mathbb{C}-c}}(G,F)=\mu hom(G,F).$
+But then, [25, Corollary 10.3.20] implies that $\mathcal{H}om_{\mathbb{P}\mu sh_{V}}(\mathcal{G},\mathcal{F})[\dim M]=\mu hom(G,F)[\dim M]$ is perverse. ∎
+6.3. Perverse microsheaves on complex contact and symplectic manifolds
+We now define notion of perverse t-structure for microsheaves on complex symplectic and contact manifolds for the canonical microsheaves over $\mathcal{C}=R-mod$, for a discrete commutative ring $R$.
+We postpone the general discussion later in Definition 6.16 for conceptual clarity.
+Recall by Corollary 1.3, for any microsheaf $\mathcal{F}$ on $V$, there is a well-defined object $\omega_{p}^{-1}(F)\in\mathcal{C}$.
+Definition 6.12.
+Let $V$ be a contact manifold of complex dimension 2n12n-1.
+We define the pair of subcategories $((\mu sh_{\widetilde{V},\mathbb{C}-c})^{\leq 0},(\mu sh_{\widetilde{V},\mathbb{C}-c})^{\geq 0})$ by constraining the microstalks:
+$(\mu sh_{\widetilde{V},\mathbb{C}-c})^{\leq 0}:=\{\mathcal{F}\in\mu sh_{V,\mathbb{C}-c}\,|\,p\in\operatorname{supp}(\mathcal{F})^{\circ}\implies\omega_{p}^{-1}\mathcal{F}[-n]\in\mathcal{C}^{\leq 0}\}$
+$(\mu sh_{\widetilde{V},\mathbb{C}-c})^{\geq 0}:=\{\mathcal{F}\in\mu sh_{V,\mathbb{C}-c}\,|\,p\in\operatorname{supp}(\mathcal{F})^{\circ}\implies\omega_{p}^{-1}\mathcal{F}[-n]\in\mathcal{C}^{\geq 0}\}$
+We define similarly the corresponding notions for objects supported in some fixed (singular) Lagrangian, and define as always the corresponding notions on $V$ by pushforward.
+[^5] The $[n]$ is just a convention, set to match the usual conventions for perverse sheaves.
+Theorem 6.13.
+$(\mu sh_{V,\mathbb{C}-c})^{\leq 0}$, $(\mu sh_{V,\mathbb{C}-c})^{\geq 0}$ determine a $t$-structure on $\mu sh_{V,\mathbb{C}-c}$.
+In particular, $(\mu sh_{V,\mathbb{C}-c})^{\heartsuit}$ is a sheaf of abelian categories.
+Furthermore, for $\mathcal{F},\mathcal{G}\in\mu sh_{V,\mathbb{C}-c}^{\heartsuit}$, the sheaf of morphisms
+$\mathcal{H}om_{\mu sh_{V}}(\mathcal{G},\mathcal{F})[\frac{1}{2}\dim\widetilde{V}$]
+is a perverse sheaf on the symplectization $\widetilde{V}$.
+Proof.
+Per Lemma 6.5, a pair of subcategories $(\mathcal{C}^{\leq},\mathcal{C}^{\geq})$ being a t-structure can be checked on open covers.
+Similarly, being a perverse sheaf if also a local condition.
+Thus, we reduce to Theorem 6.11 by taking Darboux charts.
+∎
+We deduce Theorem 1.6 from the introduction.
+Proof of Theorem 1.6.
+By Theorem 5.16, there is an equivalence $(({\gamma_{\mathbb{C}}})_{*}\mathbb{P}\mu sh_{W\times\{0\}})^{\mathbb{C}^{*}}\simeq({\gamma_{\mathbb{C}}})_{*}\mu sh_{W}.$ The $\mathbb{C}^{*}$-action manifestly preserves the subspaces $\mu sh_{W,\mathbb{C}-c}^{\leq 0}$ and $\mu sh_{W,\mathbb{C}-c}^{\geq 0}$, so the $t$-structure passes to the invariants (taking hearts is a pullback ($\mathcal{F}^{\heartsuit}:=\mathcal{F}^{\leq 0}\cap\mathcal{F}^{\geq 0}$) hence a limit, and hence commutes with taking $G$-invariants which is also a limit).
+Now, we consider the perversity of sheaf Hom.
+We recall that objects here admits two different interpretations: as microsheaves on $\widetilde{{W\times\mathbb{C}}}$ and as microsheaves on the underlying real $W$ since Theorem 5.16 is needed to define $\mu sh_{W,\mathbb{C}-c}^{\heartsuit}$.
+Denote by $\widetilde{\mathcal{F}}$, $\widetilde{\mathcal{G}}$ for the former and $\mathcal{F}$, $\mathcal{G}$ for the latter, then we see from the contact case that $\mu hom(\widetilde{\mathcal{G}},\widetilde{\mathcal{F}})[\frac{1}{2}(\dim W)+1)]$ is a perverse sheaf on $\widetilde{W\times\mathbb{C}}$.
+However, to descend them onto $W$, we recall that such objects are assumed to be locally constant along the $\mathbb{C}^{*}$-action and we take $\mathbb{C}^{*}$-invariant to quotient the extra direction.
+This process drops the dimension by $1$ and we thus conclude that $\mu hom(\mathcal{G},\mathcal{F})[\frac{1}{2}\dim W]$ is perverse.
+∎
+Now, let $\mathcal{C}$ be any symmetric monoidal category and $(\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0})$ be a $t$-structure on $\mathcal{C}$.
+We explain how one can generalize Definition 6.12 to allow more flexibility on Maslov data.
+For this purpose, we recall the notion of constrained Maslov data from Section 4.7: A collection of subcategories (of $\mathcal{C}$) $\{\mathcal{D}_{i}\}$ is said to be anchored if the submonoid $\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$ fixing each $\mathcal{D}_{i}$ is a subgroup.
 Lemma 6.14.
-The natural map
-(19) $\mu sh^{pre}(\pi^{-1}(m))\to(\mathbb{P}\mu sh)_{m}$
-is an equivalence.
+Let $(\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0})$ be a $t$-structure on $\mathcal{C}$, then the collection $\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\}$ is anchored.
 Proof.
-We first check fully-faithfulness. It is equivalent to check that the natural map
-$\mu sh^{pre}(\pi^{-1}(m))\to(\pi_{*}\mu sh)_{m}$
-is fully-faithful.
-Given objects $K,L$ in the source, Waschkies [27, Sec.
-2.4] shows that $\hom(K,L)$ in the source is computed by $(\pi_{*}\mu\operatorname{Hom}(K,L))_{m}$, which also computes $\hom(K,L)$ in the target.
-We now check essential surjectivity.
-If $K\in(\mathbb{P}\mu sh)_{m}$, then it is represented by an object $\tilde{K}\in\mathbb{P}\mu sh(U),p\in U,$ which is representable by a sheaf at $p$.
-This means that (after possibly shrinking $U$) there is a complex-contact embedding $\chi:U\to V\subset\mathbb{P}^{*}N$ so that $\chi_{*}(K)=\chi_{*}(\tilde{K})\in(\pi_{*}\mu sh)_{\chi(m)}$ is in the image of $(\pi_{*}\mu sh^{pre})_{\chi(m)}\to(\pi_{*}\mu sh)_{\chi(m)}$.
-By standard quantization of sheaf kernels (see e.g.
-[27, Thm.
-4.2.1]) we have
-[…]
-where the vertical arrows are (19). The claim follows. ∎
+It is sufficient to show that, for $x\in\operatorname{Pic}(\mathcal{C})$, tensoring $x\otimes(-)$ fixes $\mathcal{C}^{\leq 0}$ if and only if tensoring its inverse $x^{-1}\otimes(-)$ fixes $\mathcal{C}^{\geq 0}$.
+We show the “if” direction, since the argument is symmetric.
+As remarked in [30, Remark 1.2.1.3], an object $a\in\mathcal{C}$ is in $\mathcal{C}^{\geq 1}$ if $\operatorname{Hom}(b,a)=0$ for all $b\in\mathcal{C}^{\leq 0}$.
+Thus we consider such objects and compute that $\operatorname{Hom}(b,x^{-1}\otimes a)=\operatorname{Hom}(x\otimes b,a)=0$.
+This implies that $x^{-1}\otimes(-)$ fixes $\mathcal{C}^{\geq 1}$ but, since $[n]$ commutes with tensor, it fixes $\mathcal{C}^{\geq n}$ for all $n\in\mathbb{Z}$.
+∎
+Since the collection of subcategories $\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\}$ is anchored, there is a notion of constrained gradings and orientations defined in Definition 4.34.
+As explained by Equation 17, the canonical grading from the complexes structure induces a $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\})$-grading.
+Definition 6.15.
+Let $o$ be a $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\})$-orientation, i.e., a lifting of the induced grading to a $\mathcal{C}$-Maslov data.
+We denote by $\mu sh_{\widetilde{V},o}$ the associated sheaf of microsheaves on $\widetilde{V}$ and similar notations for the subsheaves with support conditions.
+Let $L$ be a complex Legendrian.
+Then Lemma 3.12 allows us to choose constrained secondary Maslov data, in the sense defined above Corollary 4.36, and the cited Proposition implies that, for any constructible microsheaf $\mathcal{F}\in\mu sh_{\widetilde{V},\mathbb{C}-c,o}$ and any smooth point $p\in\operatorname{supp}(\mathcal{F})^{\circ}$, whether the microstalk $\omega_{p}^{-1}\mathcal{F}$ is in $\mathcal{C}^{\leq 0}$ or $\mathcal{C}^{\geq 0}$ (and hence all their shiftings) is a well-defined notion.
+Thus, we have the following generalization of Definition 6.12:
+Definition 6.16.
+Fix a $t$-structure on the coefficient category $\mathcal{C}$.
+Let $V$ be a contact manifold of complex dimension 2n12n-1.
+Fix any $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\})$-orientation data $o$.
+We define subcategories of $\mu sh_{\widetilde{V},\mathbb{C}-c,o}$ by constraining the microstalks to respect the $t$-structure of $\mathcal{C}$:
+$(\mu sh_{\widetilde{V},\mathbb{C}-c,o})^{\leq 0}:=\{\mathcal{F}\in\mu sh_{V,\mathbb{C}-c,o}\,|\,p\in\operatorname{supp}(\mathcal{F})^{\circ}\implies\omega_{p}^{-1}\mathcal{F}[-n]\in\mathcal{C}^{\leq 0}\}$
+$(\mu sh_{\widetilde{V},\mathbb{C}-c,o})^{\geq 0}:=\{\mathcal{F}\in\mu sh_{V,\mathbb{C}-c,o}\,|\,p\in\operatorname{supp}(\mathcal{F})^{\circ}\implies\omega_{p}^{-1}\mathcal{F}[-n]\in\mathcal{C}^{\geq 0}\}$
+We define similarly the corresponding notions for objects supported in some fixed (singular) Lagrangian, and define as always the corresponding notions on $V$ by pushforward.
+As perversity can be checked locally, the exact same argument of Theorem 6.17 and Theorem 1.6 implies the following theorem.
 Theorem 6.17.
-[27, Thm.
-5.1.5] Assume $M$ is complex analytic, $\Lambda\subset T^{\circ}M$ is a $\mathbb{C}^{*}$ invariant (singular) Lagrangian, and $\Lambda/\mathbb{C}^{*}\to M$ is a finite map.
-Then the natural map of sheaves of categories
-$sh_{\Lambda}/loc\to\mathbb{P}\mu sh_{\Lambda}$
-is an equivalence.
-Proof.
-We check at stalks. There are natural maps:
-$(sh_{\Lambda}/loc)_{m}=sh_{\Lambda,m}/loc_{m}\to\mu sh^{pre}(\pi^{-1}(\tilde{m}))\to(\mathbb{P}\mu sh_{\Lambda})_{\tilde{m}}$
-where $m\in M$ and $\tilde{m}\in\Lambda$ its (possibly disconnected) preimage.
-In [27, Thm.
-5.1.5], Waschkies constructs the inverse using the refined microlocal cutoff of D’Agnolo [6], and shows that it is fully faithful essentially by arguing that all categories in question have homs given by the $\mu hom$ functor, similarly to [15, Thm.
-6.1.2].
-∎
+Fix a $t$-structure on the coefficient category $\mathcal{C}$.
+Let $V$ be a contact manifold of complex dimension 2n12n-1.
+Or, similarly, let $W$ be an exact complex symplectic manifold of complex dimension $2n$ with a $\mathbb{C}^{*}$-action of weight $1$.
+Given any $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0}\})$-orientation data $o$ on $V$ (resp. $W$), the pairs
+$\left((\mu sh_{V,\mathbb{C}-c,o}\right)^{\leq 0},(\mu sh_{V,\mathbb{C}-c,o})^{\geq 0})\left(\text{resp.}\ \left(\left((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c}\right)^{\geq 0}\right),\left((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c}\right)^{\leq 0}\right)$
+determine a $t$-structure on $\mu sh_{V,\mathbb{C}-c,o}$ (resp.
+$(\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c}$).
+Furthermore, for $\mathcal{F},\mathcal{G}\in\mu sh_{V,\mathbb{C}-c,o}^{\heartsuit}$ $\left(resp.\left((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c})^{\heartsuit}\right)\right)$, the sheaf of morphisms
+$\mathcal{H}om_{\mu sh_{V}}(\mathcal{G},\mathcal{F})[\frac{1}{2}\dim\widetilde{V}]\left(\text{resp.}\ \mathcal{H}om_{(\gamma_{\mathbb{C}})_{*}\mu sh_{W}}(\mathcal{G},\mathcal{F})[\frac{1}{2}\dim W]\right)$
+is a perverse sheaf on the symplectization $\widetilde{V}$ (resp. $W$).
 Remark 6.18.
-Note that one must sheafify quotients, so $(sh_{\Lambda}/loc)(U)$ is generally not $sh_{\Lambda}(U)/loc(U)$.
-However, both $sh_{\Lambda}$ and $loc$ are constructible, so for small enough open neighborhood $U$ of any $m\in M$, we have $(sh_{\Lambda}/loc)(U)=(sh_{\Lambda}/loc)_{m}=(sh_{\Lambda})_{m}/loc_{m}=sh_{\Lambda}(U)/loc(U)$.
+A $t$-structure is said to be nondegenerate if $\bigcap\mathcal{C}^{\leq 0}=0=\bigcap\mathcal{C}^{\geq 0}$.
+By co-isotropicity of microsupport, the vanishing of all microstalks implies the vanishing of an object.
+We conclude that if the $t$-structure on $\mathcal{C}$ is non-degenerate, then the $t$-structure on $\mu sh_{V,\mathbb{C}-c,o}$ is also non-degenerate.
 Remark 6.19.
-The analogue of Theorem 7.10 with complex analyticity hypotheses removed and with $\mathbb{C}^{*}$ replaced by $\mathbb{R}^{+}$ is also true (indeed easier: the cutoffs already in [15] are good enough and one does not need [6]); see [22, Lem 6.7, Prop.
-6.9].
-Corollary 7.13.
-In the situation of Theorem 7.10, there is a $t$-structure on $\pi_{*}\mu sh_{\Lambda}$ characterized by the formulas in Definition 7.7.
-Proof.
-For any $U\subset M$, the $t$-structure descends to $sh_{\Lambda}(U)/loc(U)$ by Lemma 7.2 and Lemma 7.8.
-For $p\in T^{\circ}M$, the functor $\omega_{p}$ descends to $sh_{\Lambda}(U)/loc(U)$, and clearly the quotient $t$-structure on $sh_{\Lambda}(U)/loc(U)$ is characterized exactly as in Definition 7.7.
-∎
-Remark 7.14.
-In [27, Thm.
-6.1.5], this is extended to arbitrary $\Lambda\subset\mathbb{P}\mathbb{T}^{*}M$ by noting that complex contact transformation preserves the conditions of Definition 7.7 above, using the degree shift formula [15, prop.
-7.4.6].
-In our discussion below, we don’t explicitly use this complex contact invariance of Definition 7.7 to construct the microlocal perverse $t$-structure, but would need it to check that on cotangent bundles, we have defined the usual perverse $t$-structure.
-7.3. Perverse microsheaves on complex contact and symplectic manifolds
-Definition 7.15.
-Fix a $t$-structure on the coefficient category $C$, e.g.
-the standard $t$-structure on the category of $R$-modules, for a commutative ring $R$.
-Let $V$ be a contact manifold of complex dimension $12n-$ 1.
-Fix any $(C,C^{\leq 0})$ orientation data $o$.
-(If $C=R-mod$, we may always choose the canonical orientation data.)
-We define subcategories of $\mu sh_{\tilde{V},\mathbb{C}-c,o}$ by constraining the microstalks to respect the $t$-structure of $C$. In the notation of Definition 5.11:
-$(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{\leq 0}:=(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{C^{\leq 0}[2n]}$
-$(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{\geq 0}:=(\mu sh_{\tilde{V},\mathbb{C}-c,o})^{C^{\geq 0}[2n]}$
-We define similarly the corresponding notions for objects supported in some fixed (singular) Lagrangian, and define as always the corresponding notions on $V$ by pushforward.
-The $[2n]$ is just a convention, set to match the usual conventions for perverse sheaves.
-Theorem 7.16.
-$(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\leq 0}$, $(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\geq 0}$ determine a $t$-structure on $\mathbb{P}\mu sh_{V,\mathbb{C}-c,o}$.
-Proof.
-Being a $t$-structure is an objectwise condition.
-So it suffices to show, for every fixed singular Lagrangian $\Lambda$, that objects supported in $\Lambda$ satisfy the corresponding conditions; it suffices therefore to show $(\mathbb{P}\mu sh_{\Lambda,o})^{\leq 0}$, $(\mathbb{P}\mu sh_{\Lambda,o})^{\geq 0}$ define a $t$-structure on $\mathbb{P}\mu sh_{\Lambda,o}$.
-We may check locally.
-Thus we may assume $\Lambda$ is an arbitrarily small and contractible neighborhood of some point $p\in\Lambda$.
-By the Kashiwara–Kawai’s general position theorem (see [13, Sec.
-1.6]), we may choose a complex manifold $M$ and an embedding $\iota:\Lambda\to\mathbb{P}T^{*}M$ such that $\Lambda\to M$ is finite, hence an embedding at generic points of $\Lambda$.
-Let $\eta$ be the Maslov datum used on $V$; recall that the grading part is canonical.
-Writing $\phi$ for the fiber Maslov datum on $T^{*}M$, we use the canonical identification of the ‘grading’ parts of $\eta$ and $\iota^{*}\phi$, and any arbitrary homotopy to identify the orientation data parts (which certainly exists since $\Lambda$ is contractible).
-Thus we have an identification $\iota^{*}\mu sh_{\iota({\Lambda}),o}\cong\mu sh_{\Lambda,o}$.
-Per 5.10 and Theorem 7.10, this identification intertwines the microstalk functors used to characterize the perverse $t$-structure on the left hand side with the functors we used in Definition 5.11, which therefore defines a $t$-structure as well.
-∎
-We immediately deduce Theorem 1.4 from the introduction.
-Proof of Theorem 1.4.
-As noted in Remark 6.4, $\mathbb{C}^{*}$ acts on $W\times\mathbb{C}$ by contactomorphism fixing $W\times\{0\}$ set-wise.
-Hence the inclusion $({\gamma_{\mathbb{C}}})_{*}\mathbb{P}\mu sh_{W\times\{0\}}\hookrightarrow({\gamma_{\mathbb{C}}})_{*}\pi_{*}\mu sh_{\widetilde{W\times\mathbb{C}}}$ is stable under this action.
-Taking $\mathbb{C}^{*}$-invariants, we get a fully faithful inclusion $\mu_{\mathbb{C}}sh_{W}=(({\gamma_{\mathbb{C}}})_{*}\mathbb{P}\mu sh_{W\times\{0\}})^{\mathbb{C}^{*}}\hookrightarrow(({\gamma_{\mathbb{C}}})_{*}\pi_{*}\mu sh_{\widetilde{W\times\mathbb{C}}})^{\mathbb{C}^{*}}=({\gamma_{\mathbb{C}}})_{*}\mu sh_{W}$.
-The $\mathbb{C}^{*}$-action manifestly preserves the subspaces $\mu sh_{W,\mathbb{C}-c}^{\leq 0}$ and $\mu sh_{W,\mathbb{C}-c}^{\geq 0}$, so the $t$-structure passes to the invariants (taking hearts is a pullback ($\mathcal{F}^{\heartsuit}:=\mathcal{F}^{\leq 0}\cap\mathcal{F}^{\geq 0}$) hence a limit, and hence commutes with taking $G$-invariants which is also a limit).
-∎
-As a consequence of Theorem 7.16 and Lemma 7.6, we conclude:
-Corollary 7.17.
-$(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\heartsuit}$ is a stack of abelian categories; in particular, the global sections $(\mathbb{P}\mu sh_{V,\mathbb{C}-c,o})^{\heartsuit}(V)$ is an abelian category.
-Remark 7.18.
-A $t$-structure is said to be nondegenerate if $\bigcap C^{\leq 0}=0=\bigcap C^{\geq 0}$.
-By co-isotropicity of microsupport, the vanishing of all microstalks implies the vanishing of an object.
-We conclude that if the $t$-structure on $C$ is non-degenerate, then the $t$-structure on $\mathbb{P}\mu sh_{V,\mathbb{C}-c,o}$ is also non-degenerate.
-Remark 7.19.
-Choose a stable (not necessarily presentable) subcategory $D\subset C$ to which the $t$-structure restricts.
-Require $(C,C^{\leq 0},D)$-orientation data.
-Then it is evident from the definitions that the truncation functors preserve, hence define a $t$-structure on, the subcategory of objects with all microstalks in $D$, characterized by the same formulas save only with e.g.
-$C^{\leq 0}$ replaced by $C^{\leq 0}\cap D$.
+Choose a stable (not necessarily presentable) subcategory $\mathcal{D}\subset\mathcal{C}$ to which the $t$-structure restricts.
+Require $(\mathcal{C},\{\mathcal{C}^{\leq 0},\mathcal{C}^{\geq 0},\mathcal{D}\})$-orientation data.
+Then it is evident from the definitions that the truncation functors preserve, hence define a $t$-structure on, the subcategory of objects with all microstalks in $\mathcal{D}$, characterized by the same formulas save only with e.g.
+$\mathcal{C}^{\leq 0}$ replaced by $\mathcal{C}^{\leq 0}\cap\mathcal{D}$.
 For instance, we can take various bounded categories e.g.
-$D=C^{+}=\bigcup C^{\geq n}$, or $D=C^{-}=\bigcap C^{\leq n}$, or $D=C^{b}=C^{+}\cap C^{-}$, or ask the microstalks to be compact objects $D=C^{c}$.
+$\mathcal{D}=\mathcal{C}^{+}=\bigcup\mathcal{C}^{\geq n}$, or $\mathcal{D}=\mathcal{C}^{-}=\bigcap\mathcal{C}^{\leq n}$, or $\mathcal{D}=\mathcal{C}^{b}=\mathcal{C}^{+}\cap\mathcal{C}^{-}$, or ask the microstalks to be compact objects $\mathcal{D}=\mathcal{C}^{c}$.
 Remark 6.20.
 Fix $D\subset C$ as above, and assume $D^{\heartsuit}$ is Artinian.
-(E.g., $C=k-mod$ for a field $k$ and $D=C^{c}$.) Then the full subcategory of $(\mathbb{P}\mu sh_{V,o})^{D,\heartsuit}$ on objects with finitely stratified (rather than just locally finite) support is also Artinian.
+(E.g., C=kmodC=k-mod for a field $k$ and $D=C^{c}$.) Then the full subcategory of $(\mu sh_{V,\mathbb{C}-c,o})^{D,\heartsuit}$ on objects with finitely stratified (rather than just locally finite) support is also Artinian.
 Indeed, any descending chain must have eventually stabilizing microsupports; we may restrict attention to one each in the finitely many connected components of the smooth locus of the support, hence by some point, all have stabilized.
@@
 Let $V$ be a complex contact manifold and let $L\subset V$ be a complex Legendrian.
-Then $\mathbb{P}\mu sh_{V,L}(-)$ is a sheaf of stable categories while $\mathbb{P}erv_{V,L}(-)$ is a sheaf of abelian categories.
+Then $\mu sh_{V,L}(-)$ is a sheaf of stable categories while $\mathbb{P}erv_{V,L}(-)$ is a sheaf of abelian categories.
 Be warned however that $D(\mathbb{P}erv_{V,L}(-))$ is only a presheaf of stable categories.
-In particular, the natural map $D(\mathbb{P}erv_{V,L}(-))\to\mathbb{P}\mu sh_{V,L}(-)$ may restrict to an equivalence on stalks without being an equivalence on global sections.
+In particular, the natural map $D(\mathbb{P}erv_{V,L}(-))\to\mu sh_{V,L}(-)$ may restrict to an equivalence on stalks without being an equivalence on global sections.
 A very special case: let $V=T^{*}S^{2}\times\mathbb{\mathbb{C}}$ and $L=0_{S^{2}}\times\{0\}$.
-Then $\mathbb{P}\mu sh_{V,L}(L)=loc(S^{2})\otimes loc(\mathbb{C}^{*})$ while $\mathbb{P}erv_{V,L}(L)=vect_{\mathbb{C}}\otimes loc(\mathbb{C}^{*})$, due to $S^{2}$ being simply connected.
-Similarly, $\mu_{\mathbb{C}}sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})=loc(S^{2})$ while $\mu_{\mathbb{C}}sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})^{\heartsuit}=vect_{\mathbb{C}}$.
+Then $\mu sh_{V,L}(L)=loc(S^{2})\otimes loc(\mathbb{C}^{*})$ while $\mathbb{P}erv_{V,L}(L)=vect_{\mathbb{C}}\otimes loc(\mathbb{C}^{*})$, due to $S^{2}$ being simply connected.
+Similarly, $\mu sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})=loc(S^{2})$ while $\mu sh_{T^{*}\mathbb{S}^{2},0_{S^{2}}}(0_{S^{2}})^{\heartsuit}=vect_{\mathbb{C}}$.
+Appendix A Existence of orientation data for real symplectic manifolds
+            by Sanath Devalapurkar
+We keep the notation of Section 3.
+We write $\operatorname{Sq}^{2}:B(\mathbb{Z}/2)\to B^{3}(\mathbb{Z}/2)$ for the map between infinite loop spaces representing the second Steenrod operation $\operatorname{Sq}^{2}:H^{*}(-;\mathbb{Z}/2)\rightarrow H^{*+2}(-;\mathbb{Z}/2)$.
+Lemma A.1 ( [9, (2.3)]).
+The connecting morphism for the exact triangle
+$B^{2}(\mathbb{Z}/2)\to\tau_{\leq 2}(U/O)\to\tau_{\leq 1}(U/O)=B\mathbb{Z}\to$
+is the composition $B\mathbb{Z}\to B(\mathbb{Z}/2)\xrightarrow{\operatorname{Sq}^{2}}B^{3}(\mathbb{Z}/2)$. ∎
+We consider the following diagram:
+[…]
+The map from $U(1)\to B\sqrt{SU}$ is the composition $U(1)\to BSU\to B\sqrt{SU}$, where the first map is the connecting map of the fiber sequence $SU\to U\xrightarrow{det}U(1)$.
+The bottom row is a fiber sequence, so the map $\alpha$ is induced by the fact that the composition $B\sqrt{SU}\to\pi_{\leq 2}B(U/O)\simeq B^{2}\mathbb{Z}$ is null.
+To establish the existence of the dotted arrow $U(1)\to B\mathbb{Z}$, it is enough to prove that the composition $B(SU)\to B(\sqrt{SU})\to B^{3}(\mathbb{Z}/2)$ is null (since $U(1)\to B\sqrt{SU}$ factors through $B(SU)\to B(\sqrt{SU})$).
+To this end, consider the fiber sequence $SU\to U\to U(1)$.
+Since $U(1)=\tau_{\leq 2}U$, it follows that $SU=\tau_{\geq 3}U$.
+Hence $B(SU)$ is $3$-connected, so any map into $B^{3}(\mathbb{Z}/2)$ is null.
+Lemma A.2.
+The map $\alpha:B\sqrt{SU}\rightarrow B^{3}(\mathbb{Z}/2)$ factors as a map of infinite loop spaces
+$B\sqrt{SU}\xrightarrow{B\det}B(\mathbb{Z}/2)\xrightarrow{\operatorname{Sq}^{2}}B^{3}(\mathbb{Z}/2).$
+Proof.
+This follows by contemplating the diagram:
+[…]
+The commutativity of the leftmost square comes from the fiber sequences
+$U{U}U(1){{U(1)}}BSU{{B\sqrt{SU}}}$$\scriptstyle{\operatorname{det}}$$scriptstyle{z\mapsto z^{2}}$$\scriptstyle{\operatorname{det}^{2}}$
+and the rightmost square comes from (25).
+It remains to explain the commutativity of the bottom triangle.
+Since $B(SU)\to B(\sqrt{SU})\to B(\mathbb{Z}/2)$ is a fiber sequence, it is enough to prove that the composition $B(SU)\to B(\sqrt{SU})\to B^{3}(\mathbb{Z}/2)$ is null, which was proved above.
+∎
+Proposition A.3.
+A choice of null-homotopy of $\operatorname{Sq}^{2}:B\mathbb{Z}/2\to B^{3}\mathbb{Z}/2$ gives a section of the forgetful map from grading/orientation data to grading data.
+Proof.
+A grading is a lift of the natural map $X\to BU$ to some $f:X\to B\sqrt{SU}$.
+Giving orientation data is giving a null-homotopy of the map $\alpha\circ f$, which by Lemma A.2, can be written as $\operatorname{Sq}^{2}\circ B\det\circ f$.
+∎
+While $\operatorname{Sq}^{2}$ is nonzero as a map of infinite loop spaces (as the Steenrod square is a nontrivial operation on cohomology), it is null as a map of spaces.
+Indeed, under the identifications
+(26) $\pi_{0}\left(\operatorname{Map}(B(\mathbb{Z}/2),B^{3}(\mathbb{Z}/2)\right)=H^{3}(B(\mathbb{Z}/2);\mathbb{Z}/2)=H^{3}(\mathbb{R}\mathbb{P}^{\infty},\mathbb{Z}/2)=\mathbb{Z}/2\langle w_{1}^{3}\rangle,$
+the (homotopy class of the) map $\operatorname{Sq}^{2}:B(\mathbb{Z}/2)\rightarrow B^{3}(\mathbb{Z}/2)$ corresponds to the element $\operatorname{Sq}^{2}(w_{1})\in\mathbb{Z}/2\langle w_{1}^{3}\rangle$.
+But $\operatorname{Sq}^{2}(w_{1})=0$, since Steenrod squares have the well-known property that $\operatorname{Sq}^{n}(x)=0$ if $n>\deg(x)$.
+Now, the homotopy classes of null-homotopies of $\operatorname{Sq}^{2}$ are given by $B(\mathbb{Z}/2),B^{2}(\mathbb{Z}/2)]=\mathrm{H}^{2}(\mathbb{R}\mathbb{P}^{\infty},\mathbb{Z}/2)=\mathbb{Z}/2$.
+Consider the inclusion $O\to\sqrt{SU}$; to study secondary orientation data, we will be interested in null-homotopies of the composition $\alpha^{\prime}:BO\to B\sqrt{SU}\xrightarrow{\alpha}B^{3}(\mathbb{Z}/2\mathbb{Z})$.
+The space of such null-homotopies is a torsor for $\operatorname{Map}(BO,B^{2}(\mathbb{Z}/2\mathbb{Z}))$, the homotopy classes of which are $H^{2}(BO,\mathbb{Z}/2\mathbb{Z})=\mathbb{Z}/2\langle w_{1}^{2},w_{2}\rangle$.
+One such null-homotopy arises from the null-homotopy of the factorization $O\to\sqrt{SU}\to\sqrt{SU}/O=B^{2}O\xrightarrow{Bw_{1}}B^{2}(\mathbb{Z}/2\mathbb{Z})$; let us denote this null-homotopy by $\tau$.
+Two more such null-homotopies arise by noticing that $\alpha^{\prime}$ factors through $BO\to B\sqrt{SU}\to B(\mathbb{Z}/2)\xrightarrow{\operatorname{Sq}^{2}}B^{3}\mathbb{Z}/2$, and then composing with one of the two null-homotopies of $\operatorname{Sq}^{2}$.
+Lemma A.4.
+The two null-homotopies of $\alpha^{\prime}:BO\to B^{3}(\mathbb{Z}/2\mathbb{Z})$ induced by composition with null-homotopies of $\operatorname{Sq}^{2}$ represent the classes $\tau+w_{2}$ and $\tau+w_{2}+w_{1}^{2}$.
+Proof.
+Denote by $\bar{\tau}$ the null-homotopy $BU\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$ by pre-composing $\tau$ with $BU\rightarrow BO$.
+As the pullback map $H^{2}(BO,\mathbb{Z}/2)\rightarrow H^{2}(BU,\mathbb{Z}/2)$ is given by $w_{2}\mapsto c_{1}$ and $w_{1}^{2}\mapsto 0$, it is sufficient to show that the two null-homotopies both go to $\bar{\tau}+c_{1}$ after further composing to $BU\rightarrow BO\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$.
+Thus, we consider the composition $U\rightarrow O\to\sqrt{SU}\to\sqrt{SU}/O$.
+The commutative diagram right above Definition 3.1 provides the following:
+[…]
+We note that the dashed map exists, since any map of the form $BSp\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$ factorizes to $BSp\rightarrow\tau_{\leq 3}BSp\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$ which is canonically homotopic as $\tau_{\leq 3}BSp=0$.
+In particular, the two homotopies induced from null-homotopy of $\alpha$ becomes the same after composing with $BU\rightarrow BO$ and is given by the fiber sequence $BSp\rightarrow B(Sp/U)\rightarrow B^{2}U$.
+On the other hand, the left three horizontal maps form a map between fiber sequences.
+Thus, $\bar{\tau}$ comes from post-composing the fiber sequence $BU\rightarrow BSp\rightarrow B(Sp/U)$.
+By the proof of Proposition 3.10, we see that their difference is given by $BU\xrightarrow{c_{2}}B^{2}(\mathbb{Z})\rightarrow B^{2}(\mathbb{Z}/2\mathbb{Z})$, which is exactly what we want.
+∎
+We write $\nu_{+}$ for the null-homotopy with class $\tau+w_{2}$ and $\nu_{-}$ for the one with class $\tau+w_{2}+w_{1}^{2}$.
+Lemma A.5.
+If $\nu_{\pm}$ is used to define orientation data on a symplectic manifold $X$, then secondary orientation data on a Lagrangian $L$ is a $Pin_{\pm}$ structure on $L$.
+Proof.
+Follows from Lemma A.4 by arguing as in Lemma 3.13. ∎
+Lemma A.6.
+Fix a stable quaternionic bundle $X\to BSp$.
+Then, any grading/orientation datum obtained from applying Proposition A.3 to the canonical grading of Definition 3.8 is canonically identified with the canonical grading/orientation datum of Definition 3.8.
+Proof.
+Any map $BSp\to\tau_{\leq 3}B(U/O)$ canonically factors through $\tau_{\leq 3}BSp=0$; in particular, the space of such maps is contractible.
+We defined the canonical grading/orientation data by taking the corresponding null-homotopy of the composition $X\to BSp\to BU\to\tau_{\leq 3}B(U/O)$.
+Meanwhile the grading/orientation data from Proposition A.3 are induced by choices of null-homotopy of the map $\alpha:B\sqrt{SU}\rightarrow B^{3}(\mathbb{Z}/2)$.
+Any such null-homotopy induces a null-homotopy out of $BSp$ by the pre-composition
+$BSp\rightarrow B\sqrt{SU}\rightarrow B^{3}(\mathbb{Z}/2)$
+as explained in Lemma 3.2. ∎
+Appendix B $t$-structures on Fukaya categories
+Here we translate our main results across the sheaf/Fukaya correspondence of [11], in order to construct $t$-structures on Fukaya categories of certain complex exact symplectic manifolds with contracting weight 11 $\mathbb{C}^{*}$ action, such as conic symplectic resolutions and moduli of Higgs bundles.
+Recall that a Liouville manifold is a (real) exact symplectic manifold $(W,\lambda)$ which is modeled at infinity on the symplectization of a contact manifold.
+The negative flow of the Liouville vector field $Z$ (defined by $\lambda=d\lambda(Z,\cdot)$) retracts $W$ onto a compact subset $\mathfrak{c}_{W}$ called the core.
+Fix if desired some larger closed conic subset $\Lambda\supset\mathfrak{c}_{W}$.
+We say an exact Lagrangian $L\subset W$ is admissible if it is closed and, outside a compact set, it is conic and disjoint from $\Lambda$.
+For example, when $Z$ is Smale and gradientlike for a Morse function, the ascending trajectories from maximal index critical points (“cocores”) are admissible.
+Any admissible Lagrangian disk which meets $\Lambda$ transversely at a single smooth point is termed a generalized cocore; we say a set $\{\Delta_{\alpha}\}$ of generalized cocores is complete if it meets every connected component of the smooth locus of $\Lambda$.
+If $Z$ is gradient-like for a Morse-Bott function, then $\Lambda$ is known to admit a complete set of generalized cocores.
+Fix grading and orientation data.
+We recall that one source of such data is a polarization of the stable symplectic normal (or equivalently tangent) bundle as explained in Section 3 below; see also [11, Sec.
+5.3].
+Then one can define a partially wrapped Fukaya category $Fuk(W,\partial_{\infty}\Lambda)$ [12].
+Objects are provided admissible Lagrangians equipped structures corresponding to the grading and orientation data.
+The object associated to a generalized cocore is unique up to grading shift.
+The completion of $Fuk(W,\partial_{\infty}\Lambda)$ with respect to exact triangles is generated by any any complete collection of generalized cocores [5, 12].
+We further complete with respect to idempotents, and still denote the resulting category $Fuk(W,\partial_{\infty}\Lambda)$.
+Definition B.1.
+Fix a collection $\{\Delta_{\alpha}\}$ of generalized cocores equipped with grading data. We define:
+$Fuk(W,\partial\Lambda)^{\geq 0}=\{L\,|\,\operatorname{Hom}(L,\Delta_{\alpha})\mbox{ is concentrated in degrees $\geq 0$}\}$
+$Fuk(W,\partial\Lambda)^{\leq 0}=\{L\,|\,\operatorname{Hom}(L,\Delta_{\alpha})\mbox{ is concentrated in degrees $\leq 0$}\}$
+It is natural to ask when Definition B.1 determines a $t$-structure.
+Suppose now that $W$ is a complex manifold and that $d\lambda=\operatorname{Re}\omega_{\mathbb{C}}$, for some complex symplectic structure $\omega_{\mathbb{C}}$ on $W$.
+As we have remarked above, and will explain in detail in Section 3 below, such a $W$ carries canonical grading data, which agrees with the grading induced by any stable complex Lagrangian polarization of the stable complex symplectic normal bundle (viewed as a real polarization of the real stable symplectic normal bundle), and any complex Lagrangian carries a canonical secondary grading.
+More generally, consider a real Lagrangian (or union of Lagrangians) $L\subset W$.
+Then $TL$ determines a section of $LGr(TW)|_{L}$.
+By a $g$-complex structure on $L$, we mean a simply connected neighborhood of $LGr_{\mathbb{C}}(TW)|_{L}\subset LGr(TW)|_{L}$ containing $TL$.
+An $g$-complex Lagrangian has a canonical secondary grading.
+Evidently any Lagrangian disk transverse to a complex Lagrangian admits a canonical $g$-complex structure.
+We recall the main result of [11].
+Assume $W$ is real analytic and Liouville, and $\Lambda\supset\mathfrak{c}_{W}$ is subanalytic, Lagrangian at smooth points, and admits a complete collection of generalized cocores $\{\Delta_{\alpha}\}$.
+Fix grading and orientation data coming from a stable polarization.
+[^6] The dependence on polarizations here and henceforth could be removed by a version of [35, Sec.
+11] for Fukaya categories.
+Then $Fuk(W,\partial_{\infty}\Lambda)\cong\mu sh_{\Lambda}(\Lambda)^{c,op}$ carrying $\Delta_{\alpha}$ to co-representatives of microstalk functors (here $(-)^{c}$ means that we take compact objects).
+Chasing definitions reveals that under [11], the normalized microstalks of Corollary 1.3 are carried to the canonically graded $\Delta_{\alpha}$ (up to some universal shift).
+Therefore:
+Corollary B.2.
+Suppose $(W,\lambda)$ is a complex exact symplectic manifold with weight 11 $\mathbb{C}^{*}$-action and $\Lambda\subset W$ a conic complex (singular) Lagrangian.
+Assume $(W,\operatorname{re}(\lambda))$ is Liouville and $\Lambda$ admits a complete collection of generalized cocores $\{\Delta_{\alpha}\}$.
+Fix grading and orientation data induced by a stable complex Lagrangian polarization of $(W,d\lambda)$.
+Then the equivalence of [11] carries the $t$-structure of Theorem 1.6 to a shift of Definition B.1, which therefore provides a $t$-structure.
+Also by Theorem 1.6 translated through [11], if $L,M\subset\mathfrak{c}_{X}$ are spin compact (necessarily conic) smooth Lagrangians, their Floer cohomology matches the cohomology of a (shifted) perverse sheaf supported on $L\cap M$.
+The hypotheses of Corollary B.2 are obviously satisfied for $W$ a cotangent bundle of a complex manifold.
+More generally, there are many examples of holomorphic symplectic manifolds with a weight 11 $\mathbb{C}^{*}$ action scaling the symplectic form – (coloop-free) quiver varieties, moduli of Higgs bundles, etc.
+– which satisfy all the hypotheses (the Liouville flow is gradientlike for the moment map for $S^{1}\subset\mathbb{C}^{*}$, which is Morse); cf.
+[44].
+The Fukaya category has the advantage that non-conic Lagrangians directly define objects, to which we thus have more direct geometric access.
+In particular, we can now construct objects in the heart.
+The point is that for index reasons, if $L\cup M$ is $g$-complex, then the Floer homology between $L$ and $M$ must be concentrated in degree zero.
+Similar considerations, for cotangent bundles, appear in [16].
+Thus:
+Corollary B.3.
+Retain the hypotheses of Corollary B.2.
+Let $L\subset W$ be an exact Lagrangian.
+Assume $L$ is compact, or more generally, that $\partial_{\infty}L$ wraps into $\partial_{\infty}\Lambda$ without passing through any $\partial_{\infty}\Delta_{\alpha}$.
+Suppose $L\cup\Delta_{\alpha}$ is $g$-complex for each $\alpha$.
+Then $L\in Fuk(W,\partial_{\infty}\Lambda)^{\heartsuit}=\mu sh_{\Lambda}(\Lambda)^{\heartsuit}$.
+The wrapping hypotheses appears to ensure that Hom in the wrapped Fukaya category is in fact computed without any wrapping, i.e.
+just by the Floer homology.
+One can imagine applying the corollary by taking an exact holomorphic Lagrangian is asymptotic to $\partial_{\infty}\Lambda$, and cutting off and straightening; such a process would plausibly produce a Lagrangian satisfying the wrapping hypothesis.
+We contemplate this process because such asymptotically conical exact holomorphic Lagrangians appear frequently in examples, but do not literally provide objects of $Fuk(X,\partial_{\infty}\Lambda)$.
+In a different direction, it was observed in [40] that the equivalence of microsheaf and Fukaya categories [11] remains true after enlarging the Fukaya category to contain unobstructed compact nonexact Lagrangians, of course taking coefficients in the Novikov field.
+As there, this observation is profitably combined with the fact [41] that holomorphic Lagrangians are unobstructed in hyperkähler manifolds.
+We conclude:
+Corollary B.4.
+Retain the hypotheses of Corollary B.2.
+Assume $W$ is hyperkähler and let $L\subset W$ be a compact holomorphic Lagrangian.
+Suppose $L\cup\Delta_{\alpha}$ is $g$-complex for each $\alpha$.
+Then $L\in Fuk(W,\partial_{\infty}\Lambda)^{\heartsuit}=\mu sh_{\Lambda}(\Lambda)^{\heartsuit}$, all categories taken with coefficients over the Novikov field.
+Corollary B.3 and Corollary B.4 can be stated more generally as defining fully faithful functors from appropriate abelian category of local systems on $L$ to the heart of the $t$-structure.